### v1

# Towards Human-AI Deliberation: Design and Evaluation of LLM-Empowered Deliberative AI for AI-Assisted Decision-Making

Abstract.
In AI-assisted decision-making, humans often passively review AI’s suggestion and decide whether to accept or reject it as a whole. In such a paradigm, humans are found to rarely trigger analytical thinking and face difficulties in communicating the nuances of conflicting opinions to the AI when disagreements occur. To tackle this challenge, we propose Human-AI Deliberation, a novel framework to promote human reflection and discussion on conflicting human-AI opinions in decision-making. Based on theories in human deliberation, this framework engages humans and AI in dimension-level opinion elicitation, deliberative discussion, and decision updates. To empower AI with deliberative capabilities, we designed Deliberative AI, which leverages large language models (LLMs) as a bridge between humans and domain-specific models to enable flexible conversational interactions and faithful information provision. An exploratory evaluation on a graduate admissions task shows that Deliberative AI outperforms conventional explainable AI (XAI) assistants in improving humans’ appropriate reliance and task performance. Based on a mixed-methods analysis of participant behavior, perception, user experience, and open-ended feedback, we draw implications for future AI-assisted decision tool design.
AI-Assisted Decision-making, Human-AI Collaboration, Appropriate Reliance
ccs: Human-centered computing Empirical studies in HCI
1. Introduction
With remarkable technological advancements, AI has been increasingly used to support people in making decisions in various domains, including criminal justice (Dodge et al., 2019; Dressel and Farid, 2018), admissions (Cheng et al., 2019; Zhang et al., 2023), financial investment (Green and Chen, 2019), and medical diagnosis (Cai et al., 2019b; Lee et al., 2021), among others. Concerns surrounding AI’s accuracy, safety, ethics, and accountability (Cai et al., 2019b; Lee et al., 2021; Binns et al., 2018) have led to the widespread adoption of the AI-assisted decision-making paradigm in real-world applications (Buçinca et al., 2021; Zhang et al., 2020; Wang and Yin, 2021; Bansal et al., 2021). In this paradigm, AI performs an assistive role by providing a recommendation, while human decision-makers can choose to accept or reject AI’s suggestion in their final decision (Lai et al., 2021).
Research in recent years, however, identified two challenges within the existing AI-assisted decision-making paradigm. First, a battery of empirical studies found that people rarely trigger analytical thinking (i.e., System 2 thinking (Kahneman, 2011)) when directly presented with AI’s suggestions (Buçinca et al., 2020; Rastogi et al., 2022; Bertrand et al., 2022). As a result, people frequently over-rely on the AI’s incorrect recommendations (over-reliance) or mistakenly ignore AI’s correct suggestions (under-reliance) (Buçinca et al., 2021; Ma et al., 2023; Wang and Yin, 2021). Although some solutions have been proposed, such as displaying AI explanations (Bansal et al., 2021) and forcing people to think (e.g., applying cognitive forcing functions (Buçinca et al., 2021)), the results are mixed at best (Lai et al., 2022; Poursabzi-Sangdeh et al., 2021) (Challenge 1). Second, instead of full consensus or complete divergence, human and AI decision rationales often exhibit partial alignment (Sivaraman et al., 2023; Wang and Yin, 2023). While they may concur on certain aspects, differences may persist on others (Miller, 2023). However, in current AI-assisted decision-making systems, AI consistently provides a fixed recommendation regardless of human thoughts and humans can only accept or reject AI’s recommendation as a whole (Lai et al., 2021), with limited support for resolving conflicts or engaging in a meaningful exchange of ideas with the AI system (Miller, 2023). For example, as shown in Figure 1 (a), when the human decision-maker’s prediction is inconsistent with the AI model’s recommendation and the human only partially agrees with the AI’s reasoning (e.g., explanation), existing AI-assisted decision-making interfaces do not support any communication between humans and AI regarding conflicting opinions. This limitation may impede the effective utilization of both human and AI knowledge, hindering collaborative and complementary human-AI team performance (Challenge 2).
Figure 1. An illustration of Human-AI Deliberation. (A) In traditional AI-assisted decision-making, when humans disagree with AI’s suggestions (and only find parts of AI’s reasons applaudable), it is difficult for humans to decide whether and how much to adopt AI’s suggestion. (B) In our proposed Human-AI Deliberation, we provide opportunities for the human and the AI model to deliberate on conflicting opinions by discussing related evidence and arguments. Then, AI and humans can update their thoughts (when find it necessary) and reach final predictions.
[…]
Deliberation, characterized by thoughtful and reasoned discussion, plays a pivotal role in facilitating constructive discourse and consensus-building across various contexts (Steenbergen et al., 2003; Bächtiger and Parkinson, 2019). Deliberation proves effective in facilitating various human decision-making tasks, including deliberative politics (Black et al., 2010; Habermas, 2005; Thompson, 2008), clinical diagnosis (Schaekermann et al., 2019; Inguaggiato et al., 2019; Preisz, 2019), criminal justice (Devine et al., 2001; Van Berkel et al., 2019), and more. It offers individuals an opportunity to rigorously evaluate different perspectives, including their own, which can potentially address Challenge 1 in AI-assisted decision-making. Moreover, deliberation allows participants to refine their viewpoints through informed discussions about opinion discrepancies (Peirce, 2014; Good, 1950; Gough, 2007). Such a structured process may also enable humans and AI to engage in detailed discussions, potentially mitigating Challenge 2. Despite the potential benefits of deliberation, how to design mechanisms to facilitate deliberative conversation between humans and AI and how deliberations influence AI-assisted decision-making remain to be explored.
In this paper, based on guidelines and theories in human deliberation (Steenbergen et al., 2003; Bächtiger and Parkinson, 2019; Bächtiger et al., 2009) and decision-making (Alvarez-Melis et al., 2021; Weed, 2005; Baumann et al., 2014), we introduce a novel framework, Human-AI Deliberation for AI-assisted decision-making (Figure 1 (b)). Instead of presenting a fixed AI suggestion for humans to accept or reject, our framework encourages humans to externalize their thoughts, enables an interactive deliberation process between humans and AI around the conflicting points of their opinions and rationales, and fosters dynamic, fine-grained updates of humans and AI’s decisions. The key component of this framework is Deliberative AI, which has the ability to locate viewpoint dissimilarities, stimulate comprehensive deliberation with human decision-makers, and make necessary changes, even compromises, in its own suggestion as the constructive discussion unfolds. To design such an AI assistant, we propose to integrate the strength of domain-specific models (for reliable assistant information generation) and Large Language Models (LLMs, for interactivity and conversation capability). We elaborate on the design of Human-AI Deliberation framework and Deliberative AI in Section 3 and demonstrate how to instantiate the framework in an illustrative task in Section 4.
We realize that Human-AI Deliberation might be especially suitable for handling task cases that are challenging for both humans and AI. If both parties are very good at the task cases (both have high decision accuracy) or one party is very proficient in the task cases, deliberation may not be necessary. Therefore, we study the impact of Human-AI Deliberation when the task case is challenging for both parties. Specifically, using our proposed Human-AI Deliberation as a research probe, we aim to explore the following research questions.
When dealing with task cases that are challenging for both humans and AI:
  • RQ1: How will Human-AI Deliberation affect task performance and humans’ reliance (and its appropriateness) on AI suggestions?
  • RQ2: How will Human-AI Deliberation affect humans’ perceptions of the AI and user experience?
  • RQ3: How will humans engage in the deliberation process when working with Deliberative AI?
  • RQ4: How will humans perceive the effectiveness of the proposed Human-AI Deliberation and what can be improved for future Human-AI Deliberation design?
To answer these questions, we conducted an exploratory study in the context of an illustrative task (college graduate admission). We recruited participants with graduate admissions experience (at least once admitted to a graduate program) on Prolific and asked them to predict an applicant’s chance of getting an offer based on the applicant’s profile. Based on a pilot study, we selected some task cases that were edge cases and therefore challenging for both humans and AI (e.g., their accuracy on these task cases was around 50-55%). We compared the proposed Deliberative AI with traditional explainable AI (XAI) and human alone baselines. Our experimental results revealed that Human-AI Deliberation has the potential to enhance decision accuracy and promote appropriate reliance on AI recommendations compared to traditional XAI assistants. We then derived valuable insights by analyzing participants’ questionnaire ratings, human behaviors, conversation logs, and open-ended feedback.
In summary, we make three contributions:
  • We propose a novel conceptual framework, Human-AI Deliberation, grounded in decision-making and deliberation theories, for humans and AI to deliberate their conflicting thoughts in a decision task. Accordingly, we design the architecture of Deliberative AI that is capable of deliberation with humans through a novel integration of domain-specific models and LLMs.
  • We demonstrate the instantiation of the framework and Deliberative AI in an illustrative task (college graduate admission), including the implementation of different components and interface design.
  • We conduct an exploratory study to gain an initial understanding of how Human-AI Deliberation might impact the decision-making process and how humans would perceive this novel AI assistance. Additionally, we demonstrate its potential to improve decision accuracy and promote appropriate human reliance on AI.
2. Related Work
2.1. AI-Assisted Decision-Making: Inadequacies and Challenges
Artificial Intelligence (AI) is increasingly used in decision-making across various domains (Ma et al., 2019, 2022b; Dastin, 2018; Dilsizian and Siegel, 2014; Khandani et al., 2010; Zhu and Ma, 2019). However, AI’s real-world applications are not infallible, still far from 100% accuracy (Shi et al., 2023; Gao et al., 2018; Ma et al., 2022a). This is especially concerning in high-stakes domains like medicine and criminal justice, leading to ethical and legal complexities (Cai et al., 2019b; Lee et al., 2021; Binns et al., 2018). To address this, the prevalent paradigm of AI-assisted decision-making has emerged, drawing substantial attention in the Human-Computer Interaction (HCI) and AI communities (Buçinca et al., 2021; Zhang et al., 2020; Wang and Yin, 2021; Bansal et al., 2021). In this paradigm, AI takes on a supportive role, offering recommendations for human decision-makers to accept or reject in their final decisions (Lai et al., 2021).
Research in AI-assisted decision-making encompasses various objectives, such as team performance, decision fairness, efficacy, efficiency, understanding of AI, trust, reliance on AI, and subjective experience (Zhang et al., 2020; Ma et al., 2023; Chiang et al., 2023; Wang et al., [n.d.]; Bansal et al., 2019a; Dressel and Farid, 2018; Wang and Yin, 2021; Cheng et al., 2019; Buçinca et al., 2021; Ma et al., 2023; Yang et al., 2020; Schemmer et al., 2023; He et al., 2023; Buçinca et al., 2020; Gajos and Mamykina, 2022; Kaur et al., 2020; Liao and Varshney, 2021; Ma et al., 2022a). One of the key challenges in AI-assisted decision-making is achieving complementary performance, where the collaborative decision outcome surpasses what either human or AI can achieve alone (Bansal et al., 2021; Lai and Tan, 2019; Zhang et al., 2020; Schemmer et al., 2023). However, empirical studies reveal difficulties in achieving this, primarily due to two issues.
One challenge is the underutilization of human and AI domain knowledge (Bansal et al., 2020). Some researchers aim to leverage the complementary aspects of human and AI intelligence by training AI to complement human knowledge (Bansal et al., 2019b; Wilder et al., 2020). Moreover, existing AI-assistant interfaces do not efficiently harness the knowledge of both parties (Steyvers and Kumar, 2022). AI contributes its knowledge to humans by providing recommendations with AI explanations serving as a means of representing its detailed reasoning (Lai et al., 2021). These explanations could facilitate the collaborative synthesis of human and AI intelligence, allowing them to combine insights into different features for final decisions. However, when conflicting views arise, current interfaces provide limited support for the communication and exchange of human and AI knowledge (Liao and Varshney, 2021). To address this, we propose Human-AI Deliberation to resolve conflicts through natural discussions.
The second challenge concerns human reliance on AI suggestions (Bansal et al., 2021; Zhang et al., 2020; Ma et al., 2023, 2024b, 2024a). Achieving complementary performance relies on human decision-makers’ ability to judiciously determine when to consider AI recommendations and when to be skeptical (Zhang et al., 2020; Buçinca et al., 2021; Rastogi et al., 2022). Both over-reliance, where individuals trust AI excessively (Lee and See, 2004; Parasuraman and Riley, 1997), and under-trust, where individuals fail to utilize AI when necessary (Lee and See, 2004), can lead to adverse outcomes. Successful decision-making requires individuals to decide whether and how to rely on AI recommendations on a case-by-case basis (Zhang et al., 2020; Bansal et al., 2019a, b, 2021; Turner et al., 2022). Current approaches present AI performance indicators, explanations, outputs, and confidence levels to assist humans in making informed decisions. However, existing research has found that when people are provided with a recommendation and passively look at it, they often lack analytical thinking, leading to over-reliance or under-reliance on AI systems (Buçinca et al., 2020; Gajos and Mamykina, 2022; Kaur et al., 2020; Liao and Varshney, 2021). Human-AI Deliberation, as proposed in this paper, encourages a careful evaluation of AI rationales through discussions of conflicts in human and AI opinions. By engaging humans in the deliberation process, it promotes a more comprehensive understanding of AI insights, reducing the potential for both under-reliance and over-reliance.
2.2. The Role of Deliberation in Human Decision Making
The meaning of deliberation is “the act of thinking about or discussing something and deciding carefully” (Merriam-Webster, [n.d.]). It involves considering all relevant individuals as moral agents who must justify their viewpoints and listen to others’ reasons (Gracia, 2003). Rather than seeking consensus, the process aims to enhance individual perspectives by incorporating others’ viewpoints, thus increasing decision maturity and wisdom (Gracia, 2003). The origin of group deliberation can be traced back to public deliberation or deliberative democracy, where citizens convene to discuss policies with potential implications for their lives (Simon, 2008). Recent studies on online deliberation have showcased its ability to enhance the accuracy of crowd-working tasks (Chen et al., 2019; Drapeau et al., 2016), improve perceptions of procedural justice (Fan and Zhang, 2020), and facilitate consensus-building among participants (Lee et al., 2020; Schaekermann et al., 2019; Van Berkel et al., 2019; Xie et al., 2020). Furthermore, deliberation proves effective in facilitating various decision-making tasks, including clinical diagnosis (Schaekermann et al., 2019; Inguaggiato et al., 2019; Preisz, 2019), criminal justice (Devine et al., 2001; Van Berkel et al., 2019), and more.
Effective decision-making is of paramount importance across diverse domains, and deliberation emerges as a valuable process replete with manifold advantages. First, it enhances decision quality and problem-solving (Barabas, 2004; Landemore and Page, 2015). Deliberation involves comprehensive analysis and careful evaluation of options, fostering a holistic understanding of issues by considering various factors and potential outcomes (Jacobs et al., 2009). Research underscores that deliberative decisions tend to be wiser and more effective owing to their foundation in thorough information and thoughtful analysis (Landemore, 2012; Dijksterhuis et al., 2006). Second, deliberation promotes participation and collaboration (Nabatchi and Leighninger, 2015). It encourages engagement from multiple stakeholders, thereby facilitating cooperation and communication among decision-makers (Wojcieszak et al., 2010). This collaborative process aids in resolving complex issues, providing decision-makers with ample time for dialogue and negotiations, ultimately ensuring the widespread acceptance of decisions (Fishkin, 2018). Third, deliberation has the potential to mitigate decision biases and enhance fairness (Kramer et al., 1990; Hochman et al., 2015). It assists decision-makers in objectively analyzing issues, mitigating personal biases and emotional influences, and identifying flawed perceptions and knowledge (Tetlock, 2017).
Despite the significance of deliberation in decision-making, there is a dearth of research on its integration into AI-assisted decision-making processes. To address this gap, drawing upon theories and practices in deliberation (Black et al., 2010; Habermas, 2005; Thompson, 2008; Lord and Tamvaki, 2013; Steiner et al., 2005), we propose a framework called Human-AI Deliberation to facilitate human reflection and discussion on conflicting human-AI opinions. Based on this framework, we aim to move a first step towards designing a Deliberative AI and investigating its effects on decision processes and outcomes through an exploratory empirical study.
2.3. Existing Studies on Deliberation in AI-Assisted Decision Making
Deliberation encompasses two essential facets: the analytical and contemplative consideration of issues and active discussion (Gracia, 2003). These aspects have garnered some attention in previous research on AI-assisted decision-making.
Concerning the stimulation of analytical thinking in humans, one approach involves interventions aimed at encouraging deeper engagement in System 2 thinking (Kahneman, 2011) by controlling when AI information is presented. For instance, the concept of ”cognitive forcing” has explored strategies to compel human decision-makers to spend more time deliberating on problems (Buçinca et al., 2021; Park et al., 2019; Rastogi et al., 2022). Methods include asking individuals to make independent predictions before receiving AI suggestions (Ma et al., 2023; Buçinca et al., 2021) or implementing a ”slow algorithm” (Park et al., 2019; Buçinca et al., 2021). These cognitive forcing techniques have demonstrated the ability to reduce human reliance on AI. Additionally, researchers have investigated how and what information should be presented. For instance, Garhos et al. (Gajos and Mamykina, 2022) found that people engaged in more analytical thinking when AI explanations were provided without concrete recommendations. Furthermore, Danry et al. (Danry et al., 2023) introduced AI-framed Questioning to enhance critical thinking and human discernment of flawed statements. While these approaches yield positive results, they come with limitations, such as potentially leading to under-reliance and not addressing differences between human and AI perspectives.
Regarding discussion, some research has explored dialogues between humans and AI in cooperative games ((FAIR)† et al., 2022; Kramár et al., 2022), but these discussions have not been specifically tailored to decision-making tasks. Recent efforts have started to integrate discussions into AI-assisted decision-making, such as Zheng et al.’s (Zheng et al., 2023) inclusion of AI in group decision-making to assess student essays. However, these efforts often rely on Wizard of Oz (WoZ) setups and lack genuinely designed AI systems for human interaction. Chiang et al. (Chiang et al., 2023) examined human-AI collaborative decision-making for recidivism risk assessment, but in their work, AI only offers suggestions and does not participate in the discussion with humans. Additionally, Zhang et al. (Zhang et al., 2023) utilize AI models as boundary objects to foster deliberation among organizations, but they do not encompass deliberation between humans and AI. Perhaps the most similar work to ours is that of Slack et al. (Slack et al., 2022). They focused on AI explanation and proposed using dialogue to improve the flexibility to support arbitrary follow-up questions that users might have and enhance people’s understanding of AI. In this paper, we also use the form of dialogue to allow people to interact with the AI, and also include the part where the AI gives explanations through dialogue (such as the AI answers users’ questions). But the key difference is that we focus on deliberation design and propose Deliberative AI which can not only “explain to users” but also actively engage users in the deliberative discussions by “asking or challenging” the users, aiming to promote people’s critical thinking.
Drawing inspiration from the principles of deliberation, we introduce a novel paradigm of Human-AI Deliberation. Additionally, we design Deliberative AI which combines the conversational capabilities of large language models (LLMs) with the predictive abilities of domain-specific models (DS-Model). Deliberative AI encourages meticulous human thinking and facilitates discussions concerning disparities between human and AI viewpoints. Through an exploratory assessment, we investigate the potential impacts of deliberation on AI-assisted decision-making. Our work takes an initial step toward designing and evaluating human-AI deliberation in decision-making contexts, offering valuable insights into the integration of deliberation into AI-assisted decision-making processes.
3. Framework: Human-AI Deliberation
In this section, we first introduce the general process of decision-making with an emphasis on the step that weighs the evidence for an alternative. Then, grounded in the theoretical construct of Weight of Evidence (WoE), we propose the Human-AI Deliberation framework. Finally, based on theories and practices in human deliberation, we introduce the design considerations and structure of Deliberative AI, the core AI component of our framework.
3.1. Unpacking Decision-Making
Decision-making is a process of making a choice from a number of alternatives to achieve a desired result (Eisenfuhr, 2011). Existing literature summarizes the general decision-making process into seven steps (U Mass, [n.d.]; Lunenburg, 2010) – Step 1: Identify the problem, Step 2: Collect information, Step 3: Identify the alternatives, Step 4: Weigh the evidence, Step 5: Choose from the alternatives, Step 6: Implement action, and Step 7: Evaluate the results.
In this work, we dissect a decision-making problem and human-AI thoughts into four structured components (Figure 2 (a)). (1) Decision means the overall decision to make for a specific problem. (2) Dimension refers to an aspect that people consider when reaching an overall decision. For instance, in a graduate admission task, dimensions could be applicants’ academic excellence, research ability, etc. Taking tabular task data (commonly used in decision-making (Ghai et al., 2021; Wang and Yin, 2021)) as an example, a dimension may be an attribute or a set of related attributes. (3) Opinion on a dimension denotes one’s views on the impact of a dimension on the overall decision (e.g., an applicant’s academic excellence contributes +50% to the overall chance of being admitted). (4) Evidence is the basis of an opinion. For humans, evidence can encompass facts, heuristics, past experiences, personal beliefs, interpretations, and even creative thoughts (Simon, 1990, 1997). In contrast, AI’s evidence primarily originates from information and knowledge embedded in the training data.
3.1.1. When and Where to Deliberate?
In the scope of this paper, we propose to conduct Human-AI Deliberation in “Weigh the Evidence” (Step 4 in the seven decision-making steps). For one thing, disagreements between human’s thoughts and AI’s suggestions often surface when they assess alternatives in this step. Diverse interpretations of evidence in Step 4 are likely to lead to conflicting opinions and disparate outcomes in steps 5-7 (Lai et al., 2021). For another, in existing AI-assisted decision-making tasks (Lai et al., 2021; Lai and Tan, 2019; Bansal et al., 2021), steps 1-3 are usually settled in advance and there is no need for much deliberation. We propose to have humans and AI deliberate their opinions at the detailed Dimension level (i.e., to what extent a dimension supports or opposes the final decision) rather than merely talking about accepting/rejecting one’s overall decision. In this process, both human and AI need to support their arguments with evidence and weigh the presented evidence according to credibility and probative value (Baumann et al., 2014).
3.1.2. What to Deliberate on?
In real-world practices, a deliberation process in Step 4 is often informed by the importance of evidence interpreted in a particular context (Cartwright and Stegenga, 2011). In information theory, such “importance” is quantified as the (probabilistic) weight of evidence (WoE) for a given hypothesis (and against alternative hypotheses) (Peirce, 2014; Good, 1950; Gough, 2007). WoE is well studied and widely adopted in various decision-making domains (Alvarez-Melis et al., 2021). This is because its literal meaning is easy to understand by lay users and its probabilistic calculation is simple to implement in practical applications (Saaty, 2008; Gough, 2007).
WoE is the core content of deliberation in our proposed framework. It corresponds to “opinion” – the importance of a dimension to a final decision given one’s interpretation of evidence. In the remainder of this paper, we use the term opinion (as well as its synonyms e.g., viewpoint) and WoE interchangeably.
It is crucial to note that during evidence weighing, both humans and AI possess strengths and limitations. For instance, humans may offer novel insights not present in AI training data, yet they can be constrained by time and cognitive biases, as per the “Bounded Rationality Model” (Simon, 1990, 1997). AI excels in data-driven knowledge and computation, but is known to be susceptible to biases in training data and falters in the face of out-of-distribution cases (Liu et al., 2021). These inherent human and AI shortcomings can result in potentially incorrect, incomplete, or biased evidence weighting, leading to opinion discrepancy and even decision errors. Hence, fostering humans’ deliberation on conflicting evidence weights could be advantageous.
Figure 2. The framework for Human-AI Deliberation. (A) Illustrates the Weight of Evidence (WoE) concept in decision-making, showcasing how decision-makers assess evidence across dimensions to shape opinions and arrive at a final decision. (B) Presents the Architecture for Human-AI Deliberation, with key activities (shown in grey boxes) and potential design space (shown in dashed-line boxes).
[…]
3.2. Human-AI Deliberation: A Framework
Deliberation is beneficial for enhancing decision-making by integrating diverse perspectives, improving the quality of solutions by considering various viewpoints, fostering critical thinking, and increasing the likelihood of reaching informed and well-thought-out conclusions (Gracia, 2003). Although deliberation is commonly used in democracy and political settings (Black et al., 2010; Habermas, 2005; Thompson, 2008), its idea can be applied to other decision-making tasks such as clinical diagnosis (Schaekermann et al., 2019; Inguaggiato et al., 2019; Preisz, 2019), criminal justice (Devine et al., 2001; Van Berkel et al., 2019), etc. Building upon the WoE decision-making approach, we propose Human-AI Deliberation, a framework to stimulate deliberative processes involving both humans and AI (Figure 2). This framework comprises the following essential activities:
  • Elicitation of Thoughts: Human and AI start with articulating their dimension-level perspectives on the decision problem. While AI presenting its “thoughts” (e.g., in the form of feature importance explanation) is rather common in AI-assisted decision-making (Liao and Varshney, 2021), this activity also encourages individuals to clarify their ideas and examine their reasoning, which prompts analytical thinking in human (Milkman et al., 2009; Buçinca et al., 2021). Two aspects of this activity require careful design. First, AI thought elicitation demands a good balance between human information needs and interpretability (Mishra and Rzeszotarski, 2021; Abdul et al., 2020). Prior research suggested that certain mechanisms AI applies to make inferences and predictions, although informative, are not comprehensible to humans (Miller, 2019). Hence, the design of AI’s “thought” explanations should be critical for this activity. Second, human thought elicitation, while encouraging thoughtful reasoning (Milkman et al., 2009; Buçinca et al., 2021), can impose a potential workload. It thus demands suitable, friendly interface designs.
  • Alignment of Human-AI Thoughts: As human’s and AI’s viewpoints and the process they form those viewpoints may diverge (Kaufman and Kirsh, 2022; Miller, 2019; Holzinger et al., 2019), this activity is tasked with establishing a common language for the two parties to compare their WoE and determine the extent of discrepancy, if any. Proper assessment and presentation of human-AI WoE differences can help effectively navigate humans’ attention and efforts in the subsequent activities (Boggust et al., 2022).
  • Discussion: This activity features constructive discussions that enable humans and AI to substantiate their opinions via communication, clarifying choices of evidence, and the basis for assigned weights. It has the potential benefits of cultivating critical thinking, mitigating biases, and raising awareness of differences between multiple parties (Milkman et al., 2009; Larson et al., 1994; Hirokawa, 1985). This activity has a huge design space and usually needs to be tailored to specific decision tasks. It entails considering various discussion-related factors, including content, style, leadership (who initiates and leads the conversation), duration, and so on. A typical solution is transferring human-human discussion practices (Milkman et al., 2009; Larson et al., 1994; Hirokawa, 1985) to the human-AI discussion contexts.
  • Update of Thoughts: In-depth discussions may expose potential flaws and conflicts in the original decisions as humans and AI are both imperfect (Bansal et al., 2021). This activity provides an opportunity for them to reflect on the gaps in thinking (Milkman et al., 2009) and revise their thoughts accordingly. For AI, this means designing appropriate mechanisms to interactively update its recommendations. For humans, the interface should possess the flexibility for them to change their WoE.
In summary, the proposed Human-AI Deliberation framework consists of four interlinked activities and requires proper designs (the potential design space is illustrated in the dashed box in Figure 2 (b)) of a decision-making interface and an AI with the ability to deliberate with humans. Since the interface design is task-dependent, we only introduce the design of Deliberative AI in the next subsection. Notably, not all activities in this framework are necessarily required for Human-AI Deliberation. Furthermore, we do not intend to exhaust all possible forms of Human-AI Deliberation. Researchers are encouraged to extend this framework or propose plausible alternatives.
3.3. Deliberative AI: Design Considerations and Overall Structure
Inspired by human-human deliberation practices and measurements, we derive a set of design considerations for a Deliberative AI. Then, we introduce its overall architecture proposed in accordance with these considerations.
According to deliberative theories (Black et al., 2010; Habermas, 2005; Thompson, 2008) and practices (Lord and Tamvaki, 2013; Steiner et al., 2005), Discourse Quality Index (DQI) (Steenbergen et al., 2003) and its improved versions (Bächtiger et al., 2022, 2009) are the most comprehensive and widely used guideline for assessing human deliberation. We adapt DQI to our AI-assisted decision-making context and summarize the following design considerations (DCs):
    DC 1. Participation equality: Deliberative AI should ensure that both parties possess equal voice (Chambers, 2005) and share similar opportunities to offer opinions and reasons as well as to participate in discussions.
    DC 2. Justification rationality: Deliberative AI should adeptly provide rational justifications for its stances during interactions and encourage humans to do the same.
    DC 3. Constructive updates: Rather than rigidly adhering to its initial opinions or blindly leaning towards others, Deliberative AI should aim to facilitate compromise, reconciliation, or consensus as deliberation evolves. It should help both sides to think carefully and rationally and update their WoE in a timely manner.
    DC 4. Interactivity: Deliberative AI should be able to understand human intentions and dynamically generate appropriate responses based on human’s questions, arguments, and statements.
    DC 5. Respect and agreement: Deliberative AI must ensure polite discourse and respect for other participants, especially during discussions. Even if it disagrees with humans on some aspects, Deliberative AI should show respect and understanding, creating a positive environment for continued engagement.
Figure 3. The Architecture of Deliberative AI. Our design integrates both a domain-specific model and a Large Language Model, enabling the AI to engage in natural communication with humans while also harnessing domain knowledge derived from the specialized model.
[…]
To fulfill these considerations, we integrate Large Language Models (LLMs) and domain-specific models (DS models) to build Deliberative AI. DS models are responsible for the initial generation and subsequent refinement of AI’s WoE. DS models’ predictive power and domain knowledge offer reliable (instead of potential hallucination) information for deliberation activities. LLMs, on the other hand, bridge the interactions between humans and DS models with their conversation abilities. Overall, the architecture of Deliberative AI (illustrated in Figure 3) comprises three layers: Communication layer, Control layer, and Knowledge layer.
  • The Communication layer, empowered by LLMs, incorporates three components:
      Intention Analyzer (for DC 4) understands human intent and argument evidence, facilitating cross-referencing with the Knowledge layer through the Control layer.
      Deliberation Facilitator (for DC 2&5) encourages careful thinking and rational justifications while maintaining respectful deliberation.
      Argument Evaluator (for DC 2&3) assesses human justification rationality, which can be used to further prompt humans’ reasoning and update AI opinions.
  • The Control layer, encompassing four components, oversees:
      Dialogue Controller (for DC 1) manages the deliberative discussion process (e.g., when to elicit thoughts, when to update opinions, when to move on to the next dimension, etc.)
      Regulator (for DC 2) guides and constrains LLM output with domain-specific model insights and training data.
      Knowledge Extractor (for DC 2) extracts data insights from domain-specific models and training data based on human intent analysis.
      Opinion Update Controller (for DC 3) adjusts AI viewpoints based on human-AI dynamics (e.g., the strength of justifications, uncertainty behind AI’s opinions, etc.).
  • The Knowledge layer comprises a domain-specific model and training data, providing both domain-specific knowledge and data-derived insights.
In Sec. 4, we will empirically explore a subset of these dimensions with a specific instantiation of the proposed Human-AI Deliberation framework and Deliberative AI in a graduate admission task.
4. Instantiating the Framework: Graduate Admission Prediction
4.1. Task, Dataset and AI Model
We choose to use graduate admission as an illustrative task to demonstrate how to instantiate the proposed Human-AI Deliberation framework. In this task, participants decide on admitting or rejecting applicants to a U.S. university based on their profiles. We chose this task for two key reasons. First, this task is widely used in AI-assisted decision-making research (Cheng et al., 2019; Zhang et al., 2023; Echterhoff et al., 2022; Zhao et al., 2023), with real-world universities employing AI algorithms for decision consistency and workload reduction (Waters and Miikkulainen, 2014; Pangburn, 2019). Second, graduate admission often involves deliberation among committee members (Zhang et al., 2023), making it ideal for studying the effects of our proposed Human-AI Deliberation.
The task utilizes a synthesized dataset (Cheng et al., 2019) that simulates profiles of applicants at a U.S. public university based on publicly available aggregate statistics and distributions [^1] Due to privacy issues, public graduate admission data sets are all synthesized. We acknowledged that it may deviate from the real-world setting.. The dataset comprises 100 student applications’ profiles, featuring attributes considered by admission committees in actual scenarios, e.g., GRE Verbal, GRE Quant, GRE Writing, GPA, Statement of Purpose Strength, Diversity Statement Strength, Country, Major, Applicant’s Undergraduate Institution Rank, and Recommendation Letter Strength. The dataset also includes a decision label for each case: strong reject, weak reject, weak accept, or strong accept.
To build a domain-specific model (DS-model) that can generate suggestions, we trained a multi-category linear model using a 70% random split of the dataset as in (Cheng et al., 2019). We employed a linear regression model as a decision classifier, discretizing the predicted responses into one of the four decision labels. Consistent with common practices (Echterhoff et al., 2022), we further binarized the original labels, mapping strong/weak reject to “reject” and strong/weak accept to “accept” as the ground truth for assessing AI model’s and participants’ prediction accuracy. The trained model achieved an 80% accuracy on the remaining 30% test set. The task samples used in the study were selected from the test set. It should be noted that in order to explore the effect of Human-AI Deliberation, we specially selected cases from the test data that are challenging for both humans and AI, so that the performance of AI is around 50%. But this does not mean that the AI model used in our study is an unrealistically low-performance AI, just that we focus on difficult task cases.
4.2. Implementation of Human-AI Deliberation
4.2.1. Human-AI Thought Representation and Alignment
In our Human-AI Deliberation framework, the initial step involves both humans and AI externalizing their thoughts. We employ feature contribution (Liao and Varshney, 2021; Liao et al., 2020) to represent their weight of evidence (WoE) along each decision dimension (Alvarez-Melis et al., 2021). Feature contribution is represented by contribution scores indicating the positive or negative influence of each feature $x_{i}$ on the final prediction $y$. In graduate admission tasks, we treat each attribute in an applicant’s profile as a dimension, and feature contribution requires humans and AI to assess the influence of each dimension on the final decision. It provides a common ground for AI and humans to express, compare, and initiate discussions about their thoughts at the feature level (see Table 1).
Table 1. The mapping between AI “thoughts” and human thoughts in the Weight of Evidence framework.
Component AI “thoughts” Human thoughts
Decision Accept or reject one applicant Accept or reject one applicant
Dimension Feature/Attribute (e.g., GPA) Feature/Attribute (e.g., GPA)
Opinion A feature’s contribution (e.g., GPA: + 13%) Weight of evidence (e.g., GPA: -5%)
Evidence Training data Data or other knowledge held by people
For the human side, WoE can be interpreted as the influence of an attribute on the total chance of an applicant being admitted. For the AI side, we used SHAP (SHapley Additive exPlanations) (Lundberg and Lee, 2017), a widely-used explainable AI algorithm, to generate AI feature contributions. SHAP values indicate both the direction and strength of a feature’s impact on predictions. SHAP offers two key advantages: it captures feature interactions, mirroring human decision-making, and its additive nature resembles how humans combine evidence weights for or against options (Baumann et al., 2014). However, when applied in the context of Human-AI Deliberation, two limitations arise: (1) Raw SHAP values can be challenging to interpret directly. To better convey feature influence to non-expert humans, we converted SHAP values into probabilities by adopting a regression model and transforming the four-category label range into a 0-100% range (in this way, the generated SHAP value can be directly mapped to 0-100%)  [^2] Other conversion methods include using the TreeExplainer from the SHAP library (and setting model_output=“probability”) to generate the probabilistic SHAP value. Because the admission prediction used in this paper can be naturally converted into a probability problem, we adopted more direct category label-probability mapping. (2) SHAP values only tell users how important a feature is, but not why it is important. To bridge this gap, we propose generating “meta-explanations” that use Knowledge Extractor (Sec. 3.3) to extract evidence (e.g., data patterns) from training data for each dimension, enhancing AI’s transparency during deliberation.
4.2.2. Implementation of Deliberative AI
Next, we describe how we implemented each component of Deliberative AI.
I. Communication Layer. This layer serves as a vital bridge between humans and the DS-model, facilitating effective communication by comprehending human inputs and crafting relevant responses.
I-1. Intention Analyzer. We harnessed the language capabilities of LLMs [^3] In this study, we use GPT-3.5-turbo as the language model, which has sufficient conversational power and is freely accessible (relative to GPT-4). We will use “LLM” throughout this paper for consistency. to discern human intentions and targeted dimensions in discussion. To formulate effective prompts for intention analysis, we conducted a pilot study to gather common dialogues around graduate admission decisions, including questions, arguments, critiques, and challenges. In the pilot study, we developed a preliminary version Deliberative AI (with imperfect deliberative discussion capability) to carry out conversations with 30 participants from Prolific444www.prolific.co. Each participant expressed their opinions on various dimensions of applicant profiles and engaged in discussions with the AI, particularly focusing on conflicting opinions. This yielded 226 human deliberative statements. To extract diverse intentions from these statements, two authors conducted qualitative coding using thematic analysis (Hsieh and Shannon, 2005), and the results are summarized in Table 2. We then iteratively refined LLM prompts based on the collected data and built an ”Intention Analyzer” with a 96% accuracy in identifying themes of participant statements. Specific prompts are available in the supplementary materials.
Table 2. Qualitative analysis of the sentiment/intention category of participants’ statements (arguments, justifications, questions, critiques, etc.) in the deliberative discussion.
Themes Definitions and Examples #Participants
Distribution/Level of an attribute’s values Participants evaluate how attribute values are distributed among the pool of applicants. 35 (15%)
3.16 isn’t a bad GPA - it’s only slightly below average, sure, but it’s still fairly good” (P2)
Overall importance of an attribute Participants consider or challenge the overall importance of an attribute on the admission decision. 24 (10%)
Diversity is extremely important to the institution as a whole so the students highly rated diversity statement would highly influence their admittance.” (P33)
Contribution of an attribute Participants directly express their opinion on an attribute’s contribution or challenge the contribution given by the AI but without evidence. 47 (20%)
I know Applicant Undergraduate School Ranking has a significant impact on the chance of admission. But why is medium rank not good?” (P1)
Contrastive evaluation Participants compare an attribute’s current value with other values (often using the average) to judge an attribute’s impact. 41 (18%)
I am surprised you ranked the applicant’s GPA on a negative scale. 3.26 is not that much lower than the 3.5 of the last applicant.” (P10)
Holistic review of multiple attributes Participants evaluate how different attributes interact, taking into account the influence of certain attribute values on the strength of others. 23 (10%)
The engineering major is incredibly difficult and any GPA above a 3.5 is considered successful.” (P3)
I said 2% positive influence because this individual went to a top rank school, which I assume is harder academically than some lower ranked schools.” (P22)
Data-irrelevant questions/arguments Participants give data-irrelevant statements based on their heuristics, past experiences, personal beliefs, etc. 77 (34%)
Statement of purpose is the only part of the application process where the applicant gets to show us who they really are in their own words - not just a score or some data value. I ranked these higher for this reason.” (P5)
I-2. Deliberation Facilitator. This component addresses DC2 (Justification Rationality) and DC5 (Respect and Agreement), as discussed in Sec 3.3, by designing corresponding LLM prompts. In particular, we instruct LLM to (1) Demonstrate a nuanced understanding of the human’s statement; (2) analyze the specific content of the person’s statement; and (3) provide a thoughtful and critical response. For detailed prompts, please refer to the supplementary materials.
I-3. Argument Evaluator. The main function of this component is to assess the strength of a person’s statement, which informs updates to AI opinions. Drawing from established theories in human argumentation evaluation (Harris, 1997; Van Eemeren et al., 2002; Van Eemeren and Henkemans, 2016), we devised a comprehensive scoring mechanism with nine key items: Clarity, Relevance, Evidence, Logic, Consistency, Counterarguments, Depth, Credibility, and Alignment. These criteria are integrated into a prompt, guiding the LLM to evaluate human statements. We then average and scale the scores to obtain the overall human argument strength $S_{Human}$ (from 0 to 1; 0: weakest, 1: strongest). Additional details, including scoring schemas and prompts, can be found in the supplementary materials.
In summary, the communication layer can engage in general interactions with humans. To imbue it with specific model opinions and knowledge, we require a control layer to mediate between the LLM and the DS-model.
II. Control Layer. This layer manages the querying and extraction of specific DS-model opinions and knowledge while controlling the entire conversation flow.
II-1. Dialogue/Discussion Controller. This component serves as the control center for the discussion process, orchestrating a structured deliberation flow as shown in Figure 4. It unfolds as follows: [Thought Elicitation] Participants express their WoE on each dimension; AI responds with its perspectives. [Discussion] AI highlights commonalities and discrepancies, inviting participants to provide justifications or question differing viewpoints. AI responds with critical insights. All three components of the Communication layer (Intention Analyzer, Deliberation Facilitator, and Argument Evaluator) play vital roles in this phase. After one round of discussion, AI offers input options for participants to update, maintain, or continue the discussion. AI proceeds based on participants’ choices. If they wish to move to the next dimension, AI summarizes any pending dimensions, highlighting differences. Participants can choose to explore untouched dimensions, revisit previous discussions, or skip this round. Participants have the flexibility to initiate dialogues on any dimension at any time, using quick input options or free text. They can refine their views on the decision interface independently of AI opinion updates.
II-2. Knowledge extractor. Based on the attributes/dimensions and intent types identified by the “Intention Analyzer” (see Table 2), we developed a series of query functions to extract relevant data knowledge from the DS-Model. These functions help pull evidence for the LLM to generate responses in deliberative discussions appropriately. We established a mapping between the recognized intent type and the query function and called different query functions based on the recognized intent type. Below is a brief overview of the designed functions corresponding to different human intent types. Please refer to the supplementary materials for detailed codes and examples.
  • Distribution/Level of an attribute’s value: Function get_distribution (attribute) calculates attribute value percentiles within the applicant pool, along with contextual comparisons (with minimum, maximum, quartiles, mean, and median).
  • Overall attribute importance: Function get_global_feature_importance (attribute) returns global importance. Function get_correlation (attribute) provides Pearson correlation. Function get_influence_on_admission_chance (attribute) calculates admission chance changes for varying attribute values.
  • Contribution of an attribute: Function get_current_value_influence (attribute) calculates admission chance differences when an attribute is randomized.
  • Contrastive Evaluation: Function get_contrastive_admission_chance (attribute, contrastive) computes admission chance differences with a contrastive value.
  • Holistic review of multiple attributes: Function get_holistic_analysis (attributes, fixed_attributes) evaluates attribute impacts in specific scenarios, e.g., GPA percentile in top-ranked schools.
II-3. Regulator. The primary objective of this component is to harness the expertise of the DS-Model to regulate the responses generated by LLM. This approach makes certain that LLM’s responses always align with the DS-Model’s knowledge and decisions. To achieve this goal, we created consistency-ensuring prompts based on three key elements: (1) the findings extracted by the Knowledge Extractor, (2) the overarching decisions made by the DS-Model, and (3) the DS-Model’s viewpoint on the current attribute under discussion.
Figure 4. The conversation flow for the deliberative discussion.
[…]
II-4. Opinion Update Controller: We updated the AI’s opinions by taking into consideration: (1) the current opinions of both the human ($O_{Human}$) and the AI ($O_{AI}$) on the discussed attribute, (2) the strength of the human’s argument ($S_{Human}$, see Argument Evaluator), and (3) the AI’s uncertainty ($U_{AI}$) measured and calibrated via Uncertainty Quantification 360 toolbox (Ghosh et al., 2021) (the uncertainty ranges from 0 to 1: the closer to 1, the more uncertain AI’s prediction is). We propose the following formula to update the AI’s opinions ($\hat{O}_{AI}$) on an attribute based on these factors, inspired by result aggregation in crowd intelligence (Lyon and Pacuit, 2013; Galton, 1907):
(1) $\hat{O}_{AI}=\frac{1-U_{AI}}{1-U_{AI}+S_{Human}}\cdot O_{AI}+\frac{S_{Human}}{1-U_{AI}+S_{Human}}\cdot O_{Human},$
III. Knowledge Layer
This layer primarily consists of the DS-Model and the training dataset. The DS-Model offers comprehensive perspectives and facilitates real-time decision predictions. The training dataset offers necessary information (e.g., data distributions and patterns) for the Knowledge Extractor to perform real-time calculations and queries.
4.2.3. An Example
Figure 5 provides the details of a conversation between a human and a Deliberative AI discussing how an applicant’s GPA affects admissions chances. Here’s a step-by-step breakdown:
  1. (1)
    The user inputs GPA-related arguments in the dialogue interface.
  2. (2)
    The system packages the user’s input as a prompt for the Intention Analyzer in the Communication Layer.
  3. (3)
    The Intention Analyzer recognizes attributes and intentions and saves in JSON format, then forwards it to the Knowledge Extractor in the Control Layer.
  4. (4)
    The Knowledge Extractor generates a query function and fetches statistical results from the DS-Model and training data.
  5. (5)
    The statistical results are transmitted to the Regulator.
  6. (6)
    Regulator crafts a constraint prompt ensuring consistency between the LLM’s output and the DS-Model’s prediction, feeding it to the LLM-based Deliberation Facilitator.
  7. (7)
    The Deliberation Facilitator generates responses to the user’s initial arguments.
Overall, in this framework, LLM is used for language understanding and generation. The opinions and evidence used by LLM are retrieved in real time from the DS-Model and training data through our logic code (like retrieval augmented generation (Lewis et al., 2020)). In this way, the LLM is used in a controllable and responsible manner, minimizing the potential hallucination.
Figure 5. An illustration of how Deliberative AI processes humans’ inputs and how it generates outputs. The prompts used are simplified in this figure for illustration purposes. (for the complete prompts, please check our supplementary materials)
[…]
4.3. Interface Design
The interface for the graduate admission task is structured into three main regions:
Figure 6. The interface of Deliberative AI. The interface contains three parts. The top part (A) is the applicant’s profile. The bottom left part (B) is the region for humans and AI to indicate (and update) their opinions. The bottom right part (C) is the discussion region where humans and AI can discuss conflicting opinions. (All the dashed lines are only for illustration)
[…]
  • Profile Region (Figure 6 (A)) displays the applicant’s profile, providing a table with the current value and possible range of each attribute. Users can access attribute definitions and basic data distribution statistics (minimum, maximum, average, and median values) by hovering over pink circular markers.
  • Opinion and Prediction Region (Figure 6 (B)) is dedicated to thought elicitation by both users and the AI.
      The upper part displays aggregate predictions from both humans and AI. This includes a legend (Figure 6-1) and two slide bars (Figure 6-2 and -3) representing AI’s and the user’s overall predictions, respectively. Each slide bar shows three line indicators: a white line representing the average admission probability of all applicants, a green line showing the initial predictions made by humans/AI, and a yellow line denoting the updated prediction by humans/AI (only shown after an update is made).
      The bottom part of this region allows both humans and AI to express specific opinions on each decision dimension (i.e., applicant attribute). A simplified profile is displayed in the middle (Figure 6-5) to minimize users’ need to switch attention. Each dimension is accompanied by a status indicator, denoting whether it has been discussed (green), is currently being discussed (orange), or is yet to be discussed (gray). Separate “concrete opinion” slide bars are presented next to each attribute (Figure 6-4 and Figure 6-6) for AI and humans to indicate their dimension-level opinions.
      Each dimension’s slide bar starts in a central position (0% contribution). Users can drag the slider any time to the right to increase the weight on an attribute toward a positive “admit” decision or to the left to reduce its contribution. Alternatively, users can directly input contribution values in a box below the slider.
    Slide bars in Figure 6-2 and Figure 6-4 are interconnected, so as those in Figure 6-3 and Figure 6-6. Values within the “overall prediction bar” reflect the cumulative values from the “concrete opinion bars.” Any changes in the dimension-level bars immediately update the overall prediction. AI’s and the user’s opinions are displayed side by side for easy comparison. Note that at the beginning of each case, users have to complete their opinion inputs and click the [Submit Opinion] button to see AI’s initial (overall and concrete) suggestions.
  • Discussion Region (Figure 6 (C)) is where all deliberative dialogues take place. Users can type out their opinion arguments, questions, disagreements with AI, responses to AI queries, and more. Importantly, changes made in the Opinion Region are seamlessly integrated by AI and reflected in ongoing discussions, and conversely, any viewpoint changes mentioned in the dialogue are instantly updated in the Opinion Region.
5. Exploratory User Study
To get an initial understanding of the impact of the Human-AI Deliberation mechanism on AI-assisted decision-making, we conducted a mixed-methods study in the context of graduate admission decisions. We call this study exploratory because we did not seek to specifically prove whether Human-AI Deliberation is effective for actual graduate admissions. In other words, graduate admission is NOT our focus. Instead, we just use graduate admission as an illustrative task to explore the potential effects of Human-AI Deliberation on AI-assisted decision-making.
5.1. Task Setup
We used the graduate admission task as our testbed. To ensure manageable study durations and prevent participant fatigue, we selected five task cases based on the length of the pilot study. These cases included one for the tutorial and four for the main tasks. Among the main tasks, two represented US applicants (US cases), and the other two represented Asian applicants (Asia cases). To keep the Deliberative AI discussions engaging and avoid repetition, we carefully chose tasks with diverse attributes. It’s worth noting that the AI model is biased toward US applicants due to training data biases. This bias results in a “+11%” feature contribution for “Country belongs to the US“ and a “-4%” feature contribution for “Country belongs to Asia”. So it can be interesting to test whether participants will discover such biases when collaborating with the AI.
5.2. Conditions
We compared the proposed Deliberative AI with the traditional explainable AI assistant (XAI) and Human Alone.
  • Deliberative AI (DAI): Participants share their thoughts on various dimensions before viewing AI recommendations. We present AI’s “thoughts” on each dimension afterward. After comparing conflicting viewpoints, we offer a dialogue interface for participants and AI to discuss any of the perspectives, as shown in Figure 6.
  • Explainable AI (XAI): After individuals provide their predictions, they receive AI recommendations (along with feature contribution-based explanations) and then make their final judgments (see Figure 15 in Appendix).
  • Human Alone: Participants need to make predictions independently without any AI assistance.
5.3. Procedure and Participants
5.3.1. Procedure
We conducted a between-subjects study. After obtaining consent, participants completed a background questionnaire gathering demographic data and assessing their AI expertise. Participants then went to an interactive tutorial, practiced with one example task, and received distribution and summary statistics for each attribute of the applicant’s profile. Following the tutorial, qualification questions were asked to check participants’ understanding of the task, with only those answering all questions correctly proceeding to the main task. The main task involved tackling four graduate admission task cases, including providing predictions/opinions, receiving AI suggestions/opinions, engaging in discussions with AI, and confirming final predictions. Finally, we collected participants’ perceptions, experiences, and feedback on the AI system and the discussion process in the exit survey.
5.3.2. Participants
Before participant recruitment, we conducted a power analysis to determine the required sample size for using G*Power (Faul et al., 2009). We specified a default effect size $f$=0.25 (indicating a moderate effect), a significance threshold $\alpha$=0.05, and a statistical power $1-\beta$=0.8. This resulted in a required total sample size of 159 participants for the three conditions. After obtaining institutional IRB approval, we recruited a total of 174 participants from Prolific4. To ensure high-quality responses, participants had to meet specific criteria: (1) residing in the United States; (2) having been admitted to a US graduate program before (as the task involved predicting graduate admission in a US university); (3) having at least a 99% approval rate with at least 1000 previous submissions; (4) using English as their first language; and (5) using a desktop computer for the experiment. After filtering based on attention-check questions, we obtained 153 valid responses (Deliberative AI: 48, XAI: 51, Human Alone: 54). Among the final participants, 84 self-reported as male, 67 as female, and 2 as other. There were 23 participants aged 24-29, 42 aged 30-39, 33 aged 40-49, 30 aged 50-59, and 25 aged over 59. Participants also rated their knowledge of artificial intelligence: 9 had no knowledge, 86 knew basic AI concepts, 50 had experience using AI algorithms, and 8 were AI experts. Participants in the Deliberative AI condition received bonuses based on the actual study length. To motivate high-quality work, participants received a \$0.50 bonus if their overall accuracy exceeded 75%. On average, participants earned about \$12 per hour.
Table 3. Measurements used in our user study. We collected participants’ objective decision and behavior data, subjective questionnaire data, and qualitative open-ended feedback.
Aspect Metrics Detailed Meaning and Questions
Objective Measures
Performance Decision Accuracy Accuracy of participants’ final predictions.
Reliance Agreement Fraction Percentage of tasks where participants’ final prediction agreed with AI’s prediction. $\rm\frac{Number\ of\ final\ decisions\ same\ as\ the\ AI\ suggestion}{Total\ number\ of\ decisions}$
Switch Fraction Percentage of tasks where AI’s prediction was used when initial disagreement existed. $\rm\frac{Number\ of\ decisions\ user\ switched\ to\ agree\ with\ the\ AI\ model}{Total\ number\ of\ decisions\ with\ initial\ disagreement}$
Over-reliance Ratio Fraction of tasks where participants used an incorrect AI prediction. $\rm\frac{Number\ of\ incorrect\ human\ final\ decisions\ with\ incorrect\ AI\ suggestions}{Total\ number\ of\ incorrect\ AI\ suggestions}$
Under-reliance Ratio Fraction of tasks where participants did not use a correct AI prediction. $\rm\frac{Number\ of\ incorrect\ human\ final\ decisions\ with\ correct\ AI\ suggestions}{Total\ number\ of\ correct\ AI\ suggestions}$
Behaviors Opinion Changes WOA = $\rm\frac{human\ finally\ predicted\ admission\ chance-human\ initially\ predicted\ admission\ chance}{AI\ initially\ predicted\ admission\ chance-human\ initially\ predicted\ admission\ chance}$
Frequency of Opinion Changes Count of opinion changes per prediction.
Dialogue Rounds Number of conversation rounds in a decision-making task.
Conflicting Dimensions Count of dimensions with conflicting information.
Dimensions Discussed Count of dimensions considered in a prediction.
Time Spent Average time spent per decision.
Subjective Measures
Perceptions of AI Helpfulness “I think the AI model’s assistance is helpful/useful for me to make good decisions.”
Trustworthiness “The AI model can be trusted to provide reliable decision support.”
Understanding “I understand how the AI model works to predict an applicant’s chance of being admitted.”
User Experience Decision Confidence “I feel confident in the decisions I made.”
Mental Demand “The decision-making process is mentally demanding.”
Effort “I have to work hard (mentally and physically) to accomplish my level of performance.”
Complexity “The decision-making process and the interaction with AI models are complex.”
Satisfaction “I am satisfied with the AI model’s assistance and the decision-making process.”
Open-ended Feedback Perception of helpfulness “Do you think the discussion with AI is (or not) helpful? Could you tell us the reasons why you think the discussion is helpful (or not helpful)?”
Perception of AI update “What do you think of the AI updating its own views during the discussion?”
Potential Improvement “To make a better discussion, which parts do you think the current AI needs to be improved, and how should it be improved?”
5.4. Measurement
To comprehensively assess the impact of human-AI deliberation, we evaluated the following aspects. Detailed meanings and questions for each metrics can be seen in Table 3.
  • Task Performance. We evaluated decision accuracy using Decision Accuracy (Zhang et al., 2020; Bansal et al., 2021; Ma et al., 2023).
  • Reliance. We assessed participants’ reliance on AI suggestions using Agreement Fraction (Zhang et al., 2020; Ma et al., 2023; He et al., 2023) and Switch Fraction (Zhang et al., 2020; He et al., 2023). Also, we measured the appropriateness of human reliance by Over-reliance Ratio (Wang and Yin, 2021; Visser et al., 2014; Parasuraman and Riley, 1997) and Under-reliance Ratio (Wang and Yin, 2021; Visser et al., 2014; Parasuraman and Riley, 1997).
  • Behaviors. We examined participants’ Opinion Changes (Bonaccio and Dalal, 2006; Poursabzi-Sangdeh et al., 2021; Gino and Moore, 2007), Frequency of Opinion Changes (Yaniv, 2004), Conflicting Dimensions, Dimensions Discussed, and Time Spent (Yang et al., 2020; Abdul et al., 2020). Specifically, for Opinion Changes, we use weight of advice (WOA) (Gino and Moore, 2007) to measure how humans’ overall opinions are influenced by AI’s opinions [^5] WOA is 0 when a human’s final prediction matches their initial one, 1 when it matches the AI’s prediction, and 0.5 when the individual averages their estimate with AI advice. A positive WOA means human opinions align more with AI (with values above 1 indicating over-adjustment), while a negative WOA signifies a divergence from AI opinions..
  • Perceptions of AI. We measured participants’ perceived Helpfulness (Laugwitz et al., 2008; Cai et al., 2019a; Buçinca et al., 2020), Trustworthiness (Ghai et al., 2021; Buçinca et al., 2021), and Understanding (Wang and Yin, 2021) through 7-point Likert scale (1: Strongly disagree; 7: Strongly agree. The same below).
  • User Experience. We measured participants’ Decision Confidence (Miller et al., 2015). Meanwhile, as deliberation requires extra efforts, we evaluated participants’ self-reported Mental Demand (Hart, 2006; Lai et al., 2022; Ghai et al., 2021; Buçinca et al., 2021), Effort (Hart, 2006), Complexity (Buçinca et al., 2021), and Satisfaction (Buçinca et al., 2021; Ghai et al., 2021) via 7-point Likert scale.
  • Open-ended Feedback. As we aim to gain an in-depth understanding of participants’ perceptions of the Deliberative AI and the deliberative decision-making process, we added three open-ended questions in the exit survey.
5.4.1. Analysis Methods
We conducted mixed-methods analyses to examine the aforementioned metrics. For the quantitative analysis, we first performed normality tests (Shapiro-Wilk) and found that the data did not fit the normality assumption. Therefore we ran the non-parameter tests. Specifically, to compare Deliberative AI and XAI (such as humans’ reliance on AI, and their perceptions of AI), we run Mann-Whitney U test. To compare all three conditions (such as task performance, and user experience), we employed Kruskal–Wallis tests with Bonferroni post-hoc correction. For qualitative analysis, two authors independently coded participants’ open-ended feedback and conversation logs using an inductive thematic analysis approach (Hsieh and Shannon, 2005). The final themes emerged through discussions and harmonization over several iterations. We also identified representative examples from the source texts for demonstration in this paper.
6. Results
In this section, we report our exploratory findings regarding the four research questions: (RQ1) how Human-AI Deliberation affects task performance and human reliance (and reliance appropriateness) on AI, (RQ2) how Human-AI Deliberation affects human perceptions and task experience, (RQ3) how humans engage in the deliberation process, and (RQ4) how humans perceive the effectiveness of Human-AI Deliberation and what should be improved.
6.1. RQ1: How will Human-AI Deliberation affect task performance and humans’ reliance (and its appropriateness) on AI suggestions?
6.1.1. Decision Accuracy
As shown in Figure 7, we find that participants in the Deliberative AI condition exhibited significantly higher decision accuracy ($M$=0.598, $SD$=0.169) compared to those in the XAI condition ($M$=0.524, $SD$=0.16, $p$¡0.05). Although there is no statistically significant difference, a trend emerges where Deliberative AI tends to surpass Human Alone performance, while XAI generally underperforms compared to Human Alone. This finding indicates that in scenarios where tasks are difficult for both humans and AI, traditional Explainable AI (XAI) might not enhance performance and could even lead to detrimental effects. Conversely, Human-AI Deliberation has the potential to provide positive benefits, even though the performance of underlying AI models in these challenging task cases is not good enough.
Figure 7. Task performance in different conditions. The error bars represent 95% confidence interval. (*: $p$¡0.05, **: $p$¡0.01, ***: $p$¡0.001)
[…]
6.1.2. Reliance
We measured participants’ objective reliance by agreement fraction and switch fraction. As shown in Figure 8 (a), participants agreed significantly less with AI’s suggestions in Deliberative AI ($M$=0.57, $SD$=0.24) than in XAI ($M$=0.68, $SD$=0.27, $p$¡0.05), and switched significantly less to AI’s predictions in Deliberative AI ($M$=0.23, $SD$=0.35) than in XAI ($M$=0.51, $SD$=0.41, $p$¡0.001). Combined with participants’ open-ended feedback (Sec. 6.4), this may be because people invest more in independent thinking in the process of deliberation with AI and realize the problematic aspects of AI’s perspective.
6.1.3. Appropriateness of Reliance
We further measured the appropriateness of participants’ reliance on AI’s suggestion by under-reliance and over-reliance (Figure 8 (b)). Results show that there is no significant difference between Deliberative AI ($M$=0.32, $SD$=0.28) and XAI ($M$=0.29, $SD$=0.30) in terms of under-reliance. While significantly less over-reliance was observed in Deliberative AI ($M$=0.47, $SD$=0.31) than in XAI ($M$=0.65, $SD$=0.33, $p$¡0.001), which means that participants had more appropriate reliance on AI when collaborating with our proposed Deliberative AI.
Figure 8. Participants’ reliance and the appropriateness of their reliance. (A) Participants’ reliance on AI’s suggestions was measured by agreement fraction and switch fraction. (B) The appropriateness of participants’ reliance on AI’s suggestions, including under-reliance (the ratio where participants did not use a correct AI suggestion) and over-reliance (the ratio where participants used an incorrect AI suggestion). The error bars represent 95% confidence interval. (*: $p$¡0.05, **: $p$¡0.01, ***: $p$¡0.001)
Figure 9. Learning curve measured by decision accuracy in each task round. The error bars represent 95% confidence interval.
[…]
6.1.4. An exploratory analysis of learning effects
In addition to the accuracy and reliance, we observed learning effects when humans worked with Deliberative AI to make decisions (Figure 9). To quantify the effect, following (Cabrera et al., 2023) we fitted a linear model of average participant accuracy for each decision round. We found a learning effect trend for Deliberative AI (although only marginally significant, $p$=0.051). However, no significant learning effects were found in XAI. This result reveals participants may gain better knowledge about the task and how to collaborate with the AI partner to deal with the task as the deliberation progresses.
In summary, Deliberative AI has a positive effect in improving decision accuracy. Besides, since the AI model’s accuracy is only 50%, participants in Deliberative AI relied less on AI’s incorrect suggestions which in turn leads to more appropriate reliance (less over-reliance). However, our current results do not mean that Deliberative AI will also lead to better decision performance and more appropriate human reliance than XAI when the AI performance is high. Future work is needed to further explore the effects when the AI performance varies.
6.2. RQ2: How will Human-AI Deliberation affect humans’ perceptions of the AI and user experience?
We measured the effects of different AI conditions on participants’ perceptions and user experience via a 7-point Likert scale (1: strongly disagree, 7: strongly agree).
Figure 10. Participants’ perceptions of the AI assistant. The error bars represent 95% confidence interval. (*: $p$¡0.05, **: $p$¡0.01, ***: $p$¡0.001)
[…]
Figure 11. Effects on user experience. The error bars represent 95% confidence interval. (*: $p$¡0.05, **: $p$¡0.01, ***: $p$¡0.001)
[…]
6.2.1. Perceptions of AI
Figure 10 shows participants’ perceptions of the AI model. There were no significant differences in perceived helpfulness and understanding between Deliberative AI and XAI. However, participants reported significantly less trust in Deliberative AI ($M$=4.47, $SD$=1.68) compared to XAI ($M$=5.52, $SD$=1.27, $p$¡0.01), aligning with their reliance behaviors (Sec. 6.1.2). This difference may be attributed to participants identifying more AI flaws through deliberation than by solely observing AI’s explanations, supported by conversation logs analysis (see Sec. 6.4 for details).
6.2.2. User experience
First, we want to see participants’ decision confidence. As indicated in Figure 11, participants in XAI reported significantly higher confidence ($M$=6, $SD$=1.10) in their predictions than those in Human Alone ($M$=5.59, $SD$=1.12, $p$¡0.05). However, from Figure 7 (a) we found that the final accuracy of participants in XAI is even lower than those in Human Alone (although not significant). This indicates that the traditional XAI might lead to humans’ illusionary confidence, which could prevent humans from making optimal decisions.
Given that the Deliberative AI requires participants to externalize thoughts at a dimension level and engage in deliberative discussions on conflicting opinions, it’s crucial to explore how these activities influence the user experience. Results showed no significant difference among the three conditions concerning Mental Demand, Effort, and Perceived System Complexity. However, we find participants reported significantly lower Satisfaction in Deliberative AI than in Human Alone. Combined with the qualitative results (Sec. 6.4), we speculate that this may be because Deliberative AI exposes more conflicts between humans and AI and they need to carry out deliberation around these conflicts. This can easily reduce people’s satisfaction compared to simply making a decision. We also observed a trend that Deliberative AI is more complex than the two baselines. These findings suggest that there is a trade-off between encouraging users’ deliberative thinking and optimizing their user experience, in line with findings in previous studies (Buçinca et al., 2021).
Figure 12. Descriptive statistics of participants’ decision-making process. (a) Descriptive statistics of conversations in Deliberative AI. (b) Time spent per decision in the three conditions.
Figure 13. The changes in participants’ opinions measured by Weight of Advice. (*: $p$¡0.05, **: $p$¡0.01, ***: $p$¡0.001)
[…]
6.3. RQ3: How will humans engage in the deliberation process when working with Deliberative AI?
To gain deeper insights into the impact of our Deliberative AI, we analyzed how participants engaged with this technology during the decision-making process.
As shown in Figure 12 (a), each decision-making instance typically involved a median of 2.0 rounds of conversation between participants and the Deliberative AI. During these discussions, participants’ viewpoints changed a median of 1.0 times, and they encountered a median of 1.0 conflicting dimension with the AI model (based on our pilot study, we defined “conflicting” by a 15% opinion difference threshold). It’s interesting to note that although Deliberative AI only highlighted a median of 1 conflicting dimension, participants discussed a median of 2.0 dimensions with the AI during these interactions, indicating proactive engagement in conversations, even when not explicitly prompted.
Regarding the time spent in making a decision, as shown in Figure 12 (b), participants took the most time to make a decision in Deliberative AI (a median of 236 seconds), followed by a median of 33 seconds in XAI, and a median of 23 seconds in Human Alone, all pair-wise comparisons are significant ($p$¡0.001). These statistics indicate that humans make “quick” decisions with traditional XAI. Compared to without AI assistance, participants only spent an average of 10 seconds longer on each decision in the XAI condition. This also echoes previous findings that people will mistakenly regard explanations as a manifestation of AI capabilities (Bertrand et al., 2022), resulting in over-reliance. Since in Deliberative AI humans need to spend more time in the deliberative discussion process, it might force humans to make “slow” decisions (i.e., invoke more System 2 thinking (Kahneman, 2011)), which is more appropriate for high-stakes decision-making.
We also assessed how participants’ opinions were influenced by AI’s opinions in both the Deliberative AI and XAI conditions by calculating the Weight of Advice (WOA). Figure 13 shows a clear contrast: the WOA in the Deliberative AI condition had a notably lower median (0.085) compared to XAI (median: 0.538, $p$¡0.001). This indicates that participants in the Deliberative AI condition tended to rely more on their own predictions. It’s noteworthy that the lower quartile for Deliberative AI is 0, implying that at least 25% of participants did not alter their opinions after seeing AI suggestions. These findings align with participants’ lower reliance on AI observed in Figure 8.
6.4. RQ4: How will humans perceive the effectiveness of the proposed Human-AI Deliberation and what can be improved for future Human-AI Deliberation design?
In addition to quantitative measures, we aim to comprehend participants’ in-depth perceptions regarding the helpfulness of the proposed Human-AI Deliberation and the feature of Deliberative AI (especially for the opinion updating feature). We also seek valuable insights to enhance the future design of Human-AI Deliberation. We analyzed participants’ open-ended feedback using their conversation logs as support. We summarize our key findings in Figure 14.
Figure 14. The main results of our thematic analysis of the open-ended questions.
[…]
6.4.1. Perceptions of the Deliberative AI and the discussion process
Overall, participants provided both positive and negative feedback on Deliberative AI and the deliberative discussion process. 43 out of 48 participants thought the deliberative discussion helped them make more informed decisions. Specifically, discussion with AI help humans “identify AI’s problems” (21/48), “consider different aspects” (10/48), “reflect and realize their mistakes” (15/48). For instance, Participant 1 (Male, 32) pointed out issues with AI’s knowledge:
”The AI relies way too much on GPA scores but underestimates the role of recommendation letters. It did not give me convincing justifications. I cannot rely too much on the AI’s opinion on these two dimensions.” (P1, Male, 32)
Participant 19 (Male, 42) appreciated how the discussions inspired self-reflection:
”The AI can cause you to doubt what you believe to be sufficient for an opinion, especially according to the information/data given.” (P19, Male, 42)
These findings align with the content of the human-AI conversations we collected. In 32% of the conversations, participants expressed doubts or questions to AI, in 17% of the conversations, participants acknowledged AI’s arguments, and in 15% of the conversations, participants engaged in reflection and self-correction.
Besides, 18 participants acknowledged that discussing with AI can “offer new knowledge, insights, and fresh perspectives”. For example, P5 (Male, 45) said
”It provided some information that I actually did not know (such as the percentile, how similar stats impacted other decisions in the past, etc.), which I thought were extremely helpful.”
Moreover, nine participants explicitly acknowledged that discussion “helps them identify biases”. For example,
”There were times where it made me realize the AI probably had an inherent bias so all in all I think it helped to actually make me pause and reflect.” (P35, Female, 29)
This observation is substantiated by an examination of participants’ conversation histories during the discussion process. Remember that the AI model assigns a positive feature importance to [Country belongs to the US] while a negative feature importance to feature [Country belongs to Asia]. We are gratified to note that a substantial portion of participants (15 out of 48) identified biases in AI’s opinions. For instance, P3 (Male, 42) questioned, “Why does their country matter? They are being penalized for being from a certain country, which could be interpreted as being racist.” Similarly, P17 (Female, 60) inquired, “Could you explain why you see an engineering major and Asian origin as negatives?” These insightful conversations indicate that with Deliberative AI, participants could identify the AI model’s biases in the decision-making process.
What’s more, 15 participants thought the discussion “promoted a balance between AI’s objective knowledge and human’s subjective opinions”. For example,
”It provides an opportunity to consider multiple perspectives and statistical data, which can lead to a more informed and balanced decision-making process.” (P2, Male, 42)
”I think that receiving the feedback with the data behind the AI is a helpful visual reminder of how much above/below the average the applicant was. Being able to rely on the AI for this and focusing on my own opinions that are more subjective helped me to balance.” (P48, Male, 28)
However, we also identified limitations in the current discussion process based on participants’ feedback. Regarding the discussion’s limitations, five participants felt their “agency gets reduced” as the AI prompted them to think other than passively waiting for their questions. Additionally, four participants found the deliberative discussion “mentally demanding” and believed the AI’s responses were overly verbose. These findings align with the results of the effects on participants’ user experience (see Sec. 6.2.2).
Regarding the Deliberative AI’s limitations, two common concerns were that it “relies too heavily on data over logic” (mentioned by 15 participants) and “fails to consider human reasoning” (mentioned by 10 participants). This feedback aligns with our analysis of participants’ conversations, where humans integrated personal experiences and logic into their decision-making. For instance, they considered factors such as “Business applications are highly competitive, so a higher GPA is required.” (P17, Female, 60), “A good SOP represents that the applicant has an in-depth understanding of the project and a strong willingness.” (P32, Male, 28), “Middle-ranking schools may give students higher GPA, so a GPA of 3.6 may not have a particularly positive impact.” (P29, Male, 50). In contrast, the AI’s responses lacked in-depth analysis of these human justifications. Furthermore, some participants expressed concerns about the AI’s inability to remember historical discussions (five), its failure to consider the bigger picture (three), and its lack of transparency (three).
6.4.2. Perceptions of Deliberative AI’s updating
Participants generally appreciated the feature of dynamical updating. 25 out of 48 participants found that the AI updates “improved the user experience.” They described the experience as more interactive, akin to “interacting with a real person” (6 participants), where their opinions were “listened to and respected” (19 participants). Furthermore, these dynamic updates made the discussion process “more engaging” (8 participants) and the overall decision-making process “more collaborative” (5 participants).
However, perceptions of how AI updated its opinions varied. While 11 participants felt the magnitude of AI updates was too sensible (e.g., “the AI updated too much”), 13 participants expressed uncertainty about “why and how the AI updated.” For instance, one participant mentioned,
”I’m not sure if the AI really changed its opinions because of what I said or because it was programmed to do so in response to my judgments”. (P2, Male, 42)
Interestingly, some participants believed that AI changed its views too easily, while others felt it didn’t compromise enough. Additionally, AI’s opinion updates influenced participants’ perceptions of AI. Eight participants believed that “AI could take qualitative information into account”, and five participants thought “AI is learning human knowledge from the discussion process”, which could “improve the quality of decision-making” (3 participants). Three participants reported that the lack of transparency in the AI update process made them “doubt the quality of the opinions given by the AI at the beginning”. Therefore, although AI updates imitate the behavior of deliberation between humans, in order to make AI updates more meaningful, it is necessary to consider the perceptions of different users and conduct a more targeted design.
6.4.3. Opportunities for future improvements of Deliberative AI
Participants offered insightful suggestions for enhancing Deliberative AI in the future.
  • Deeper Understanding of Humans: 15 participants stressed the need for AI to “better comprehend humans’ thoughts and arguments”, desiring more nuanced and context-aware communication.
  • Enhanced Memory: 8 participants emphasized the importance of AI “remembering previous interactions for more personalized and coherent conversations”.
  • Holistic Problem Solving: 12 participants recommended that AI “think beyond statistics” and “provide arguments with broader considerations”, offering more comprehensive suggestions.
  • Deeper Analysis: 10 participants advocated for AI to delve deeper into discussions, “conducting more thorough probing and “analyzing the implications beyond the superficial meaning of an attribute within a given context”.
  • More Transparency: 8 participants expressed the desire for more transparency from AI, seeking explanations for “why AI’s opinions changed during conversations”. This transparency was seen as essential for building trust and understanding.
Regarding the design of the human-AI discussion process, 8 participants suggested making the conversation “more conversational and human-like”, with 11 participants favoring “concise AI responses”. Additionally, 2 participants recommended “progressively disclosing AI’s arguments”, while 3 participants desired more user agency in leading the discussion process. In addition, 3 participants also suggested improving the dialogue generation speed and visual information display.
Things to Note about Interpreting the Results. We explored the possible impact of Human-AI Deliberation on decision-making from different aspects. However, we caution readers to consider these results as exploratory findings. Since we did not recruit real members of the graduate admissions committee to participate in our study, we do not know whether Deliberative AI can help with real graduate admissions decisions. In addition, the generalizability of these results has yet to be verified. In our study, Deliberative AI only had 50% accuracy on the selected “challenging” task cases, so it is uncertain whether similar findings can appear on “easy” cases where AI has high performance. In addition, due to the length of discussion, we only asked participants to solve four task cases with Deliberative AI, which also limits the generalizability of the results. Nonetheless, we believe these exploratory findings are valuable in providing an initial understanding of the potential of Deliberative AI and the impact it may have on decision-making. Common issues identified in this study can shed light on future designs of Deliberative AI.
7. Discussion
In this paper, we introduce Human-AI Deliberation, which aims to address two common challenges in AI-assisted decision-making. First, we tackle the issue of humans’ insufficient analytical thinking when a recommendation is directly presented. Second, we address the problem of limited interaction support for resolving human-AI disagreements. Existing interfaces only allow individuals to accept or reject AI suggestions as a whole, hindering the utilization of human-AI nuanced knowledge and collaborative problem-solving. Our exploratory assessment in the context of an illustrative task reveals that Human-AI Deliberation has the potential to enhance decision accuracy and promote appropriate reliance on AI recommendations compared to traditional explainable AI assistants. Analysis of questionnaire ratings, human behaviors, and open-ended feedback provides valuable insights. We now delve into our key findings and discuss design implications and critical limitations.
7.1. Deliberation as a New Paradigm Complementing Existing (X)AI Assistance
The proposed Human-AI Deliberation aims at engaging humans in the deliberation process through conflict-oriented discussions. It can serve as a powerful supplement to existing AI assistance. On the one hand, Deliberative AI still plays the role of an assistant and its underlying component is based on existing XAI. Although it will refute, it always respects human opinions. People have the full agency as the final decision-maker. Additionally, Deliberative AI can be seen as an enhanced version of an AI assistant that engages people in analytical thinking. Some interaction design of Deliberative AI aligns with prior work, such as encouraging participants to form independent opinions before seeing AI suggestions (e.g., Cognitive Forcing Function (Buçinca et al., 2021)) and utilizing AI-generated questions to stimulate critical thinking (e.g., employing AI-framed questioning (Danry et al., 2023)).
On the other hand, Human-AI Deliberation makes important contributions to stepping outside the boundaries of traditional AI/XAI for AI-assisted decision-making. People are no longer just recipients of AI suggestions and explanations, but need to actively participate in discussions and explain their own ideas, which contributes to mutual communication and transparency between humans and AI. This communication can lead to more informed and mature decisions as deliberations progress.
We suggest that the design of future AI assistants should not only focus on the design of the AI side (such as improving the performance of AI or improving the presentation of AI auxiliary information), but also draw on human-to-human decision-making methods and theories to explore design opportunities, especially from the perspective of communication between humans and AI.
7.2. Reducing Human Over-Reliance by Exposing AI Mistakes
Human-AI Deliberation significantly diminishes participants’ tendency to over-rely on incorrect AI suggestions. A pivotal element is the adoption of cognitive forcing theory, as proposed by Buccinca et al. (Buçinca et al., 2021), which motivates individuals to formulate their opinions before being exposed to AI recommendations. This method effectively counters the anchoring bias—the initial influence of AI’s erroneous predictions on judgment—thereby encouraging more analytical, System 2 thinking  (Kahneman, 2011). Notably, our Explainable AI (XAI) baseline also requires participants to articulate their perspectives independently before receiving AI counsel and explanations. Despite this stringent baseline, our Deliberative AI approach shows enhanced effectiveness in reducing over-reliance. Thus, the observed decrease in over-reliance within our Deliberative AI extends beyond mere cognitive forcing, attributed also to the promotion of thoughtful deliberation between individuals and AI amidst contrasting viewpoints.
Within the Deliberative AI, as participants navigate through conflicting opinions with AI, they become increasingly cognizant of AI limitations, further mitigating over-reliance. Furthermore, 31% of conversations involve participants expressing skepticism towards the AI’s opinions and logic. However, the decline in over-reliance is not due to people blindly doubting AI. Notably, our strategy significantly reduces over-reliance on AI without increasing under-reliance. This indicates that through deliberative discourse, participants critically assess AI’s opinions on a case-by-case basis, integrating accurate AI insights while amending their own assessments as necessary and identifying instances where AI judgments may be erroneous. Consequently, we recommend that designers of AI-assisted decision-making systems prioritize transparency and explicitly highlight AI’s potential errors, rather than merely hinting that “AI may make errors”.
7.3. Human-AI Conflict Resolving Is The Key to Improving Collaboration
We contend that resolving conflicts is inherently more beneficial for enhancing decision-making processes than merely seeking consensus. Our methodology, Human-AI Deliberation, emphasizes the resolution of disputes between humans and AI, a dimension presently overlooked in AI-assisted decision-making frameworks. We argue that conflict serves as a crucial element in human-AI collaboration, unveiling inherent flaws and biases. This conflict-centric approach yields multiple advantages: it enhances the accuracy of decision-making by mitigating over-reliance on AI, fosters introspection, and facilitates the reconciliation of differing opinions with AI recommendations. Furthermore, by prioritizing conflict, humans can better recognize biases in AI interpretations, thereby advancing fairness in decision-making processes (Kramer et al., 1990; Hochman et al., 2015).
Nevertheless, an emphasis on conflict resolution—despite its apparent benefits—may diminish user satisfaction. Engaging in deliberation with AI, particularly around divergent viewpoints, necessitates that individuals critically evaluate their own perspectives and articulate their arguments. This process amplifies the perception of ”dissenting voices,” potentially compromising the user experience. Hence, we advocate for a balanced approach in future design paradigms: one that centers on conflict resolution while simultaneously considering strategies to enhance user engagement with conflicts without detracting from the overall experience.
7.4. Obstacles to Discussion: Humans and AI Think Differently
Humans Use Heuristics and Logic, while AI Relies on Data: Combining LLM and DS-Model empowers Deliberative AI in dynamic discussions with humans. However, AI’s data-centric approach clashes with human decision-making, which integrates personal understanding and experiences (Ghattas et al., 2014), logic (Evans, 2002), heuristics (Doswell, 2004), and creativity (Simon, 1990, 1997). To address this, we propose DS-Model guiding LLMs for data-related discussions and LLMs autonomously handling data-irrelevant matters based on DS-Model’s predictions. While effective, this approach faces limitations in engaging users with strong subjective perspectives and reasoning (Liu et al., 2023).
Existing Explanation Methods Are Insufficient to Enable Discussion: Current XAI methods, like local feature importance (Liao and Varshney, 2021), inadequately clarify specific feature contributions. Our solution, based on training data, provides data-related insights, such as data distribution and value comparisons, to empower AI in feature-level discussions. However, challenges arise. First, patterns inferred from data may not precisely align with AI’s explanations. Second, users often rely on various forms of reasoning beyond data, including subjective experiences (Doswell, 2004; Ghattas et al., 2014). This finding is aligned with Miller et al.’s work (Miller, 2019), ”probabilities are not as important as causal links.” Hence, offering data-driven insights as meta-explanations may not fully align with users’ thinking styles or replicate human-like conversations.
Therefore, future human-AI deliberation designs should try to cope with humans’ intuition and cognitive processes and design human-compatible AI explanations to empower more effective human-AI discussions.
7.5. Ethical Concerns in Deliberative AI Design
Responsibly Utilizing LLMs for Assisting Human Decision-Making. LLMs are advancing rapidly but face a critical issue known as “hallucination”, where they generate seemingly credible but false information (Ji et al., 2023). Engaging in discussions with users solely based on LLM-generated opinions, without proper fine-tuning of task data, is highly irresponsible. Even with fine-tuning, unpredictable responses can occur. To address this, we advocate for a responsible approach that combines LLMs and DS-Model. DS-Model’s output tightly controls LLMs’ responses, with LLMs primarily acting as a communication bridge between users and DS-Model. While this approach restricts LLMs’ flexibility to some extent, it enhances security and controllability. However, it’s essential to note that even with DS-Model and training data guiding LLMs’ responses, the potential for error information remains. Therefore, researchers planning to employ LLMs in decision-making should exercise caution and transparency. It’s crucial to inform users about the possibility of inaccurate information and potential pitfalls associated with LLM-generated responses.
Concerns Regarding AI Updating. Research on human-AI discussions highlights a desire for AI to actively engage with human opinions (Zheng et al., 2023). To address this, we’ve implemented an AI opinion update mechanism, allowing AI to adapt its stance based on user arguments’ rationality and prediction uncertainty. Although users generally appreciate the AI’s responsiveness and the feeling of being heard and respected, ethical concerns have emerged, particularly regarding accountability. One pivotal question arises: Who should be held accountable if the AI initially had the correct prediction but updated to an incorrect one after engaging in discussions with the human? Furthermore, the AI’s ability to adapt its viewpoints may give humans the impression of learning and progress, despite the underlying AI model remaining unchanged. This can impact user expectations and potentially lead to an overestimation of AI capability. While the AI’s capacity to update opinions is promising, it’s crucial to remain vigilant about potential adverse outcomes during the design phase. Most importantly, future designs should prioritize transparency in the AI updating mechanism to help users fully comprehend the principles governing AI updates and effectively manage their expectations.
7.6. Applicability and Generalizability
Although the human-AI deliberation framework we introduced is potentially applicable and can theoretically be adapted to various decision-making scenarios, deliberation may not be suitable for all tasks. First, the deliberative process requires additional time and cognitive resources, making it less suitable for repetitive, low-stakes decisions such as content moderation (Lai et al., 2022). Instead, it may be more beneficial in high-stakes decision-making environments, such as healthcare (Lee et al., 2021), financial investments (Green and Chen, 2019), and criminal justice (Chiang et al., 2023), especially when dealing with complex or edge cases. Secondly, Human-AI Deliberation is better suited to handle challenging task cases that are difficult for both humans and AI. Furthermore, our proposed method of combining large language models (LLMs) with domain-specific (DS) models can be adapted to other decision-making scenarios. It does not limit the form of the DS model, which can range from a simple machine learning classifier to a complex neural network model or even another specific large model.
7.7. Limitations and Future Work
The study design of this paper has several limitations. First, the illustrative task (college admission decision) we chose disconnects with real college admission. On the one hand, limited by the available public data set, the data set we used does not contain the raw texts of some important dimensions, such as purpose statements, diversity statements, recommendation letters, etc. Thus, in our study, participants could only discuss the ratings on them rather than what information is included in them. Furthermore, the participants we recruited were not members of the actual admissions committee. Although we require participants to have at least a master’s degree and graduate application experience, there is still a considerable gap in expertise. Therefore, the graduate admission task is just an illustrative task to explore the instantiation and potential of the proposed Human-AI Deliberation framework. We need future studies to truly explore whether and why Deliberative AI yields improvement in real-world college admission decision-making.
Second, the number of decision task samples used in our study is relatively limited. We tried to include 8 or 10 decision task samples in our pilot study, which resulted in participants taking more than one hour to complete all decision tasks. In addition, the users’ engagement dropped substantially after completing three or four tasks, and they simply dragged the slider on the interface to give their opinions without discussing the conflicting points with AI. We infer that long tasks lead to user fatigue, which will affect the validity of the study. Therefore, due to concerns about the length of the study, we only selected four task samples, which to some extent limits the generalizability and significance reliability of the experimental results. Future work needs to test the effects of Human-AI Deliberation through a long-term study so that more deliberative decision data can be collected.
Third, the AI in our study is of low performance (50% accuracy), not a conventional AI that performs better than humans. We made this design choice because Human-AI Deliberation is more suitable for dealing with task cases that are challenging for both humans and AI. However, we acknowledge that the decision-making dynamics assisted by AI with different performances can be different. Thus, all our results are based on the premise that AI’s performance is relatively low and on par with humans’ independent decision-making performance. Therefore, we cannot say whether some of the positive outcomes of Deliberative AI (such as a reduction in people’s overdependence) will persist when the AI performs very well. Future work can further consider the variation of human and AI decision performance.
Finally, our study shows the potential of Deliberative AI compared to XAI, but it is still unclear which components of Deliberative AI contribute to the benefits. We acknowledge that Deliberative AI has more features than the XAI condition. For example, Deliberative AI can present distribution information of the training data (e.g., percentiles) in the conversation (as evidence of AI in the Human-AI Deliberation framework), which XAI does not have. This is because we chose to apply the most commonly used XAI display method as the baseline. These XAI systems typically only provide suggestions and explanations without providing data distribution information. In addition, Deliberative AI consists of multiple critical components. A future ablation study could help investigate the role and importance of each component for a systematic comparison.
8. Conclusion
In this paper, we introduce the framework of Human-AI Deliberation, as a new paradigm of AI assistance for decision-making. Human-AI Deliberation encourages the externalization of thoughts, facilitates interactive deliberation between humans and AI, and allows for dynamic updates of decisions. To enable the deliberation, we present a novel AI assistant called Deliberative AI, which can identify differences in viewpoints, engage in comprehensive deliberation, and adapt its suggestions during discussions. We apply this framework to an illustrative task (graduate admissions decisions) and conduct an exploratory study to assess its potential impact on decision-making processes, outcomes, user perceptions, and experiences. Results indicate the potential of Deliberative AI to improve decision accuracy and promote more appropriate human reliance on AI. Additionally, we analyze participants’ behaviors and open-ended feedback to gain deeper insights into how users use and perceive Deliberative AI, uncovering areas for improvement. Lastly, we offer design implications based on our findings, highlighting the potential of Human-AI Deliberation as a valuable supplement to current AI-assisted decision-making.
Appendices
Appendix A Base Interface
Figure 15. The baseline (traditional explainable AI) interface in our user study. The interface contains three parts. The top (A) is the applicant’s profile. The bottom left part (B) shows AI’s feature contribution explanation. The bottom right part (C) is for humans to (1) indicate their initial predictions, (2) see AI’s suggestions, and (3) indicate their final decisions. Note that AI’s suggestions and explanations are only shown after humans make their initial predictions. (All the dashed lines are only for illustration)
[…]

### v2

@@
 Abstract.
-In AI-assisted decision-making, humans often passively review AI’s suggestion and decide whether to accept or reject it as a whole.
-In such a paradigm, humans are found to rarely trigger analytical thinking and face difficulties in communicating the nuances of conflicting opinions to the AI when disagreements occur.
-To tackle this challenge, we propose Human-AI Deliberation, a novel framework to promote human reflection and discussion on conflicting human-AI opinions in decision-making.
-Based on theories in human deliberation, this framework engages humans and AI in dimension-level opinion elicitation, deliberative discussion, and decision updates.
-To empower AI with deliberative capabilities, we designed Deliberative AI, which leverages large language models (LLMs) as a bridge between humans and domain-specific models to enable flexible conversational interactions and faithful information provision.
-An exploratory evaluation on a graduate admissions task shows that Deliberative AI outperforms conventional explainable AI (XAI) assistants in improving humans’ appropriate reliance and task performance.
-Based on a mixed-methods analysis of participant behavior, perception, user experience, and open-ended feedback, we draw implications for future AI-assisted decision tool design.
-AI-Assisted Decision-making, Human-AI Collaboration, Appropriate Reliance
-ccs: Human-centered computing Empirical studies in HCI
+Traditional AI-assisted decision-making systems often provide fixed recommendations that users must either accept or reject entirely, limiting meaningful interaction—especially in cases of disagreement.
+To address this, we introduce Human-AI Deliberation, an approach inspired by human deliberation theories that enables dimension-level opinion elicitation, iterative decision updates, and structured discussions between humans and AI.
+At the core of this approach is Deliberative AI, an assistant powered by large language models (LLMs) that facilitates flexible, conversational interactions and precise information exchange with domain-specific models.
+Through a mixed-methods user study, we found that Deliberative AI outperforms traditional explainable AI (XAI) systems by fostering appropriate human reliance and improving task performance.
+By analyzing participant perceptions, user experience, and open-ended feedback, we highlight key findings, discuss potential concerns, and explore the broader applicability of this approach for future AI-assisted decision-making systems.
+AI-Assisted Decision-making, Human-AI Collaboration, Deliberation, Appropriate Reliance, Large Language Models
+journalyear: 2025copyright: acmlicensedconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3713423isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing Empirical studies in HCI
 1. Introduction
@@
 Concerns surrounding AI’s accuracy, safety, ethics, and accountability (Cai et al., 2019b; Lee et al., 2021; Binns et al., 2018) have led to the widespread adoption of the AI-assisted decision-making paradigm in real-world applications (Buçinca et al., 2021; Zhang et al., 2020; Wang and Yin, 2021; Bansal et al., 2021).
-In this paradigm, AI performs an assistive role by providing a recommendation, while human decision-makers can choose to accept or reject AI’s suggestion in their final decision (Lai et al., 2021).
-Research in recent years, however, identified two challenges within the existing AI-assisted decision-making paradigm.
-First, a battery of empirical studies found that people rarely trigger analytical thinking (i.e., System 2 thinking (Kahneman, 2011)) when directly presented with AI’s suggestions (Buçinca et al., 2020; Rastogi et al., 2022; Bertrand et al., 2022).
-As a result, people frequently over-rely on the AI’s incorrect recommendations (over-reliance) or mistakenly ignore AI’s correct suggestions (under-reliance) (Buçinca et al., 2021; Ma et al., 2023; Wang and Yin, 2021).
-Although some solutions have been proposed, such as displaying AI explanations (Bansal et al., 2021) and forcing people to think (e.g., applying cognitive forcing functions (Buçinca et al., 2021)), the results are mixed at best (Lai et al., 2022; Poursabzi-Sangdeh et al., 2021) (Challenge 1).
-Second, instead of full consensus or complete divergence, human and AI decision rationales often exhibit partial alignment (Sivaraman et al., 2023; Wang and Yin, 2023).
-While they may concur on certain aspects, differences may persist on others (Miller, 2023).
-However, in current AI-assisted decision-making systems, AI consistently provides a fixed recommendation regardless of human thoughts and humans can only accept or reject AI’s recommendation as a whole (Lai et al., 2021), with limited support for resolving conflicts or engaging in a meaningful exchange of ideas with the AI system (Miller, 2023).
-For example, as shown in Figure 1 (a), when the human decision-maker’s prediction is inconsistent with the AI model’s recommendation and the human only partially agrees with the AI’s reasoning (e.g., explanation), existing AI-assisted decision-making interfaces do not support any communication between humans and AI regarding conflicting opinions.
-This limitation may impede the effective utilization of both human and AI knowledge, hindering collaborative and complementary human-AI team performance (Challenge 2).
+In this paradigm, AI performs an assistive role by providing a recommendation, while human decision-makers can choose to accept or reject it in their final decision (Lai et al., 2021).
 Figure 1.
@@
 […]
+Research in recent years, however, identified two challenges within the existing AI-assisted decision-making paradigm.
+[Challenge 1] First, a battery of empirical studies found that people rarely trigger analytical thinking when directly presented with AI’s suggestions (Buçinca et al., 2020; Rastogi et al., 2022; Bertrand et al., 2022).
+As a result, people frequently over-rely on the AI’s incorrect recommendations (over-reliance) or mistakenly ignore AI’s correct suggestions (under-reliance) (Buçinca et al., 2021; Ma et al., 2023; Wang and Yin, 2021).
+Although some solutions have been proposed, such as displaying AI explanations (Bansal et al., 2021) and forcing people to think more effortfully (Buçinca et al., 2021), the results are mixed at best (Lai et al., 2022; Poursabzi-Sangdeh et al., 2021).
+[Challenge 2] Second, instead of full consensus or complete divergence, human and AI decision rationales often exhibit partial alignment (Sivaraman et al., 2023; Wang and Yin, 2023).
+While they may concur on certain aspects, differences may persist on others (Miller, 2023).
+However, in current AI-assisted decision-making systems, AI always provides a fixed recommendation regardless of human thoughts and humans can only accept or reject AI’s recommendation as a whole (Lai et al., 2021), with limited support for resolving conflicts or engaging in a meaningful exchange of ideas with the AI system (Miller, 2023).
+For example, as shown in Figure 1 (a), when the human decision-maker’s prediction is inconsistent with the AI model’s recommendation and the human only partially agrees with the AI’s reasoning (e.g., explanation), existing AI-assisted decision-making interfaces do not support any communication between humans and AI regarding conflicting opinions.
+This limitation may impede the effective utilization of both human and AI knowledge, hindering collaborative and complementary human-AI team performance.
 Deliberation, characterized by thoughtful and reasoned discussion, plays a pivotal role in facilitating constructive discourse and consensus-building across various contexts (Steenbergen et al., 2003; Bächtiger and Parkinson, 2019).
-Deliberation proves effective in facilitating various human decision-making tasks, including deliberative politics (Black et al., 2010; Habermas, 2005; Thompson, 2008), clinical diagnosis (Schaekermann et al., 2019; Inguaggiato et al., 2019; Preisz, 2019), criminal justice (Devine et al., 2001; Van Berkel et al., 2019), and more.
+Deliberation proves effective in facilitating diverse human decision-making tasks, including deliberative politics (Black et al., 2010; Habermas, 2005; Thompson, 2008), clinical diagnosis (Schaekermann et al., 2019; Inguaggiato et al., 2019; Preisz, 2019), criminal justice (Devine et al., 2001; Van Berkel et al., 2019), among others.
 It offers individuals an opportunity to rigorously evaluate different perspectives, including their own, which can potentially address Challenge 1 in AI-assisted decision-making.
@@
 Despite the potential benefits of deliberation, how to design mechanisms to facilitate deliberative conversation between humans and AI and how deliberations influence AI-assisted decision-making remain to be explored.
-In this paper, based on guidelines and theories in human deliberation (Steenbergen et al., 2003; Bächtiger and Parkinson, 2019; Bächtiger et al., 2009) and decision-making (Alvarez-Melis et al., 2021; Weed, 2005; Baumann et al., 2014), we introduce a novel framework, Human-AI Deliberation for AI-assisted decision-making (Figure 1 (b)).
-Instead of presenting a fixed AI suggestion for humans to accept or reject, our framework encourages humans to externalize their thoughts, enables an interactive deliberation process between humans and AI around the conflicting points of their opinions and rationales, and fosters dynamic, fine-grained updates of humans and AI’s decisions.
-The key component of this framework is Deliberative AI, which has the ability to locate viewpoint dissimilarities, stimulate comprehensive deliberation with human decision-makers, and make necessary changes, even compromises, in its own suggestion as the constructive discussion unfolds.
+In this paper, building on established guidelines for enhancing discourse quality and identifying common ground in human deliberation (Steenbergen et al., 2003; Bächtiger and Parkinson, 2019; Bächtiger et al., 2009), as well as the weight-of-evidence approach in decision-making (Alvarez-Melis et al., 2021; Weed, 2005; Baumann et al., 2014), we propose a novel solution: Human-AI Deliberation for AI-assisted decision-making (Figure 1 (b)).
+Instead of presenting a fixed AI suggestion for humans to accept or reject, our approach encourages humans to externalize their thoughts, enables an interactive deliberation process between humans and AI around the conflicting points of their opinions and rationales, and fosters dynamic, fine-grained updates of humans and AI’s decisions.
+The key component of this approach is Deliberative AI, which has the ability to locate viewpoint dissimilarities, stimulate comprehensive deliberation with human decision-makers, and make necessary changes, even compromises, in its own suggestion as the constructive discussion unfolds.
 To design such an AI assistant, we propose to integrate the strength of domain-specific models (for reliable assistant information generation) and Large Language Models (LLMs, for interactivity and conversation capability).
-We elaborate on the design of Human-AI Deliberation framework and Deliberative AI in Section 3 and demonstrate how to instantiate the framework in an illustrative task in Section 4.
-We realize that Human-AI Deliberation might be especially suitable for handling task cases that are challenging for both humans and AI.
-If both parties are very good at the task cases (both have high decision accuracy) or one party is very proficient in the task cases, deliberation may not be necessary.
-Therefore, we study the impact of Human-AI Deliberation when the task case is challenging for both parties.
-Specifically, using our proposed Human-AI Deliberation as a research probe, we aim to explore the following research questions.
-When dealing with task cases that are challenging for both humans and AI:
-  • RQ1: How will Human-AI Deliberation affect task performance and humans’ reliance (and its appropriateness) on AI suggestions?
-  • RQ2: How will Human-AI Deliberation affect humans’ perceptions of the AI and user experience?
-  • RQ3: How will humans engage in the deliberation process when working with Deliberative AI?
-  • RQ4: How will humans perceive the effectiveness of the proposed Human-AI Deliberation and what can be improved for future Human-AI Deliberation design?
-To answer these questions, we conducted an exploratory study in the context of an illustrative task (college graduate admission).
+We elaborate on the architecture design of Human-AI Deliberation and Deliberative AI in Section 3 and demonstrate how to instantiate the architecture in an illustrative task in Section 4.
+Since the primary purpose of deliberation is to resolve conflicts between human and AI perspectives, we intentionally selected task cases with notable human-AI disagreements for our user study, drawing from insights in our pilot study.
+As human-AI deliberation is designed to both resolve conflicts and mitigate inappropriate human reliance on AI, we investigate its impact on task performance and human reliance on AI.
+Additionally, since deliberation explicitly highlights these conflicts, we are particularly interested in examining its effect on human perceptions of AI and the overall decision-making experience.
+Specifically, using our proposed concept of Human-AI Deliberation as a research probe, we aim to explore the following research questions.
+  • RQ1: How will Human-AI Deliberation affect task performance and humans’ reliance on AI suggestions?
+  • RQ2: How will Human-AI Deliberation affect humans’ perceptions of the AI partner and their user experience?
+  • RQ3: How will humans perceive the effectiveness of the proposed Human-AI Deliberation and what can be improved for future Human-AI Deliberation design?
+To answer these questions, we conducted an exploratory study using a graduate admissions task.
 We recruited participants with graduate admissions experience (at least once admitted to a graduate program) on Prolific and asked them to predict an applicant’s chance of getting an offer based on the applicant’s profile.
-Based on a pilot study, we selected some task cases that were edge cases and therefore challenging for both humans and AI (e.g., their accuracy on these task cases was around 50-55%).
 We compared the proposed Deliberative AI with traditional explainable AI (XAI) and human alone baselines.
 Our experimental results revealed that Human-AI Deliberation has the potential to enhance decision accuracy and promote appropriate reliance on AI recommendations compared to traditional XAI assistants.
-We then derived valuable insights by analyzing participants’ questionnaire ratings, human behaviors, conversation logs, and open-ended feedback.
+We conclude by discussing key implications and addressing generalizability concerns based on our design and study findings.
 In summary, we make three contributions:
-  • We propose a novel conceptual framework, Human-AI Deliberation, grounded in decision-making and deliberation theories, for humans and AI to deliberate their conflicting thoughts in a decision task.
-Accordingly, we design the architecture of Deliberative AI that is capable of deliberation with humans through a novel integration of domain-specific models and LLMs.
-  • We demonstrate the instantiation of the framework and Deliberative AI in an illustrative task (college graduate admission), including the implementation of different components and interface design.
-  • We conduct an exploratory study to gain an initial understanding of how Human-AI Deliberation might impact the decision-making process and how humans would perceive this novel AI assistance.
+  • We propose a novel architecture, Deliberative AI, designed to enable Human-AI Deliberation—a collaborative process where humans and AI deliberate together to resolve conflicting perspectives in decision-making tasks—by seamlessly integrating domain-specific models with large language models (LLMs).
+  • We demonstrate the instantiation of the Deliberative AI in an illustrative task (college graduate admission), including the implementation of different components and interface design.
+  • We conduct an exploratory study to gain an initial empirical understanding of how Human-AI Deliberation might impact the decision-making process and how humans would perceive this novel AI assistance.
 Additionally, we demonstrate its potential to improve decision accuracy and promote appropriate human reliance on AI.
 2. Related Work
-2.1. AI-Assisted Decision-Making: Inadequacies and Challenges
-Artificial Intelligence (AI) is increasingly used in decision-making across various domains (Ma et al., 2019, 2022b; Dastin, 2018; Dilsizian and Siegel, 2014; Khandani et al., 2010; Zhu and Ma, 2019).
+2.1. AI-Assisted Decision-Making: Objectives and Challenges
+Artificial Intelligence (AI) is increasingly used in decision-making across various domains (Ma et al., 2019, 2022b; Dastin, 2018; Dilsizian and Siegel, 2014).
 However, AI’s real-world applications are not infallible, still far from 100% accuracy (Shi et al., 2023; Gao et al., 2018; Ma et al., 2022a).
@@
 To address this, the prevalent paradigm of AI-assisted decision-making has emerged, drawing substantial attention in the Human-Computer Interaction (HCI) and AI communities (Buçinca et al., 2021; Zhang et al., 2020; Wang and Yin, 2021; Bansal et al., 2021).
-In this paradigm, AI takes on a supportive role, offering recommendations for human decision-makers to accept or reject in their final decisions (Lai et al., 2021).
-Research in AI-assisted decision-making encompasses various objectives, such as team performance, decision fairness, efficacy, efficiency, understanding of AI, trust, reliance on AI, and subjective experience (Zhang et al., 2020; Ma et al., 2023; Chiang et al., 2023; Wang et al., [n.d.]; Bansal et al., 2019a; Dressel and Farid, 2018; Wang and Yin, 2021; Cheng et al., 2019; Buçinca et al., 2021; Ma et al., 2023; Yang et al., 2020; Schemmer et al., 2023; He et al., 2023; Buçinca et al., 2020; Gajos and Mamykina, 2022; Kaur et al., 2020; Liao and Varshney, 2021; Ma et al., 2022a).
-One of the key challenges in AI-assisted decision-making is achieving complementary performance, where the collaborative decision outcome surpasses what either human or AI can achieve alone (Bansal et al., 2021; Lai and Tan, 2019; Zhang et al., 2020; Schemmer et al., 2023).
-However, empirical studies reveal difficulties in achieving this, primarily due to two issues.
+In this paradigm, AI takes on an assistant role, offering recommendations for human decision-makers to accept or reject in their final decisions (Lai et al., 2021).
+Research in AI-assisted decision-making spans a range of objectives, including enhancing team performance (Zhang et al., 2020; Ma et al., 2023), promoting decision fairness (Chiang et al., 2023; Wang et al., [n.d.]), improving efficacy and efficiency (Cheng et al., 2019), fostering understanding of AI (Wang and Yin, 2021; Cheng et al., 2019), building trust and appropriate reliance on AI (Schemmer et al., 2023; He et al., 2023), and enriching subjective user experiences (Liao and Varshney, 2021; Ma et al., 2022a).
+Among these, a key goal is achieving complementary performance—where the collaborative outcomes of human-AI teams exceed what either humans or AI could achieve independently (Bansal et al., 2021).
+Despite its importance, recent empirical studies highlight persistent difficulties in reaching this goal (Zhang et al., 2020; Bansal et al., 2021; Rastogi et al., 2022), driven by two primary challenges.
 One challenge is the underutilization of human and AI domain knowledge (Bansal et al., 2020).
@@
 Furthermore, deliberation proves effective in facilitating various decision-making tasks, including clinical diagnosis (Schaekermann et al., 2019; Inguaggiato et al., 2019; Preisz, 2019), criminal justice (Devine et al., 2001; Van Berkel et al., 2019), and more.
-Effective decision-making is of paramount importance across diverse domains, and deliberation emerges as a valuable process replete with manifold advantages.
-First, it enhances decision quality and problem-solving (Barabas, 2004; Landemore and Page, 2015).
-Deliberation involves comprehensive analysis and careful evaluation of options, fostering a holistic understanding of issues by considering various factors and potential outcomes (Jacobs et al., 2009).
-Research underscores that deliberative decisions tend to be wiser and more effective owing to their foundation in thorough information and thoughtful analysis (Landemore, 2012; Dijksterhuis et al., 2006).
-Second, deliberation promotes participation and collaboration (Nabatchi and Leighninger, 2015).
-It encourages engagement from multiple stakeholders, thereby facilitating cooperation and communication among decision-makers (Wojcieszak et al., 2010).
-This collaborative process aids in resolving complex issues, providing decision-makers with ample time for dialogue and negotiations, ultimately ensuring the widespread acceptance of decisions (Fishkin, 2018).
-Third, deliberation has the potential to mitigate decision biases and enhance fairness (Kramer et al., 1990; Hochman et al., 2015).
-It assists decision-makers in objectively analyzing issues, mitigating personal biases and emotional influences, and identifying flawed perceptions and knowledge (Tetlock, 2017).
+Effective decision-making is crucial across various domains, and deliberation offers significant advantages.
+It improves decision quality and problem-solving by enabling comprehensive analysis and evaluation of options, fostering a deeper understanding of issues and outcomes (Barabas, 2004; Landemore and Page, 2015).
+Deliberative decisions are often wiser and more effective due to their basis in thorough analysis (Landemore, 2012; Dijksterhuis et al., 2006).
+Additionally, deliberation promotes participation and collaboration, encouraging stakeholder engagement and facilitating communication, which aids in resolving complex issues and ensuring decision acceptance (Nabatchi and Leighninger, 2015; Wojcieszak et al., 2010; Fishkin, 2018).
+Finally, it helps mitigate decision biases and enhance fairness by enabling objective analysis and reducing emotional influences (Kramer et al., 1990; Hochman et al., 2015; Tetlock, 2017).
 Despite the significance of deliberation in decision-making, there is a dearth of research on its integration into AI-assisted decision-making processes.
-To address this gap, drawing upon theories and practices in deliberation (Black et al., 2010; Habermas, 2005; Thompson, 2008; Lord and Tamvaki, 2013; Steiner et al., 2005), we propose a framework called Human-AI Deliberation to facilitate human reflection and discussion on conflicting human-AI opinions.
-Based on this framework, we aim to move a first step towards designing a Deliberative AI and investigating its effects on decision processes and outcomes through an exploratory empirical study.
+To address this gap, drawing upon theories and practices in deliberation (Black et al., 2010; Habermas, 2005; Thompson, 2008; Lord and Tamvaki, 2013; Steiner et al., 2005), we propose Human-AI Deliberation to facilitate human reflection and discussion on conflicting human-AI opinions.
+Based on this approach, we aim to move a first step towards designing a Deliberative AI and investigating its effects on decision processes and outcomes through an exploratory empirical study.
 2.3. Existing Studies on Deliberation in AI-Assisted Decision Making
-Deliberation encompasses two essential facets: the analytical and contemplative consideration of issues and active discussion (Gracia, 2003).
-These aspects have garnered some attention in previous research on AI-assisted decision-making.
-Concerning the stimulation of analytical thinking in humans, one approach involves interventions aimed at encouraging deeper engagement in System 2 thinking (Kahneman, 2011) by controlling when AI information is presented.
-For instance, the concept of ”cognitive forcing” has explored strategies to compel human decision-makers to spend more time deliberating on problems (Buçinca et al., 2021; Park et al., 2019; Rastogi et al., 2022).
-Methods include asking individuals to make independent predictions before receiving AI suggestions (Ma et al., 2023; Buçinca et al., 2021) or implementing a ”slow algorithm” (Park et al., 2019; Buçinca et al., 2021).
-These cognitive forcing techniques have demonstrated the ability to reduce human reliance on AI.
-Additionally, researchers have investigated how and what information should be presented.
-For instance, Garhos et al.
-(Gajos and Mamykina, 2022) found that people engaged in more analytical thinking when AI explanations were provided without concrete recommendations.
-Furthermore, Danry et al.
-(Danry et al., 2023) introduced AI-framed Questioning to enhance critical thinking and human discernment of flawed statements.
-While these approaches yield positive results, they come with limitations, such as potentially leading to under-reliance and not addressing differences between human and AI perspectives.
-Regarding discussion, some research has explored dialogues between humans and AI in cooperative games ((FAIR)† et al., 2022; Kramár et al., 2022), but these discussions have not been specifically tailored to decision-making tasks.
-Recent efforts have started to integrate discussions into AI-assisted decision-making, such as Zheng et al.’s (Zheng et al., 2023) inclusion of AI in group decision-making to assess student essays.
-However, these efforts often rely on Wizard of Oz (WoZ) setups and lack genuinely designed AI systems for human interaction.
-Chiang et al.
-(Chiang et al., 2023) examined human-AI collaborative decision-making for recidivism risk assessment, but in their work, AI only offers suggestions and does not participate in the discussion with humans.
-Additionally, Zhang et al.
-(Zhang et al., 2023) utilize AI models as boundary objects to foster deliberation among organizations, but they do not encompass deliberation between humans and AI.
-Perhaps the most similar work to ours is that of Slack et al.
-(Slack et al., 2022).
-They focused on AI explanation and proposed using dialogue to improve the flexibility to support arbitrary follow-up questions that users might have and enhance people’s understanding of AI.
-In this paper, we also use the form of dialogue to allow people to interact with the AI, and also include the part where the AI gives explanations through dialogue (such as the AI answers users’ questions).
-But the key difference is that we focus on deliberation design and propose Deliberative AI which can not only “explain to users” but also actively engage users in the deliberative discussions by “asking or challenging” the users, aiming to promote people’s critical thinking.
-Drawing inspiration from the principles of deliberation, we introduce a novel paradigm of Human-AI Deliberation.
-Additionally, we design Deliberative AI which combines the conversational capabilities of large language models (LLMs) with the predictive abilities of domain-specific models (DS-Model).
-Deliberative AI encourages meticulous human thinking and facilitates discussions concerning disparities between human and AI viewpoints.
-Through an exploratory assessment, we investigate the potential impacts of deliberation on AI-assisted decision-making.
-Our work takes an initial step toward designing and evaluating human-AI deliberation in decision-making contexts, offering valuable insights into the integration of deliberation into AI-assisted decision-making processes.
-3. Framework: Human-AI Deliberation
-In this section, we first introduce the general process of decision-making with an emphasis on the step that weighs the evidence for an alternative.
-Then, grounded in the theoretical construct of Weight of Evidence (WoE), we propose the Human-AI Deliberation framework.
-Finally, based on theories and practices in human deliberation, we introduce the design considerations and structure of Deliberative AI, the core AI component of our framework.
-3.1. Unpacking Decision-Making
-Decision-making is a process of making a choice from a number of alternatives to achieve a desired result (Eisenfuhr, 2011).
-Existing literature summarizes the general decision-making process into seven steps (U Mass, [n.d.]; Lunenburg, 2010) – Step 1: Identify the problem, Step 2: Collect information, Step 3: Identify the alternatives, Step 4: Weigh the evidence, Step 5: Choose from the alternatives, Step 6: Implement action, and Step 7: Evaluate the results.
-In this work, we dissect a decision-making problem and human-AI thoughts into four structured components (Figure 2 (a)).
-(1) Decision means the overall decision to make for a specific problem.
-(2) Dimension refers to an aspect that people consider when reaching an overall decision.
-For instance, in a graduate admission task, dimensions could be applicants’ academic excellence, research ability, etc.
-Taking tabular task data (commonly used in decision-making (Ghai et al., 2021; Wang and Yin, 2021)) as an example, a dimension may be an attribute or a set of related attributes.
-(3) Opinion on a dimension denotes one’s views on the impact of a dimension on the overall decision (e.g., an applicant’s academic excellence contributes +50% to the overall chance of being admitted).
-(4) Evidence is the basis of an opinion.
-For humans, evidence can encompass facts, heuristics, past experiences, personal beliefs, interpretations, and even creative thoughts (Simon, 1990, 1997).
-In contrast, AI’s evidence primarily originates from information and knowledge embedded in the training data.
-3.1.1. When and Where to Deliberate?
-In the scope of this paper, we propose to conduct Human-AI Deliberation in “Weigh the Evidence” (Step 4 in the seven decision-making steps).
-For one thing, disagreements between human’s thoughts and AI’s suggestions often surface when they assess alternatives in this step.
-Diverse interpretations of evidence in Step 4 are likely to lead to conflicting opinions and disparate outcomes in steps 5-7 (Lai et al., 2021).
-For another, in existing AI-assisted decision-making tasks (Lai et al., 2021; Lai and Tan, 2019; Bansal et al., 2021), steps 1-3 are usually settled in advance and there is no need for much deliberation.
-We propose to have humans and AI deliberate their opinions at the detailed Dimension level (i.e., to what extent a dimension supports or opposes the final decision) rather than merely talking about accepting/rejecting one’s overall decision.
-In this process, both human and AI need to support their arguments with evidence and weigh the presented evidence according to credibility and probative value (Baumann et al., 2014).
-3.1.2. What to Deliberate on?
-In real-world practices, a deliberation process in Step 4 is often informed by the importance of evidence interpreted in a particular context (Cartwright and Stegenga, 2011).
-In information theory, such “importance” is quantified as the (probabilistic) weight of evidence (WoE) for a given hypothesis (and against alternative hypotheses) (Peirce, 2014; Good, 1950; Gough, 2007).
-WoE is well studied and widely adopted in various decision-making domains (Alvarez-Melis et al., 2021).
-This is because its literal meaning is easy to understand by lay users and its probabilistic calculation is simple to implement in practical applications (Saaty, 2008; Gough, 2007).
-WoE is the core content of deliberation in our proposed framework.
-It corresponds to “opinion” – the importance of a dimension to a final decision given one’s interpretation of evidence.
-In the remainder of this paper, we use the term opinion (as well as its synonyms e.g., viewpoint) and WoE interchangeably.
-It is crucial to note that during evidence weighing, both humans and AI possess strengths and limitations.
-For instance, humans may offer novel insights not present in AI training data, yet they can be constrained by time and cognitive biases, as per the “Bounded Rationality Model” (Simon, 1990, 1997).
-AI excels in data-driven knowledge and computation, but is known to be susceptible to biases in training data and falters in the face of out-of-distribution cases (Liu et al., 2021).
-These inherent human and AI shortcomings can result in potentially incorrect, incomplete, or biased evidence weighting, leading to opinion discrepancy and even decision errors.
-Hence, fostering humans’ deliberation on conflicting evidence weights could be advantageous.
+Deliberation enhances decision-making by integrating diverse perspectives, improving solution quality, and fostering critical thinking (Gracia, 2003).
+It involves analytical reflection and active discussion(Gracia, 2003), both of which have been explored separately in research on AI-assisted decision making.
+To stimulate analytical thinking, different interventions have been designed to encourage deeper engagement with System 2 thinking (Kahneman, 2011), such as “cognitive forcing” techniques that prompt more deliberation.
+Examples include asking individuals to make independent predictions before receiving AI input (Buçinca et al., 2021; Park et al., 2019; Rastogi et al., 2022) or using “slow algorithms” to reduce reliance on AI.
+Additionally, providing AI explanations without concrete recommendations (Gajos and Mamykina, 2022) and using AI-framed questioning (Danry et al., 2023) have been shown to enhance critical thinking.
+However, these approaches may lead to under-reliance and may not fully address differences between human and AI perspectives.
+Some studies have explored human-AI dialogues in cooperative games ((FAIR)† et al., 2022; Kramár et al., 2022), but these were not tailored to decision-making tasks.
+Recent work has begun integrating discussions into AI-assisted decision-making.
+For example, Zheng et al.
+(Zheng et al., 2023) included AI in group decisions for student essay evaluations, though these efforts often rely on Wizard of Oz setups rather than purpose-built AI systems.
+Similarly, Chiang et al.
+(Chiang et al., 2023) studied collaboration between AI and two humans in recidivism risk assessment but limited the AI’s role to providing suggestions without active discussion.
+Other efforts, such as Zhang et al.
+(Zhang et al., 2023), used AI models as tools to facilitate deliberation among organizations but did not address direct deliberation between humans and AI.
+Perhaps the most relevant work is by Slack et al.
+(Slack et al., 2022), who explored dialogue-based AI explanations to handle follow-up user questions and improve understanding.
+However, the key difference is that we focus on deliberation design and propose Deliberative AI which can not only “explain to users” but also actively engage users in the deliberative discussions by “asking or challenging” the users, aiming to promote people’s critical thinking.
+Previous work has touched upon the idea of having AI serve deliberation among humans, but to our knowledge, no research has directly facilitated deliberation between humans and AI.
+Our work takes an initial step toward designing and evaluating human-AI deliberation in decision-making contexts, offering valuable insights into integrating deliberation into AI-assisted decision-making processes.
+3. Deliberative AI for Human-AI Deliberation
+This section introduces deliberation into the decision-making process, focusing on weighing the evidence.
+Building on the Weight of Evidence (WoE) framework, we propose a Human-AI Deliberation architecture and present the design considerations and architecture of Deliberative AI, an AI assistant capable of engaging in deliberation with humans.
+3.1. Integrating Deliberation into Decision-Making
+Decision-making involves selecting the best choice from a range of alternatives to achieve a desired outcome (Eisenfuhr, 2011).
+The process is typically summarized in seven steps [^1] Seven-step decision making: Step 1: Identify the problem, Step 2: Collect information, Step 3: Identify the alternatives, Step 4: Weigh the evidence, Step 5: Choose from the alternatives, Step 6: Implement action, and Step 7: Evaluate the results., with weighing the evidence being a critical step due to its direct influence on subsequent decision outcomes (U Mass, [n.d.]; Lunenburg, 2010).
+We propose conducting Human-AI Deliberation during this phase because disagreements often arise between human perspectives and AI suggestions in this phase and these human-AI disagreements can lead to conflicting opinions and divergent outcomes in later steps (Lai et al., 2021).
+To structure the deliberation process, we break down decision-making problems and human-AI thoughts into four components (Figure 2 (a)): (1) Decision: The overall choice to be made for a specific problem; (2) Dimension: An aspect considered when forming the overall decision.
+In tabular datasets commonly used in decision-making (Ghai et al., 2021; Wang and Yin, 2021), dimensions often correspond to attributes or sets of related attributes, such as academic excellence or research ability in a graduate admission task; (3) Opinion on a dimension: The assessment of a dimension’s impact on the overall decision (e.g., academic excellence contributes +50% to admission probability); (4) Evidence: The foundation for forming an opinion.
+For humans, evidence may include facts, heuristics, or personal experiences (Simon, 1990, 1997).
+For AI, evidence is often rooted in the information encoded in its training data.
+We propose deliberation at the dimension level, focusing on how each dimension supports or opposes the overall decision.
+Both humans and AI must substantiate their opinions with evidence and assess its credibility and probative value (Baumann et al., 2014).
+This process generates the Weight of Evidence (WoE) (Cartwright and Stegenga, 2011), which quantifies evidence significance/importance relative to alternatives (Peirce, 2014; Good, 1950; Gough, 2007).
+WoE is widely used in decision-making for its intuitive meaning and practical implementation (Saaty, 2008; Gough, 2007).
+In the rest of this paper, we will use “opinion” and “WoE” interchangeably.
 Figure 2.
-The framework for Human-AI Deliberation.
+The architecture of Human-AI Deliberation.
 (A) Illustrates the Weight of Evidence (WoE) concept in decision-making, showcasing how decision-makers assess evidence across dimensions to shape opinions and arrive at a final decision.
@@
 […]
-3.2. Human-AI Deliberation: A Framework
-Deliberation is beneficial for enhancing decision-making by integrating diverse perspectives, improving the quality of solutions by considering various viewpoints, fostering critical thinking, and increasing the likelihood of reaching informed and well-thought-out conclusions (Gracia, 2003).
-Although deliberation is commonly used in democracy and political settings (Black et al., 2010; Habermas, 2005; Thompson, 2008), its idea can be applied to other decision-making tasks such as clinical diagnosis (Schaekermann et al., 2019; Inguaggiato et al., 2019; Preisz, 2019), criminal justice (Devine et al., 2001; Van Berkel et al., 2019), etc.
-Building upon the WoE decision-making approach, we propose Human-AI Deliberation, a framework to stimulate deliberative processes involving both humans and AI (Figure 2).
-This framework comprises the following essential activities:
+3.2. An Architecture of Human-AI Deliberation
+Based on the WoE-centered decision-making approach, we propose Human-AI Deliberation, an architecture to stimulate deliberative processes between humans and AI (Figure 2).
+This architecture comprises the following essential activities:
   • Elicitation of Thoughts: Human and AI start with articulating their dimension-level perspectives on the decision problem.
@@
 First, AI thought elicitation demands a good balance between human information needs and interpretability (Mishra and Rzeszotarski, 2021; Abdul et al., 2020).
-Prior research suggested that certain mechanisms AI applies to make inferences and predictions, although informative, are not comprehensible to humans (Miller, 2019).
-Hence, the design of AI’s “thought” explanations should be critical for this activity.
 Second, human thought elicitation, while encouraging thoughtful reasoning (Milkman et al., 2009; Buçinca et al., 2021), can impose a potential workload.
 It thus demands suitable, friendly interface designs.
-  • Alignment of Human-AI Thoughts: As human’s and AI’s viewpoints and the process they form those viewpoints may diverge (Kaufman and Kirsh, 2022; Miller, 2019; Holzinger et al., 2019), this activity is tasked with establishing a common language for the two parties to compare their WoE and determine the extent of discrepancy, if any.
+  • Alignment of Human-AI Thoughts: As human’s and AI’s viewpoints and the process they form those viewpoints may diverge (Kaufman and Kirsh, 2022; Miller, 2019; Holzinger et al., 2019), this activity is tasked with establishing a common language for the two parties to compare their WoE and determine the extent of discrepancy.
 Proper assessment and presentation of human-AI WoE differences can help effectively navigate humans’ attention and efforts in the subsequent activities (Boggust et al., 2022).
-  • Discussion: This activity features constructive discussions that enable humans and AI to substantiate their opinions via communication, clarifying choices of evidence, and the basis for assigned weights.
-It has the potential benefits of cultivating critical thinking, mitigating biases, and raising awareness of differences between multiple parties (Milkman et al., 2009; Larson et al., 1994; Hirokawa, 1985).
-This activity has a huge design space and usually needs to be tailored to specific decision tasks.
-It entails considering various discussion-related factors, including content, style, leadership (who initiates and leads the conversation), duration, and so on.
-A typical solution is transferring human-human discussion practices (Milkman et al., 2009; Larson et al., 1994; Hirokawa, 1985) to the human-AI discussion contexts.
+  • Discussion: This activity fosters constructive discussions where humans and AI substantiate their opinions, clarify evidence choices, and explain weight assignments.
+It promotes critical thinking, reduces biases, and highlights differences between parties (Milkman et al., 2009; Larson et al., 1994; Hirokawa, 1985).
+With a broad design space, it must be tailored to specific decision tasks, considering factors like content, style, leadership (who initiates and leads the conversation), and duration.
+A potential solution is adapting human-human discussion (Milkman et al., 2009; Larson et al., 1994; Hirokawa, 1985) to the human-AI discussion contexts.
   • Update of Thoughts: In-depth discussions may expose potential flaws and conflicts in the original decisions as humans and AI are both imperfect (Bansal et al., 2021).
@@
 For humans, the interface should possess the flexibility for them to change their WoE.
-In summary, the proposed Human-AI Deliberation framework consists of four interlinked activities and requires proper designs (the potential design space is illustrated in the dashed box in Figure 2 (b)) of a decision-making interface and an AI with the ability to deliberate with humans.
-Since the interface design is task-dependent, we only introduce the design of Deliberative AI in the next subsection.
-Notably, not all activities in this framework are necessarily required for Human-AI Deliberation.
-Furthermore, we do not intend to exhaust all possible forms of Human-AI Deliberation.
-Researchers are encouraged to extend this framework or propose plausible alternatives.
+In summary, the proposed Human-AI Deliberation architecture includes four interlinked activities and requires appropriate designs for both the decision-making interface and the AI (as shown in the dashed box in Figure 2 (b)).
+Since interface design is task-dependent, we focus on the design of Deliberative AI in the next subsection.
 3.3. Deliberative AI: Design Considerations and Overall Structure
-Inspired by human-human deliberation practices and measurements, we derive a set of design considerations for a Deliberative AI.
-Then, we introduce its overall architecture proposed in accordance with these considerations.
-According to deliberative theories (Black et al., 2010; Habermas, 2005; Thompson, 2008) and practices (Lord and Tamvaki, 2013; Steiner et al., 2005), Discourse Quality Index (DQI) (Steenbergen et al., 2003) and its improved versions (Bächtiger et al., 2022, 2009) are the most comprehensive and widely used guideline for assessing human deliberation.
-We adapt DQI to our AI-assisted decision-making context and summarize the following design considerations (DCs):
+Based on deliberative theories (Black et al., 2010; Habermas, 2005; Thompson, 2008) and practices (Lord and Tamvaki, 2013; Steiner et al., 2005), the Discourse Quality Index (DQI) (Steenbergen et al., 2003) and its improved versions (Bächtiger et al., 2022, 2009) provide a comprehensive framework for assessing human deliberation.
+We adapt DQI to AI-assisted decision-making and derive the following design considerations (DCs):
   • DC 1.
@@
 Figure 3.
-The Architecture of Deliberative AI.
-Our design integrates both a domain-specific model and a Large Language Model, enabling the AI to engage in natural communication with humans while also harnessing domain knowledge derived from the specialized model.
+The architecture of Deliberative AI which integrates a domain-specific model and a Large Language Model, enabling the AI to engage in natural communication with humans while also harnessing domain knowledge derived from the specialized model.
 […]
@@
   • The Knowledge layer comprises a domain-specific model and training data, providing both domain-specific knowledge and data-derived insights.
-In Sec. 4, we will empirically explore a subset of these dimensions with a specific instantiation of the proposed Human-AI Deliberation framework and Deliberative AI in a graduate admission task.
-4. Instantiating the Framework: Graduate Admission Prediction
+In the next section, we will provide a detailed description of how we implemented Human-AI Deliberation and Deliberative AI architecture in the context of a graduate admissions task.
+4. Instantiating the Architecture: Graduate Admission Prediction
 4.1. Task, Dataset and AI Model
-We choose to use graduate admission as an illustrative task to demonstrate how to instantiate the proposed Human-AI Deliberation framework.
+We choose to use graduate admission as an illustrative task to demonstrate how to instantiate the proposed Human-AI Deliberation architecture.
 In this task, participants decide on admitting or rejecting applicants to a U.S.
@@
 The task samples used in the study were selected from the test set.
-It should be noted that in order to explore the effect of Human-AI Deliberation, we specially selected cases from the test data that are challenging for both humans and AI, so that the performance of AI is around 50%.
-But this does not mean that the AI model used in our study is an unrealistically low-performance AI, just that we focus on difficult task cases.
-4.2. Implementation of Human-AI Deliberation
-4.2.1. Human-AI Thought Representation and Alignment
-In our Human-AI Deliberation framework, the initial step involves both humans and AI externalizing their thoughts.
-We employ feature contribution (Liao and Varshney, 2021; Liao et al., 2020) to represent their weight of evidence (WoE) along each decision dimension (Alvarez-Melis et al., 2021).
+4.2. Human-AI Thought Representation and Alignment
+In our Human-AI Deliberation architecture, the initial step involves both humans and AI externalizing their thoughts.
+We employ feature contribution (Liao and Varshney, 2021; Liao et al., 2020) to represent their weight of evidence (WoE) along each dimension (Alvarez-Melis et al., 2021).
 Feature contribution is represented by contribution scores indicating the positive or negative influence of each feature $x_{i}$ on the final prediction $y$.
 In graduate admission tasks, we treat each attribute in an applicant’s profile as a dimension, and feature contribution requires humans and AI to assess the influence of each dimension on the final decision.
-It provides a common ground for AI and humans to express, compare, and initiate discussions about their thoughts at the feature level (see Table 1).
-Table 1. The mapping between AI “thoughts” and human thoughts in the Weight of Evidence framework.
-Component AI “thoughts” Human thoughts
-Decision Accept or reject one applicant Accept or reject one applicant
-Dimension Feature/Attribute (e.g., GPA) Feature/Attribute (e.g., GPA)
-Opinion A feature’s contribution (e.g., GPA: + 13%) Weight of evidence (e.g., GPA: -5%)
-Evidence Training data Data or other knowledge held by people
-For the human side, WoE can be interpreted as the influence of an attribute on the total chance of an applicant being admitted.
-For the AI side, we used SHAP (SHapley Additive exPlanations) (Lundberg and Lee, 2017), a widely-used explainable AI algorithm, to generate AI feature contributions.
+It provides a common ground for AI and humans to express, compare, and initiate discussions.
+For the human side, Weight of Evidence (WoE) represents humans’ perceived influence of an attribute on the overall likelihood of an applicant being admitted.
+On the AI side, we utilized SHAP (SHapley Additive exPlanations) (Lundberg and Lee, 2017), a widely-used posthoc explainable AI algorithm, to assess feature contribution/importance.
 SHAP values indicate both the direction and strength of a feature’s impact on predictions.
-SHAP offers two key advantages: it captures feature interactions, mirroring human decision-making, and its additive nature resembles how humans combine evidence weights for or against options (Baumann et al., 2014).
-However, when applied in the context of Human-AI Deliberation, two limitations arise: (1) Raw SHAP values can be challenging to interpret directly.
-To better convey feature influence to non-expert humans, we converted SHAP values into probabilities by adopting a regression model and transforming the four-category label range into a 0-100% range (in this way, the generated SHAP value can be directly mapped to 0-100%)  [^2] Other conversion methods include using the TreeExplainer from the SHAP library (and setting model_output=“probability”) to generate the probabilistic SHAP value.
-Because the admission prediction used in this paper can be naturally converted into a probability problem, we adopted more direct category label-probability mapping.
-(2) SHAP values only tell users how important a feature is, but not why it is important.
-To bridge this gap, we propose generating “meta-explanations” that use Knowledge Extractor (Sec.
-3.3) to extract evidence (e.g., data patterns) from training data for each dimension, enhancing AI’s transparency during deliberation.
-4.2.2. Implementation of Deliberative AI
+SHAP offers two key advantages: First, it captures feature interactions (e.g., how multiple features jointly influence the final decision), aligning with human decision-making processes.
+Second, its additive nature mirrors how humans combine evidence for or against options (Baumann et al., 2014).
+In essence, SHAP allows feature contribution/importance to be linearly aggregate to match the model’s actual prediction no matter the AI model is linear or a complex non-linear neural network.
+However, when applied to Human-AI Deliberation, two limitations of SHAP values become evident: (1) Interpretability of Raw SHAP Values: Raw SHAP values (e.g., 2.15) can be difficult for non-experts to interpret directly.
+To address this, we converted SHAP values into probabilities by using a regression model, mapping the four-category label range to a 0-100% scale.
+This conversion allows SHAP values to be directly interpreted as probabilities.
+(2) Explanation of Feature Importance: SHAP values indicate the importance of features but do not explain why a feature is important.
+To address this, we propose generating “meta-explanations” using the Knowledge Extractor (Sec.
+3.3) to extract relevant evidence (e.g., data patterns) from the training data for each dimension.
+This approach aims to enhance the transparency of AI during the deliberation process.
 Next, we describe how we implemented each component of Deliberative AI.
-I. Communication Layer. This layer serves as a vital bridge between humans and the DS-model, facilitating effective communication by comprehending human inputs and crafting relevant responses.
+4.3. Implementation of Deliberative AI
+4.3.1. I. Communication Layer.
+This layer serves as a vital bridge between humans and the DS-model, facilitating effective communication by comprehending human inputs and crafting relevant responses.
+Note that all components in this layer are based on our designed prompts, which enable LLMs to play different roles, rather than incorporating additional predictive models.
 I-1.
 Intention Analyzer.
-We harnessed the language capabilities of LLMs [^3] In this study, we use GPT-3.5-turbo as the language model, which has sufficient conversational power and is freely accessible (relative to GPT-4).
-We will use “LLM” throughout this paper for consistency.
-to discern human intentions and targeted dimensions in discussion.
+We harnessed the language capabilities of LLMs [^3] During our experiment (conducted in August 2023), we used the GPT-3.5 model.
+For consistency, we will refer to it as “LLM” throughout this paper.
+to identify human intentions and targeted dimensions in discussion.
 To formulate effective prompts for intention analysis, we conducted a pilot study to gather common dialogues around graduate admission decisions, including questions, arguments, critiques, and challenges.
-In the pilot study, we developed a preliminary version Deliberative AI (with imperfect deliberative discussion capability) to carry out conversations with 30 participants from Prolific444www.prolific.co.
-Each participant expressed their opinions on various dimensions of applicant profiles and engaged in discussions with the AI, particularly focusing on conflicting opinions.
-This yielded 226 human deliberative statements.
-To extract diverse intentions from these statements, two authors conducted qualitative coding using thematic analysis (Hsieh and Shannon, 2005), and the results are summarized in Table 2.
-We then iteratively refined LLM prompts based on the collected data and built an ”Intention Analyzer” with a 96% accuracy in identifying themes of participant statements.
+In the pilot study, we developed a preliminary version Deliberative AI (with basic deliberative discussion capability) to carry out conversations with 30 participants from Prolific444www.prolific.co.
+Each participant expressed their opinions on various dimensions of applicant profiles and engaged in discussions with the AI.
+We collected 226 human deliberative statements.
+To extract various intentions from these statements, two authors performed qualitative coding using thematic analysis (Hsieh and Shannon, 2005), and the results are summarized in Table 1.
+We then iteratively refined LLM prompts based on the collected data and built an “Intention Analyzer” with a 96% accuracy in identifying themes of participant statements.
 Specific prompts are available in the supplementary materials.
@@
 Deliberation Facilitator.
-This component addresses DC2 (Justification Rationality) and DC5 (Respect and Agreement), as discussed in Sec 3.3, by designing corresponding LLM prompts.
+This component addresses DC2 (Justification Rationality) and DC5 (Respect and Agreement) by designing corresponding LLM prompts.
 In particular, we instruct LLM to (1) Demonstrate a nuanced understanding of the human’s statement; (2) analyze the specific content of the person’s statement; and (3) provide a thoughtful and critical response.
@@
 Argument Evaluator.
-The main function of this component is to assess the strength of a person’s statement, which informs updates to AI opinions.
+The main function of this component is to assess the strength of a person’s statement/argument, which later informs updates to AI opinions.
 Drawing from established theories in human argumentation evaluation (Harris, 1997; Van Eemeren et al., 2002; Van Eemeren and Henkemans, 2016), we devised a comprehensive scoring mechanism with nine key items: Clarity, Relevance, Evidence, Logic, Consistency, Counterarguments, Depth, Credibility, and Alignment.
@@
 We then average and scale the scores to obtain the overall human argument strength $S_{Human}$ (from 0 to 1; 0: weakest, 1: strongest).
+We conducted a pilot study to evaluate the reliability of the LLM in scoring human arguments, using data collected from our Intention Analyzer pilot study (Sec.
+4.3.1 I-1).
+Two authors independently scored the arguments using predefined schemas, resolving disagreements to reach consensus.
+We then calculated Cohen’s Kappa ($\kappa$) to measure agreement between the LLM and human scores.
+The resulting $\kappa$ value of 0.78 indicates substantial agreement, suggesting that the LLM provides reliable annotations with minimal disagreement in the context of our graduate admission task.
 Additional details, including scoring schemas and prompts, can be found in the supplementary materials.
-In summary, the communication layer can engage in general interactions with humans.
-To imbue it with specific model opinions and knowledge, we require a control layer to mediate between the LLM and the DS-model.
-II. Control Layer. This layer manages the querying and extraction of specific DS-model opinions and knowledge while controlling the entire conversation flow.
+In summary, the communication layer enables general interactions with humans. To integrate it with the DS-model’s predictions and knowledge, a control layer is required to connect the two.
+4.3.2. II. Control Layer.
+This layer manages the querying and extraction of specific DS-model opinions and knowledge while controlling the entire conversation flow.
 II-1.
@@
 We established a mapping between the recognized intent type and the query function and called different query functions based on the recognized intent type.
+In practice, for a decision-making task, an existing dataset is typically available for training the AI model.
+This same dataset can be used to extract data patterns based on task-specific features, such as the percentile of a single feature value or the combination of multiple features.
+We recommend that researchers interested in applying this approach to other tasks first identify the data patterns that users are likely to find relevant, and then design the corresponding extraction functions.
 Below is a brief overview of the designed functions corresponding to different human intent types.
 Please refer to the supplementary materials for detailed codes and examples.
-  • Distribution/Level of an attribute’s value: Function get_distribution (attribute) calculates attribute value percentiles within the applicant pool, along with contextual comparisons (with minimum, maximum, quartiles, mean, and median).
-  • Overall attribute importance: Function get_global_feature_importance (attribute) returns global importance.
-Function get_correlation (attribute) provides Pearson correlation.
-Function get_influence_on_admission_chance (attribute) calculates admission chance changes for varying attribute values.
-  • Contribution of an attribute: Function get_current_value_influence (attribute) calculates admission chance differences when an attribute is randomized.
-  • Contrastive Evaluation: Function get_contrastive_admission_chance (attribute, contrastive) computes admission chance differences with a contrastive value.
-  • Holistic review of multiple attributes: Function get_holistic_analysis (attributes, fixed_attributes) evaluates attribute impacts in specific scenarios, e.g., GPA percentile in top-ranked schools.
+  • Distribution/Level of an attribute’s value:
+    • Function get_distribution(attr_val) calculates attribute value percentiles within the applicant pool, along with contextual comparisons (with minimum, maximum, quartiles, mean, and median).
+  • Overall attribute importance:
+    • Function get_global_feature_importance(attr) returns global importance.
+    • Function get_correlation(attr) provides Pearson correlation between the selected attribute and the admission chance.
+    • Function get_influence_on_admission_chance(attr) calculates admission chance changes for varying attribute values.
+  • Contribution of an attribute:
+    • Function get_current_value_influence(attr) calculates admission chance differences when the value of a selected attribute is randomized.
+  • Contrastive Evaluation:
+    • Function get_contrastive(attr, contrast) computes admission chance differences compared to a contrastive value.
+  • Holistic review of multiple attributes:
+    • Function get_holistic_analysis(attr, fix_attr) evaluates the impacts of attribute(s) in specific scenarios, e.g., the impact of the GPA percentile in [top-ranked universities].
 II-3.
@@
 Opinion Update Controller: We updated the AI’s opinions by taking into consideration: (1) the current opinions of both the human ($O_{Human}$) and the AI ($O_{AI}$) on the discussed attribute, (2) the strength of the human’s argument ($S_{Human}$, see Argument Evaluator), and (3) the AI’s uncertainty ($U_{AI}$) measured and calibrated via Uncertainty Quantification 360 toolbox (Ghosh et al., 2021) (the uncertainty ranges from 0 to 1: the closer to 1, the more uncertain AI’s prediction is).
-We propose the following formula to update the AI’s opinions ($\hat{O}_{AI}$) on an attribute based on these factors, inspired by result aggregation in crowd intelligence (Lyon and Pacuit, 2013; Galton, 1907):
+We propose the following formula to update the AI’s opinions ($\hat{O}_{AI}$) on an attribute based on these factors, inspired by result aggregation in crowd intelligence (Lyon and Pacuit, 2013; Galton, 1907).
+It is important to clarify that the term “update” here does not refer to modifications to the domain-specific model itself, such as retraining or fine-tuning.
+Instead, it pertains to the adjustment of the AI’s expressed viewpoints regarding the current task case.
+Notably, LLM is not directly involved in updating the domain-specific model’s viewpoints.
+Rather, LLM generates a strength score for the human’s arguments, with the update being performed using Eq.
+1.
 (1) $\hat{O}_{AI}=\frac{1-U_{AI}}{1-U_{AI}+S_{Human}}\cdot O_{AI}+\frac{S_{Human}}{1-U_{AI}+S_{Human}}\cdot O_{Human},$
-III. Knowledge Layer
-This layer primarily consists of the DS-Model and the training dataset.
-The DS-Model offers comprehensive perspectives and facilitates real-time decision predictions.
-The training dataset offers necessary information (e.g., data distributions and patterns) for the Knowledge Extractor to perform real-time calculations and queries.
-4.2.3. An Example
-Figure 5 provides the details of a conversation between a human and a Deliberative AI discussing how an applicant’s GPA affects admissions chances. Here’s a step-by-step breakdown:
-  1. (1)
-    The user inputs GPA-related arguments in the dialogue interface.
-  2. (2)
-    The system packages the user’s input as a prompt for the Intention Analyzer in the Communication Layer.
-  3. (3)
-    The Intention Analyzer recognizes attributes and intentions and saves in JSON format, then forwards it to the Knowledge Extractor in the Control Layer.
-  4. (4)
-    The Knowledge Extractor generates a query function and fetches statistical results from the DS-Model and training data.
-  5. (5)
-    The statistical results are transmitted to the Regulator.
-  6. (6)
-    Regulator crafts a constraint prompt ensuring consistency between the LLM’s output and the DS-Model’s prediction, feeding it to the LLM-based Deliberation Facilitator.
-  7. (7)
-    The Deliberation Facilitator generates responses to the user’s initial arguments.
-Overall, in this framework, LLM is used for language understanding and generation.
+4.3.3. III. Knowledge Layer
+This layer comprises the DS-Model and the training dataset.
+The DS-Model provides overall predictions and opinions on each dimension, while the training dataset supplies essential information (e.g., data distributions and patterns) for the Knowledge Extractor to perform real-time calculations and queries.
+Overall, in this architecture, LLM is used for language understanding and generation.
 The opinions and evidence used by LLM are retrieved in real time from the DS-Model and training data through our logic code (like retrieval augmented generation (Lewis et al., 2020)).
 In this way, the LLM is used in a controllable and responsible manner, minimizing the potential hallucination.
-Figure 5.
-An illustration of how Deliberative AI processes humans’ inputs and how it generates outputs.
-The prompts used are simplified in this figure for illustration purposes.
-(for the complete prompts, please check our supplementary materials)
-[…]
+We provide an example of how data is processed in Deliberative AI in the Appendix.
 4.4. Interface Design
@@
 Each slide bar shows three line indicators: a white line representing the average admission probability of all applicants, a green line showing the initial predictions made by humans/AI, and a yellow line denoting the updated prediction by humans/AI (only shown after an update is made).
-      The bottom part of this region allows both humans and AI to express specific opinions on each decision dimension (i.e., applicant attribute).
-A simplified profile is displayed in the middle (Figure 6-5) to minimize users’ need to switch attention.
-Each dimension is accompanied by a status indicator, denoting whether it has been discussed (green), is currently being discussed (orange), or is yet to be discussed (gray).
-Separate “concrete opinion” slide bars are presented next to each attribute (Figure 6-4 and Figure 6-6) for AI and humans to indicate their dimension-level opinions.
+      The bottom part of this region enables both humans and AI to express opinions on each decision dimension (i.e., applicant attribute).
+A simplified profile in the middle (Figure 5-5) reduces attention shifts.
+Status indicators show if a dimension has been discussed (green), is being discussed (orange), or is yet to be discussed (gray).
+Separate “concrete opinion” sliders (Figure 5-4 and Figure 5-6) allow AI and humans to share dimension-level opinions.
       Each dimension’s slide bar starts in a central position (0% contribution).
@@
 5. Exploratory User Study
-To get an initial understanding of the impact of the Human-AI Deliberation mechanism on AI-assisted decision-making, we conducted a mixed-methods study in the context of graduate admission decisions.
-We call this study exploratory because we did not seek to specifically prove whether Human-AI Deliberation is effective for actual graduate admissions.
-In other words, graduate admission is NOT our focus.
-Instead, we just use graduate admission as an illustrative task to explore the potential effects of Human-AI Deliberation on AI-assisted decision-making.
+To gain an initial understanding of the impact of the Human-AI Deliberation, we conducted a mixed-methods study within the context of graduate admissions.
+This study is termed exploratory because our objective was not to assess the effectiveness of Human-AI Deliberation specifically for graduate admissions.
+Rather, we used graduate admissions as an illustrative task to explore the potential effects of Human-AI Deliberation on AI-assisted decision-making.
 5.1. Task Setup
@@
 These cases included one for the tutorial and four for the main tasks.
-Among the main tasks, two represented US applicants (US cases), and the other two represented Asian applicants (Asia cases).
-To keep the Deliberative AI discussions engaging and avoid repetition, we carefully chose tasks with diverse attributes.
-It’s worth noting that the AI model is biased toward US applicants due to training data biases.
-This bias results in a “+11%” feature contribution for “Country belongs to the US“ and a “-4%” feature contribution for “Country belongs to Asia”.
-So it can be interesting to test whether participants will discover such biases when collaborating with the AI.
+To investigate human-AI deliberation, we selected cases from a pilot study where human and AI opinions conflicted, often involving ambiguous applicant profiles near the admission borderline.
+As a result, predicting the admission outcome for these cases was challenging for both humans and AI, leading to performance levels between 50% and 60% for both.
+However, this does not imply that the AI model used in our study is of unrealistically low performance; rather, we focused on ambiguous task cases that require conflict resolution.
 5.2. Conditions
@@
   • Human Alone: Participants need to make predictions independently without any AI assistance.
-5.3. Procedure and Participants
-5.3.1. Procedure
-We conducted a between-subjects study.
-After obtaining consent, participants completed a background questionnaire gathering demographic data and assessing their AI expertise.
-Participants then went to an interactive tutorial, practiced with one example task, and received distribution and summary statistics for each attribute of the applicant’s profile.
-Following the tutorial, qualification questions were asked to check participants’ understanding of the task, with only those answering all questions correctly proceeding to the main task.
-The main task involved tackling four graduate admission task cases, including providing predictions/opinions, receiving AI suggestions/opinions, engaging in discussions with AI, and confirming final predictions.
+5.3. Procedure
+With our institutional IRB approval, we conducted a between-subjects study.
+After obtaining consent, we had participants complete a background questionnaire to gather demographic data and assess their AI expertise.
+We then introduced the study, explained the task, workflow, and AI’s functions, including its ability to update opinions, without delving into the specifics of the adjustment mechanism.
+This approach reflects real-world scenarios, where non-technical users focus more on functionality than technical details.
+Participants then engaged in an interactive tutorial, practiced with one example task, and received distribution and summary statistics for each attribute of the applicant’s profile.
+After the tutorial, we asked qualification questions to check participants’ understanding of the task, allowing only those who answered all questions correctly to proceed to the main task.
+In the main task, participants worked on four graduate admission task cases, which were presented randomly.
 Finally, we collected participants’ perceptions, experiences, and feedback on the AI system and the discussion process in the exit survey.
-5.3.2. Participants
-Before participant recruitment, we conducted a power analysis to determine the required sample size for using G*Power (Faul et al., 2009).
-We specified a default effect size $f$=0.25 (indicating a moderate effect), a significance threshold $\alpha$=0.05, and a statistical power $1-\beta$=0.8.
+5.4. Participants
+We first conducted a power analysis to determine the required sample size for using G*Power (Faul et al., 2009) with a default effect size $f$=0.25 (indicating a moderate effect), a significance threshold $\alpha$=0.05, and a statistical power $1-\beta$=0.8.
 This resulted in a required total sample size of 159 participants for the three conditions.
@@
 After filtering based on attention-check questions, we obtained 153 valid responses (Deliberative AI: 48, XAI: 51, Human Alone: 54).
-Among the final participants, 84 self-reported as male, 67 as female, and 2 as other.
-There were 23 participants aged 24-29, 42 aged 30-39, 33 aged 40-49, 30 aged 50-59, and 25 aged over 59.
-Participants also rated their knowledge of artificial intelligence: 9 had no knowledge, 86 knew basic AI concepts, 50 had experience using AI algorithms, and 8 were AI experts.
+Among the final participants, 84 self-identified as male, 67 as female, and 2 as others.
+The age distribution was as follows: 23 participants aged 24–29, 42 aged 30–39, 33 aged 40–49, 30 aged 50–59, and 25 aged over 59.
+Regarding education, 125 participants held an MA/MSc/MPhil degree, and 28 held a Doctorate (PhD or equivalent).
+Participants also had diverse levels of AI knowledge: 9 reported having no knowledge, 86 were familiar with basic AI concepts, 50 had experience using AI algorithms, and 8 identified as AI experts.
 Participants in the Deliberative AI condition received bonuses based on the actual study length.
@@
 On average, participants earned about \$12 per hour.
-Table 3. Measurements used in our user study. We collected participants’ objective decision and behavior data, subjective questionnaire data, and qualitative open-ended feedback.
+5.5. Measurement
+To answer the aforementioned research questions, we comprehensively measured the effects of human-AI deliberation across four dimensions: task performance, reliance, perceptions of AI, and user experience.
+  • Task Performance. We evaluated decision-making accuracy using Decision Accuracy metrics (Zhang et al., 2020; Bansal et al., 2021).
+  • Reliance.
+Participants’ reliance on AI suggestions was assessed through the Agreement Fraction (Zhang et al., 2020; He et al., 2023) and Switch Fraction (Zhang et al., 2020; He et al., 2023).
+Additionally, the appropriateness of reliance was measured using the Over-reliance Ratio (Wang and Yin, 2021; Visser et al., 2014; Parasuraman and Riley, 1997) and Under-reliance Ratio (Wang and Yin, 2021; Visser et al., 2014; Parasuraman and Riley, 1997).
+  • Perceptions of AI.
+Participants’ perceptions of AI were measured using 7-point Likert scales for Helpfulness (Laugwitz et al., 2008; Cai et al., 2019a; Buçinca et al., 2020), Trustworthiness (Ghai et al., 2021; Buçinca et al., 2021), and Understanding (Wang and Yin, 2021) (1: Strongly disagree; 7: Strongly agree).
+  • User Experience.
+Participants’ Decision Confidence was evaluated using established measures (Miller et al., 2015).
+Given that deliberation requires additional effort, we also assessed Mental Demand (Hart, 2006; Lai et al., 2022; Ghai et al., 2021; Buçinca et al., 2021), Effort (Hart, 2006), Complexity (Buçinca et al., 2021), and Satisfaction (Buçinca et al., 2021; Ghai et al., 2021), all measured on 7-point Likert scales.
+To gain a deeper understanding of participants’ perceptions of both Deliberative AI and the deliberative decision-making process, we also gathered open-ended feedback in the exit survey.
+These questions explored participants’ perceptions of the usefulness of deliberating with AI, the AI’s updates, and their expectations for system improvements.
+A detailed overview of the metrics and questions is provided in Table 2 in the Appendix.
+5.6. Analysis Methods
+We conducted mixed-methods analyses to examine the aforementioned metrics.
+For quantitative analysis, we first performed normality tests (Shapiro-Wilk) and found that the data did not fit the normality assumption.
+Therefore we ran the non-parameter tests.
+Specifically, to compare Deliberative AI and XAI (such as humans’ reliance on AI, and their perceptions of AI), we run Mann-Whitney U test.
+To compare all three conditions (such as task performance, and user experience), we employed Kruskal–Wallis tests with Bonferroni post-hoc correction and we reported adjusted p-values.
+For qualitative analysis, two authors independently coded participants’ open-ended feedback and conversation logs using an inductive thematic analysis approach (Hsieh and Shannon, 2005).
+The final themes emerged through discussions and harmonization over several iterations.
+We also identified representative examples from the source texts for demonstration in this paper.
+6. Results
+In this section, we report our exploratory findings regarding the three research questions: (RQ1) how Human-AI Deliberation affects task performance and human reliance (and reliance appropriateness) on AI, (RQ2) how Human-AI Deliberation affects human perceptions and task experience, and (RQ3) how humans perceive the effectiveness of Human-AI Deliberation and what should be improved.
+6.1. RQ1: How will Human-AI Deliberation affect task performance and humans’ reliance (and its appropriateness) on AI suggestions?
+6.1.1. Decision Accuracy
+As shown in Figure 6, participants in the Deliberative AI condition demonstrated significantly higher decision accuracy ($M$=0.598, $SD$=0.169) compared to those in the XAI condition ($M$=0.524, $SD$=0.16, $p<$0.05).
+This finding suggests that in scenarios where tasks are challenging for both humans and AI—where the individual performance of humans and AI is relatively low—traditional Explainable AI (XAI) may not improve performance and could even have negative effects.
+In contrast, Human-AI Deliberation shows potential to yield positive outcomes, even when the performance of the underlying AI models in these difficult task cases is suboptimal.
+Figure 6.
+Task performance in different conditions.
+It is important to note that we intentionally selected ambiguous task cases that are prone to conflicts between humans and AI and are challenging for both.
+As a result, the accuracy of both humans and AI individually is relatively low.
+The error bars represent 95% confidence interval.
+(*: $p<$0.05)
+[…]
+Figure 7.
+Participants’ reliance and the appropriateness of their reliance.
+(A) Participants’ reliance on AI’s suggestions was measured by agreement fraction and switch fraction.
+(B) The appropriateness of participants’ reliance on AI’s suggestions, including under-reliance (the ratio where participants did not use a correct AI suggestion) and over-reliance (the ratio where participants used an incorrect AI suggestion).
+The error bars represent 95% confidence interval.
+(*: $p<$0.05, ***: $p<$0.001)
+6.1.2. Reliance
+We measured participants’ objective reliance by agreement fraction and switch fraction.
+As shown in Figure 7 (a), participants agreed significantly less with AI’s suggestions in Deliberative AI ($M$=0.57, $SD$=0.24) than in XAI ($M$=0.68, $SD$=0.27, $p<$0.05), and switched significantly less to AI’s predictions in Deliberative AI ($M$=0.23, $SD$=0.35) than in XAI ($M$=0.51, $SD$=0.41, $p<$0.001).
+Combined with participants’ open-ended feedback (Sec.
+6.3), this may be because people invest more in independent thinking in the process of deliberation with AI and realize the problematic aspects of AI’s perspective.
+6.1.3. Appropriateness of Reliance
+We further measured the appropriateness of participants’ reliance on AI’s suggestion by under-reliance and over-reliance (Figure 7 (b)).
+Results show that there is no significant difference between Deliberative AI ($M$=0.32, $SD$=0.28) and XAI ($M$=0.29, $SD$=0.30) in terms of under-reliance.
+While significantly less over-reliance was observed in Deliberative AI ($M$=0.47, $SD$=0.31) than in XAI ($M$=0.65, $SD$=0.33, $p<$0.001), which means that participants had more appropriate reliance on AI when collaborating with our proposed Deliberative AI.
+This result aligns with existing research on cognitive forcing functions (Buçinca et al., 2021), where high cognitive effort reduces over-reliance on AI.
+However, unlike in (Buçinca et al., 2021), where cognitive effort significantly increased under-reliance (i.e., humans under high cognitive effort may blindly ignore AI’s suggestions), our findings reveal no such adverse effect.
+This indicates that our tool mitigates over-reliance not merely by increasing cognitive effort but by fostering meaningful deliberation.
+6.2. RQ2: How will Human-AI Deliberation affect humans’ perceptions of the AI and user experience?
+We measured the effects of different AI conditions on participants’ perceptions and user experience via a 7-point Likert scale (1: strongly disagree, 7: strongly agree).
+Figure 8.
+Participants’ perceptions and task experience.
+(A) Participants’ perceptions of different AI assistant.
+(B) Effects on user experience.
+The error bars represent 95% confidence interval.
+(*: $p<$0.05, **: $p<$0.01)
+6.2.1. Perceptions of AI
+Figure 8 (a) shows participants’ perceptions of the AI model.
+There were no significant differences in perceived helpfulness and understanding between Deliberative AI and XAI.
+However, participants reported significantly less trust in Deliberative AI ($M$=4.47, $SD$=1.68) compared to XAI ($M$=5.52, $SD$=1.27, $p<$0.01), aligning with their reliance behaviors (Sec.
+6.1.2).
+This difference may be attributed to participants identifying more AI flaws through deliberation than by solely observing AI’s explanations, supported by conversation logs analysis (see Sec.
+6.3 for details).
+6.2.2. User experience
+First, we want to see participants’ decision confidence.
+As indicated in Figure 8 (b), participants in XAI reported significantly higher confidence ($M$=6, $SD$=1.10) in their predictions than those in Human Alone ($M$=5.59, $SD$=1.12, $p<$0.05).
+However, from Figure 6 (a) we found that the final accuracy of participants in XAI is even lower than those in Human Alone.
+This indicates that the traditional XAI might lead to humans’ illusionary confidence, which could prevent humans from making optimal decisions.
+Given that the Deliberative AI requires participants to externalize thoughts at a dimension level and engage in deliberative discussions on conflicting opinions, it’s crucial to explore how these activities influence the user experience.
+Results showed no significant difference among the three conditions concerning Mental Demand, Effort, and Perceived System Complexity.
+However, we find participants reported significantly lower Satisfaction in Deliberative AI than in Human Alone.
+This decrease in user experience may be due to the AI exposing more conflicts for humans to resolve.
+As P3 noted, “It’s annoying because I have to try to find evidence to prove that my point of view is correct.” This result suggests that there is a trade-off between encouraging users’ deliberative thinking and optimizing their user experience, in line with findings in previous studies (Buçinca et al., 2021).
+Future work should focus on finding a balance between the benefits of deliberation and maintaining a positive user experience.
+6.3. RQ3: How will humans perceive the effectiveness of the proposed Human-AI Deliberation and what can be improved for future Human-AI Deliberation design?
+In addition to quantitative measures, we aimed to gain a deeper understanding of participants’ perceptions of the helpfulness of the proposed Human-AI Deliberation and the Deliberative AI feature, particularly the opinion-updating aspect.
+We also sought insights to inform future design improvements for Human-AI Deliberation.
+To achieve this, we analyzed participants’ open-ended feedback, supported by their conversation logs.
+Our key findings are summarized in Figure 9.
+Below, we present key insights, with themes highlighted in bold.
+Figure 9. The main results of our thematic analysis of the open-ended questions.
+[…]
+6.3.1. Perceptions of Deliberative AI and the Discussion Process
+Participants offered both positive and critical feedback on Deliberative AI and the deliberation process.
+Of the 48 participants who experienced with Deliberative AI condition, 43 felt that deliberation helped them make more informed decisions.
+Specifically, AI-assisted discussions enabled participants to “identify AI’s limitations” (21/48), “consider different perspectives” (10/48), and “reflect on and correct their own mistakes” (15/48).
+For example, P1 (Male, 32) pointed out the AI’s overreliance on GPA scores while underestimating the role of recommendation letters: ”The AI relies too much on GPA scores but undervalues recommendation letters.
+It didn’t provide convincing justifications, so I couldn’t rely on its opinion for these factors.” P19 (Male, 42) noted how the deliberation process encouraged self-reflection: ”The AI made me question what I believed to be sufficient reasoning, especially given the data.” These reflections are consistent with our analysis of participants’ conversation data with Deliberative AI, where participants expressed doubts to the AI in 32% of dialogues, acknowledged its arguments in 17% of dialogues, and engaged in self-reflection and correction in 15% of dialogues.
+Moreover, 18 participants found that discussing with the AI introduced “new knowledge, insights, and perspectives”.
+For instance, P5 (Male, 45) commented: ”The AI provided information I didn’t know, like percentiles and how similar stats influenced past decisions, which I found extremely helpful.”
+Nine participants said that deliberation helped them recognize biases.
+For example P35 (Female, 29) mentioned: ”It made me realize the AI had inherent biases, which prompted me to pause and reflect.” This aligns with findings from the participants’ conversation data with AI, where many participants (15/48) identified biases in the AI’s reasoning.
+For example, the AI exhibited bias by assigning more importance to U.S.
+applicants while downplaying those from Asia.
+P3 (Male, 42) questioned: ”Why does their country matter?
+Penalizing someone for their nationality seems biased.”
+Furthermore, 15 participants appreciated that deliberation balanced the AI’s objective data with human subjective judgment.” For instance, P2 (Male, 42) said: ”It provides an opportunity to consider multiple perspectives and statistical data, resulting in more balanced decision-making.”
+Despite the benefits, participants also identified limitations in the deliberative process.
+Five participants felt that their sense of agency was reduced as the AI prompted them to think, rather than passively waiting for their input.
+Four participants found the discussion is mentally demanding and criticized the AI’s verbose responses, reflecting the results on user experience (see Sec.
+6.2.2).
+Regarding the limitations of Deliberative AI, two main concerns emerged: that “Deliberative AI relied too heavily on data over logic” (15 participants) and “failed to consider human reasoning” (10 participants).
+This feedback aligns with our analysis, where participants integrated personal experiences and logic into their decisions, while the AI lacked this depth.
+For example, participants considered factors such as: ”Business programs are highly competitive, requiring a higher GPA.” (P17, Female, 60) and ”A strong SOP reflects a deep understanding of the project and strong commitment.” (P32, Male, 28)
+Additional concerns included the AI’s insufficient ability to faithfully remember previous discussions (five participants), failure to consider the broader context (three participants), and lack of transparency (three participants).
+6.3.2. Perceptions of Deliberative AI’s Updates
+Participants generally appreciated the dynamic updating feature.
+25 out of 48 participants felt that the AI updates improved the user experience.
+They described the experience as more interactive, comparing it to “interacting with a real person” (6 participants), where their opinions felt “heard and respected” (19 participants).
+Moreover, these updates made the discussion “more engaging” (8 participants) and the decision-making process “more collaborative” (5 participants).
+However, perceptions of AI’s updates varied.
+While 11 participants felt the AI adjusted its views too frequently (e.g., “the AI updated too much”), 13 participants expressed uncertainty about “why and how the AI was updating.” For example, P2 (Male, 42) noted: ”I’m not sure if the AI really changed its opinion based on what I said or if it was just programmed to do so in response to my input.” Interestingly, some participants thought the AI was too quick to change its stance, while others felt it wasn’t flexible enough.
+These updates also affected participants’ perceptions of the AI itself.
+Eight participants felt the AI could “incorporate qualitative information,” and five believed the AI was “learning from human knowledge during the discussion,” which could “enhance the quality of decision-making” (3 participants).
+However, three participants found the lack of transparency in the update process led them to “doubt the reliability of the AI’s initial opinions.”
+Although the AI’s dynamic updates mimic human deliberation, making these updates more meaningful requires considering diverse user perceptions and designing the feature more thoughtfully.
+Specifically, the future design of AI opinion updates should draw insights from social science (Liao and Varshney, 2021), addressing user preferences regarding the frequency and magnitude of updates.
+Recall that in the introduction before the experiment, we did not disclose the specific mechanism of AI opinion updates to the users, so this process was not transparent to them.
+Enhancing transparency and users’ understanding of AI update could alleviate their concerns.
+6.3.3. Opportunities for Future Improvements of Deliberative AI
+Participants provided several suggestions for improving Deliberative AI in future versions.
+Fifteen participants emphasized the need for AI to develop a deeper understanding of human thoughts and arguments, calling for more nuanced and context-aware communication.
+Eight participants stressed the importance of AI remembering previous interactions to enable more personalized and coherent conversations, avoiding repetition and improving the AI’s responsiveness to specific perspectives.
+Twelve participants suggested that AI should move beyond statistical reasoning to offer broader, more holistic solutions, while 10 participants advocated for deeper analysis and consideration of broader implications in each context.
+Transparency was also highlighted as crucial, with eight participants asking for clearer explanations about the AI’s changing opinions to foster trust and understanding.
+Regarding the design of human-AI discussions, eight participants preferred a more conversational and human-like interaction, while 11 favored concise AI responses.
+Two participants recommended gradually disclosing the AI’s arguments, and three wanted more user control over steering the discussion.
+Additional suggestions included improving dialogue generation speed and enhancing the display of visual information beyond just text.
+7. Discussion
+In this paper, we introduce the Human-AI Deliberation approach to address two key challenges in AI-assisted decision-making: insufficient analytical engagement with AI recommendations and limited support for resolving human-AI disagreements.
+Traditional interfaces often restrict users to accepting or rejecting AI suggestions as a whole, limiting nuanced understanding and collaboration.
+Our exploratory assessment demonstrates that Human-AI Deliberation improves decision accuracy and fosters appropriate reliance on AI compared to conventional explainable AI systems.
+In this section, we discuss key implications, generalizability, limitations, and directions for future work.
+7.1. Deliberation as a New Paradigm Complementing Existing (X)AI Assistance
+The Human-AI Deliberation approach introduces a conflict-driven discussion model that complements traditional AI assistance.
+Deliberative AI builds on explainable AI (XAI) principles, challenging human perspectives while respecting their agency as decision-makers.
+It enhances AI assistants by fostering deeper analytical thinking, encouraging users to form independent opinions before seeing AI suggestions (as in the Cognitive Forcing Function approach (Buçinca et al., 2021)) and stimulating critical thinking through AI-generated questions (similar to (Danry et al., 2023)).
+Additionally, Human-AI Deliberation involves users in active discussions, improving communication and transparency with AI.
+This engagement helps lead to more informed and nuanced decisions.
+Application Value of Deliberative AI.
+Deliberative AI is well-suited for critical decision-making tasks, such as investment or hiring, where decision quality outweighs the need for speed or convenience.
+While engaging with Deliberative AI may demand more time and effort, decision-makers appreciate its ability to support well-considered outcomes.
+Additionally, Deliberative AI fosters user reflection by encouraging individuals to examine biases in their reasoning and evaluate differences between their perspectives and the AI’s suggestions.
+Participants frequently emphasized this in their qualitative feedback, highlighting its value even in subjective decision contexts.
+Beyond enhancing objective accuracy, Deliberative AI promotes thoughtful consideration and introspection, which are vital for decisions requiring nuanced judgment.
+7.2. Reducing Human Over-Reliance by Exposing AI Mistakes
+Human-AI Deliberation significantly reduces participants’ tendency to over-rely on inaccurate AI suggestions.
+This approach utilizes cognitive forcing theory, which encourages forming independent opinions before viewing AI recommendations.
+This helps counteract anchoring bias—the undue influence of initial AI predictions on subsequent judgments—and fosters more analytical, System 2 thinking  (Kahneman, 2011).
+While our Explainable AI (XAI) baseline also promotes independent opinion formation, Deliberative AI proves more effective by involving deeper deliberative discussions, especially when facing conflicting viewpoints.
+In the Deliberative AI architecture, participants become more aware of AI’s limitations through engagement with conflicting opinions, reducing over-reliance.
+Notably, 31% of conversations involve participants questioning AI’s logic, reflecting a critical evaluation of its insights.
+This approach effectively minimizes over-reliance without increasing under-reliance, as participants adjust their judgments based on deliberative dialogue.
+We recommend that AI-assisted decision-making systems should emphasize transparency by explicitly highlighting potential AI errors rather than merely suggesting that “AI may make errors”.
+7.3. Human-AI Conflict Resolution: Key to Enhancing Collaboration
+We propose that addressing conflicts in decision-making is more beneficial than merely seeking consensus.
+Our approach, Human-AI Deliberation, prioritizes resolving disagreements between humans and AI—an often overlooked aspect in current AI-assisted decision-making.
+Conflicts serve as a lens to uncover underlying flaws and biases, making them essential to effective human-AI collaboration.
+This conflict-centered methodology offers several advantages: it enhances decision-making accuracy by mitigating over-reliance on AI, fosters deeper introspection, and facilitates reconciliation between differing human perspectives and AI-generated recommendations.
+Additionally, addressing conflicts allows for the identification of biases in AI interpretations, thereby promoting fairness in decision-making (Kramer et al., 1990; Hochman et al., 2015).
+However, prioritizing conflict resolution can influence user experience.
+Our experimental results show that while Deliberative AI demonstrates no significant differences from XAI and Human Alone in terms of mental demand, effort, or perceived complexity, it introduces a more complex decision-making process, resulting in lower user satisfaction.
+We identify three primary sources of the system’s burden and propose solutions to mitigate them:
+  • Articulating Opinions Across Dimensions: Currently, users must articulate their views comprehensively across multiple dimensions.
+Future interfaces could streamline this by letting users focus on dimensions they consider most critical or provide approximate opinions in natural language.
+AI could then identify the direction and intensity of human opinions from languages, allowing users to refine these interpretations as needed.
+  • Dialogue Effort: The AI currently initiates discussions on dimensions with significant disagreements.
+Future designs could let users take a more active role or adopt a bidirectional approach to enhance autonomy.
+Alternatively, AI could prompt users to reflect on specific dimensions and acknowledge the completion of reflection with simple actions, such as clicking an “OK” button, reducing dialogue input.
+  • Emotional Challenges in Conflict Resolution: Resolving conflicts with AI can be difficult and reduce satisfaction with Deliberative AI.
+To improve this, interfaces could reframe conflicts as opportunities for reflection and feedback, using Constructive Conflict Theory (Kriesberg, 2007) to emphasize that AI and humans share the same goals.
+Drawing from Positive Psychology (Noor et al., 2008), AI could demonstrate active listening, show appreciation for user viewpoints, and foster a supportive atmosphere to reduce dissatisfaction and encourage deeper engagement.
+7.4. Obstacles to Discussion: Humans and AI Think Differently
+Humans Use Heuristics and Logic, While AI Relies on Data: Integrating LLMs and DS-Models enhances Deliberative AI for dynamic human-AI discussions.
+However, AI’s data-centric approach often clashes with human decision-making, which relies on personal understanding, logic, heuristics, and creativity (Ghattas et al., 2014; Evans, 2002; Doswell, 2004; Simon, 1990, 1997).
+We propose using DS-Models to guide LLMs on data-related discussions, while LLMs handle non-data matters autonomously.
+Despite this, engaging users with strong subjective perspectives remains challenging (Liu et al., 2023).
+Existing Explanation Methods Fall Short: Current XAI methods, such as local feature importance (Liao and Varshney, 2021), inadequately explain specific feature contributions.
+Our approach, which provides data-related insights like distribution and value comparisons, aims to support feature-level discussions.
+Yet, challenges persist: data patterns may not match AI explanations precisely, and users often rely on subjective reasoning (Doswell, 2004; Ghattas et al., 2014).
+As Miller et al.
+note, ”probabilities are not as important as causal links” (Miller, 2019).
+Thus, data-driven insights alone may not align with users’ cognitive processes or replicate human-like conversations.
+Future human-AI deliberation designs should address these challenges by aligning with human intuition and cognitive processes, and by crafting AI explanations that facilitate more effective discussions.
+7.5. Ethical Concerns in Deliberative AI Design
+Responsible Use of LLMs in Assisting Human Decision-Making: LLMs are evolving rapidly but are prone to “hallucinations,” where they generate plausible but incorrect information (Ji et al., 2023).
+Relying solely on LLM-generated opinions without appropriate fine-tuning is irresponsible.
+Even with fine-tuning, LLMs may produce unpredictable responses.
+We advocate for a responsible approach that integrates LLMs with DS-Models, where DS-Models guide LLM responses, positioning LLMs primarily as intermediaries between users and DS-Models.
+Although this approach limits LLM flexibility, it enhances security and control.
+Nonetheless, even with DS-Model guidance, inaccuracies can still occur.
+Researchers should exercise caution and transparency, informing users about the potential for errors and the limitations of LLM-generated information.
+Ethical Issues with AI Opinion Updates: Research indicates users value AI’s responsiveness to their arguments (Zheng et al., 2023).
+Our AI opinion update mechanism, which adjusts AI stances based on user input and prediction uncertainty, aims to reflect this need.
+While users appreciate feeling heard, ethical concerns arise, particularly about accountability.
+A key issue is who should be responsible if the AI, initially correct, updates to an incorrect prediction after discussion.
+Additionally, the AI’s adaptability may create the impression of learning and progress, even if the underlying model remains unchanged, potentially leading to unrealistic expectations.
+It is crucial to design AI systems with transparent updating mechanisms to ensure users understand how updates are made and manage their expectations effectively.
+7.6. On the Generalizability of Human-AI Deliberation
+Task Suitability: Human-AI Deliberation is less suited for repetitive, low-stakes tasks like content moderation (Lai et al., 2022) but is more appropriate for high-stakes, complex tasks such as healthcare (Lee et al., 2021), finance (Green and Chen, 2019), and criminal justice (Chiang et al., 2023).
+Discussion Effort: Discussing every feature, as in our study, may be impractical for tasks with many attributes. Grouping features into broader categories could streamline discussions.
+Opinion Representation and Alignment: We used the Weight of Evidence method to quantify opinions, but this may not be intuitive for all users.
+Future designs could infer opinions from natural language, emotional intensity, or comparisons, or simplify input with rankings or pairwise comparisons (Fürnkranz and Hüllermeier, 2010).
+Data Type Applicability: While designed for tabular data, the architecture can extend to text tasks with LLM’s capability in dealing with textual data.
+Adapting it to image data may require advances in vision-language models (Zhang et al., 2024).
+Using LLM as a Communication Bridge: In Deliberative AI, the LLM acts as a deliberation facilitator, intention analyzer, and argument evaluator, relying mainly on its language understanding capabilities with minimal reliance on its reasoning abilities.
+Through two pilot studies, we validated the LLM’s effectiveness in these roles.
+However, challenges remain regarding its reasoning capabilities (Wei et al., 2022).
+Designers should leverage the LLM’s strengths in communication and avoid overburdening it with complex reasoning tasks.
+7.7. Limitations and Future Work
+The study design has several limitations.
+First, the college admissions task used for illustration does not fully capture real-world admissions processes, which typically involve in-depth discussions about a student’s materials, such as her/his statement of purpose (SOP), background, and overall fit for the department.
+However, our used public dataset quantifies the “strength” of the SOP and recommendation letter as a scale value, which inevitably oversimplifies these nuanced evaluations.
+Additionally, while participants had relevant experience, they were not actual admissions committee members, leading to an expertise gap.
+Furthermore, most participants had experience applying to master’s programs, with only a few familiar with PhD admissions.
+Given the distinct evaluation criteria for these two application types, our findings may not generalize to PhD admissions.
+Future research should assess the approach’s effectiveness in real-world admissions contexts.
+Second, graduate admissions is a subjective task that lacks a definitive ground truth.
+In our study, since the dataset labels were provided by a professional admissions committee, we used decision accuracy and over/under-reliance as objective metrics to assess decision quality “to some extent”.
+The dataset also abstracts subjective elements (e.g., SOP and recommendation letter) into numerical strength values, making the task relatively more objective.
+Nonetheless, future work should explore the effects of Deliberative AI in the context of more objective tasks.
+Third, the number of decision tasks in the study was limited.
+During the pilot, using 8-10 tasks led to participant fatigue and a drop in engagement after completing 3-4 tasks.
+To maintain engagement, we reduced the task count to four, which restricts the generalizability of our results.
+Future studies should conduct long-term evaluations to collect more deliberative decision data.
+Fourth, the independent decision accuracy of both humans and AI was relatively low (50-60%) in our selected task cases, as we intentionally chose tasks prone to conflicts that require discussion.
+The ambiguity in these cases led to suboptimal performance from both.
+However, our study did not address less-ambiguous cases, where differing but firm opinions may still arise.
+We believe Deliberative AI can still help reduce errors in such cases by prompting reflection on biases or overlooked perspectives, especially when humans’ confidence is high.
+Further research is needed to explore different task cases for a more comprehensive understanding of human-AI deliberation.
+8. Conclusion
+In this paper, we introduce Human-AI Deliberation, as a new paradigm of AI assistance for decision-making.
+Human-AI Deliberation encourages the externalization of thoughts, facilitates interactive deliberation between humans and AI, and allows for dynamic updates of decisions.
+To enable the deliberation, we present a novel AI assistant called Deliberative AI, which can identify differences in viewpoints, engage in comprehensive deliberation, and adapt its suggestions during discussions.
+We apply this architecture to an illustrative task (graduate admissions decisions) and conduct an exploratory study to assess its potential impact on decision-making processes, outcomes, user perceptions, and experiences.
+Results indicate the potential of Deliberative AI to improve decision accuracy and promote more appropriate human reliance on AI.
+Additionally, we analyze participants’ open-ended feedback to gain deeper insights into how users use and perceive Deliberative AI, uncovering areas for improvement.
+With the key insights and implications derived from our study, we aim for this work to serve as an exploratory step toward establishing a new paradigm of AI assistance that enhances decision-making.
+Acknowledgements.
+This work is supported by the Research Grants Council of the Hong Kong Special Administrative Region under General Research Fund (GRF) with Grant No. 16207923.
+Appendices
+Appendix A Baseline XAI Interface
+Figure 10 shows the baseline XAI interface.
+Figure 10.
+The baseline XAI (traditional explainable AI) interface in our user study.
+The interface contains three parts.
+The top (A) is the applicant’s profile.
+The bottom left part (B) shows AI’s feature contribution explanation.
+The bottom right part (C) is for humans to (1) indicate their initial predictions, (2) see AI’s suggestions, and (3) indicate their final decisions.
+Note that AI’s suggestions and explanations are only shown after humans make their initial predictions.
+(All the dashed lines are only for illustration)
+[…]
+Appendix B An Example Dataflow of Deliberative AI
+Figure 11 provides the details of a conversation between a human and a Deliberative AI discussing how an applicant’s GPA affects admissions chances. Here’s a step-by-step breakdown:
+  1. (1)
+    The user inputs GPA-related arguments in the dialogue interface.
+  2. (2)
+    The system packages the user’s input as a prompt for the Intention Analyzer in the Communication Layer.
+  3. (3)
+    The Intention Analyzer recognizes attributes and intentions and saves in JSON format, then forwards it to the Knowledge Extractor in the Control Layer.
+  4. (4)
+    The Knowledge Extractor generates a query function and fetches statistical results from the DS-Model and training data.
+  5. (5)
+    The statistical results are transmitted to the Regulator.
+  6. (6)
+    Regulator crafts a constraint prompt ensuring consistency between the LLM’s output and the DS-Model’s prediction, feeding it to the LLM-based Deliberation Facilitator.
+  7. (7)
+    The Deliberation Facilitator generates responses to the user’s initial arguments.
+Figure 11.
+An illustration of how Deliberative AI processes humans’ inputs and how it generates outputs.
+The prompts used are simplified in this figure for illustration purposes.
+(for the complete prompts, please check our supplementary materials)
+[…]
+Appendix C Detailed Metrics and Questions
+Table 2 shows the detailed metrics and questions used in our measurement.
+Table 2. Measurements used in our user study. We collected participants’ objective decision data, subjective questionnaire data, and qualitative open-ended feedback.
 Aspect Metrics Detailed Meaning and Questions
@@
 $\rm\frac{Number\ of\ incorrect\ human\ final\ decisions\ with\ correct\ AI\ suggestions}{Total\ number\ of\ correct\ AI\ suggestions}$
-Behaviors Opinion Changes WOA = $\rm\frac{human\ finally\ predicted\ admission\ chance-human\ initially\ predicted\ admission\ chance}{AI\ initially\ predicted\ admission\ chance-human\ initially\ predicted\ admission\ chance}$
-Frequency of Opinion Changes Count of opinion changes per prediction.
-Dialogue Rounds Number of conversation rounds in a decision-making task.
-Conflicting Dimensions Count of dimensions with conflicting information.
-Dimensions Discussed Count of dimensions considered in a prediction.
-Time Spent Average time spent per decision.
 Subjective Measures
@@
 Potential Improvement “To make a better discussion, which parts do you think the current AI needs to be improved, and how should it be improved?”
-5.4. Measurement
-To comprehensively assess the impact of human-AI deliberation, we evaluated the following aspects. Detailed meanings and questions for each metrics can be seen in Table 3.
-  • Task Performance. We evaluated decision accuracy using Decision Accuracy (Zhang et al., 2020; Bansal et al., 2021; Ma et al., 2023).
-  • Reliance.
-We assessed participants’ reliance on AI suggestions using Agreement Fraction (Zhang et al., 2020; Ma et al., 2023; He et al., 2023) and Switch Fraction (Zhang et al., 2020; He et al., 2023).
-Also, we measured the appropriateness of human reliance by Over-reliance Ratio (Wang and Yin, 2021; Visser et al., 2014; Parasuraman and Riley, 1997) and Under-reliance Ratio (Wang and Yin, 2021; Visser et al., 2014; Parasuraman and Riley, 1997).
-  • Behaviors.
-We examined participants’ Opinion Changes (Bonaccio and Dalal, 2006; Poursabzi-Sangdeh et al., 2021; Gino and Moore, 2007), Frequency of Opinion Changes (Yaniv, 2004), Conflicting Dimensions, Dimensions Discussed, and Time Spent (Yang et al., 2020; Abdul et al., 2020).
-Specifically, for Opinion Changes, we use weight of advice (WOA) (Gino and Moore, 2007) to measure how humans’ overall opinions are influenced by AI’s opinions [^5] WOA is 0 when a human’s final prediction matches their initial one, 1 when it matches the AI’s prediction, and 0.5 when the individual averages their estimate with AI advice.
-A positive WOA means human opinions align more with AI (with values above 1 indicating over-adjustment), while a negative WOA signifies a divergence from AI opinions..
-  • Perceptions of AI.
-We measured participants’ perceived Helpfulness (Laugwitz et al., 2008; Cai et al., 2019a; Buçinca et al., 2020), Trustworthiness (Ghai et al., 2021; Buçinca et al., 2021), and Understanding (Wang and Yin, 2021) through 7-point Likert scale (1: Strongly disagree; 7: Strongly agree.
-The same below).
-  • User Experience.
-We measured participants’ Decision Confidence (Miller et al., 2015).
-Meanwhile, as deliberation requires extra efforts, we evaluated participants’ self-reported Mental Demand (Hart, 2006; Lai et al., 2022; Ghai et al., 2021; Buçinca et al., 2021), Effort (Hart, 2006), Complexity (Buçinca et al., 2021), and Satisfaction (Buçinca et al., 2021; Ghai et al., 2021) via 7-point Likert scale.
-  • Open-ended Feedback.
-As we aim to gain an in-depth understanding of participants’ perceptions of the Deliberative AI and the deliberative decision-making process, we added three open-ended questions in the exit survey.
-5.4.1. Analysis Methods
-We conducted mixed-methods analyses to examine the aforementioned metrics.
-For the quantitative analysis, we first performed normality tests (Shapiro-Wilk) and found that the data did not fit the normality assumption.
-Therefore we ran the non-parameter tests.
-Specifically, to compare Deliberative AI and XAI (such as humans’ reliance on AI, and their perceptions of AI), we run Mann-Whitney U test.
-To compare all three conditions (such as task performance, and user experience), we employed Kruskal–Wallis tests with Bonferroni post-hoc correction.
-For qualitative analysis, two authors independently coded participants’ open-ended feedback and conversation logs using an inductive thematic analysis approach (Hsieh and Shannon, 2005).
-The final themes emerged through discussions and harmonization over several iterations.
-We also identified representative examples from the source texts for demonstration in this paper.
-6. Results
-In this section, we report our exploratory findings regarding the four research questions: (RQ1) how Human-AI Deliberation affects task performance and human reliance (and reliance appropriateness) on AI, (RQ2) how Human-AI Deliberation affects human perceptions and task experience, (RQ3) how humans engage in the deliberation process, and (RQ4) how humans perceive the effectiveness of Human-AI Deliberation and what should be improved.
-6.1. RQ1: How will Human-AI Deliberation affect task performance and humans’ reliance (and its appropriateness) on AI suggestions?
-6.1.1. Decision Accuracy
-As shown in Figure 7, we find that participants in the Deliberative AI condition exhibited significantly higher decision accuracy ($M$=0.598, $SD$=0.169) compared to those in the XAI condition ($M$=0.524, $SD$=0.16, $p$¡0.05).
-Although there is no statistically significant difference, a trend emerges where Deliberative AI tends to surpass Human Alone performance, while XAI generally underperforms compared to Human Alone.
-This finding indicates that in scenarios where tasks are difficult for both humans and AI, traditional Explainable AI (XAI) might not enhance performance and could even lead to detrimental effects.
-Conversely, Human-AI Deliberation has the potential to provide positive benefits, even though the performance of underlying AI models in these challenging task cases is not good enough.
-Figure 7. Task performance in different conditions. The error bars represent 95% confidence interval. (*: $p$¡0.05, **: $p$¡0.01, ***: $p$¡0.001)
-[…]
-6.1.2. Reliance
-We measured participants’ objective reliance by agreement fraction and switch fraction.
-As shown in Figure 8 (a), participants agreed significantly less with AI’s suggestions in Deliberative AI ($M$=0.57, $SD$=0.24) than in XAI ($M$=0.68, $SD$=0.27, $p$¡0.05), and switched significantly less to AI’s predictions in Deliberative AI ($M$=0.23, $SD$=0.35) than in XAI ($M$=0.51, $SD$=0.41, $p$¡0.001).
-Combined with participants’ open-ended feedback (Sec.
-6.4), this may be because people invest more in independent thinking in the process of deliberation with AI and realize the problematic aspects of AI’s perspective.
-6.1.3. Appropriateness of Reliance
-We further measured the appropriateness of participants’ reliance on AI’s suggestion by under-reliance and over-reliance (Figure 8 (b)).
-Results show that there is no significant difference between Deliberative AI ($M$=0.32, $SD$=0.28) and XAI ($M$=0.29, $SD$=0.30) in terms of under-reliance.
-While significantly less over-reliance was observed in Deliberative AI ($M$=0.47, $SD$=0.31) than in XAI ($M$=0.65, $SD$=0.33, $p$¡0.001), which means that participants had more appropriate reliance on AI when collaborating with our proposed Deliberative AI.
-Figure 8.
-Participants’ reliance and the appropriateness of their reliance.
-(A) Participants’ reliance on AI’s suggestions was measured by agreement fraction and switch fraction.
-(B) The appropriateness of participants’ reliance on AI’s suggestions, including under-reliance (the ratio where participants did not use a correct AI suggestion) and over-reliance (the ratio where participants used an incorrect AI suggestion).
-The error bars represent 95% confidence interval.
-(*: $p$¡0.05, **: $p$¡0.01, ***: $p$¡0.001)
-Figure 9. Learning curve measured by decision accuracy in each task round. The error bars represent 95% confidence interval.
-[…]
-6.1.4. An exploratory analysis of learning effects
-In addition to the accuracy and reliance, we observed learning effects when humans worked with Deliberative AI to make decisions (Figure 9).
-To quantify the effect, following (Cabrera et al., 2023) we fitted a linear model of average participant accuracy for each decision round.
-We found a learning effect trend for Deliberative AI (although only marginally significant, $p$=0.051).
-However, no significant learning effects were found in XAI.
-This result reveals participants may gain better knowledge about the task and how to collaborate with the AI partner to deal with the task as the deliberation progresses.
-In summary, Deliberative AI has a positive effect in improving decision accuracy.
-Besides, since the AI model’s accuracy is only 50%, participants in Deliberative AI relied less on AI’s incorrect suggestions which in turn leads to more appropriate reliance (less over-reliance).
-However, our current results do not mean that Deliberative AI will also lead to better decision performance and more appropriate human reliance than XAI when the AI performance is high.
-Future work is needed to further explore the effects when the AI performance varies.
-6.2. RQ2: How will Human-AI Deliberation affect humans’ perceptions of the AI and user experience?
-We measured the effects of different AI conditions on participants’ perceptions and user experience via a 7-point Likert scale (1: strongly disagree, 7: strongly agree).
-Figure 10. Participants’ perceptions of the AI assistant. The error bars represent 95% confidence interval. (*: $p$¡0.05, **: $p$¡0.01, ***: $p$¡0.001)
-[…]
-Figure 11. Effects on user experience. The error bars represent 95% confidence interval. (*: $p$¡0.05, **: $p$¡0.01, ***: $p$¡0.001)
-[…]
-6.2.1. Perceptions of AI
-Figure 10 shows participants’ perceptions of the AI model.
-There were no significant differences in perceived helpfulness and understanding between Deliberative AI and XAI.
-However, participants reported significantly less trust in Deliberative AI ($M$=4.47, $SD$=1.68) compared to XAI ($M$=5.52, $SD$=1.27, $p$¡0.01), aligning with their reliance behaviors (Sec.
-6.1.2).
-This difference may be attributed to participants identifying more AI flaws through deliberation than by solely observing AI’s explanations, supported by conversation logs analysis (see Sec.
-6.4 for details).
-6.2.2. User experience
-First, we want to see participants’ decision confidence.
-As indicated in Figure 11, participants in XAI reported significantly higher confidence ($M$=6, $SD$=1.10) in their predictions than those in Human Alone ($M$=5.59, $SD$=1.12, $p$¡0.05).
-However, from Figure 7 (a) we found that the final accuracy of participants in XAI is even lower than those in Human Alone (although not significant).
-This indicates that the traditional XAI might lead to humans’ illusionary confidence, which could prevent humans from making optimal decisions.
-Given that the Deliberative AI requires participants to externalize thoughts at a dimension level and engage in deliberative discussions on conflicting opinions, it’s crucial to explore how these activities influence the user experience.
-Results showed no significant difference among the three conditions concerning Mental Demand, Effort, and Perceived System Complexity.
-However, we find participants reported significantly lower Satisfaction in Deliberative AI than in Human Alone.
-Combined with the qualitative results (Sec.
-6.4), we speculate that this may be because Deliberative AI exposes more conflicts between humans and AI and they need to carry out deliberation around these conflicts.
-This can easily reduce people’s satisfaction compared to simply making a decision.
-We also observed a trend that Deliberative AI is more complex than the two baselines.
-These findings suggest that there is a trade-off between encouraging users’ deliberative thinking and optimizing their user experience, in line with findings in previous studies (Buçinca et al., 2021).
-Figure 12. Descriptive statistics of participants’ decision-making process. (a) Descriptive statistics of conversations in Deliberative AI. (b) Time spent per decision in the three conditions.
-Figure 13. The changes in participants’ opinions measured by Weight of Advice. (*: $p$¡0.05, **: $p$¡0.01, ***: $p$¡0.001)
-[…]
-6.3. RQ3: How will humans engage in the deliberation process when working with Deliberative AI?
-To gain deeper insights into the impact of our Deliberative AI, we analyzed how participants engaged with this technology during the decision-making process.
-As shown in Figure 12 (a), each decision-making instance typically involved a median of 2.0 rounds of conversation between participants and the Deliberative AI.
-During these discussions, participants’ viewpoints changed a median of 1.0 times, and they encountered a median of 1.0 conflicting dimension with the AI model (based on our pilot study, we defined “conflicting” by a 15% opinion difference threshold).
-It’s interesting to note that although Deliberative AI only highlighted a median of 1 conflicting dimension, participants discussed a median of 2.0 dimensions with the AI during these interactions, indicating proactive engagement in conversations, even when not explicitly prompted.
-Regarding the time spent in making a decision, as shown in Figure 12 (b), participants took the most time to make a decision in Deliberative AI (a median of 236 seconds), followed by a median of 33 seconds in XAI, and a median of 23 seconds in Human Alone, all pair-wise comparisons are significant ($p$¡0.001).
-These statistics indicate that humans make “quick” decisions with traditional XAI.
-Compared to without AI assistance, participants only spent an average of 10 seconds longer on each decision in the XAI condition.
-This also echoes previous findings that people will mistakenly regard explanations as a manifestation of AI capabilities (Bertrand et al., 2022), resulting in over-reliance.
-Since in Deliberative AI humans need to spend more time in the deliberative discussion process, it might force humans to make “slow” decisions (i.e., invoke more System 2 thinking (Kahneman, 2011)), which is more appropriate for high-stakes decision-making.
-We also assessed how participants’ opinions were influenced by AI’s opinions in both the Deliberative AI and XAI conditions by calculating the Weight of Advice (WOA).
-Figure 13 shows a clear contrast: the WOA in the Deliberative AI condition had a notably lower median (0.085) compared to XAI (median: 0.538, $p$¡0.001).
-This indicates that participants in the Deliberative AI condition tended to rely more on their own predictions.
-It’s noteworthy that the lower quartile for Deliberative AI is 0, implying that at least 25% of participants did not alter their opinions after seeing AI suggestions.
-These findings align with participants’ lower reliance on AI observed in Figure 8.
-6.4. RQ4: How will humans perceive the effectiveness of the proposed Human-AI Deliberation and what can be improved for future Human-AI Deliberation design?
-In addition to quantitative measures, we aim to comprehend participants’ in-depth perceptions regarding the helpfulness of the proposed Human-AI Deliberation and the feature of Deliberative AI (especially for the opinion updating feature).
-We also seek valuable insights to enhance the future design of Human-AI Deliberation.
-We analyzed participants’ open-ended feedback using their conversation logs as support.
-We summarize our key findings in Figure 14.
-Figure 14. The main results of our thematic analysis of the open-ended questions.
-[…]
-6.4.1. Perceptions of the Deliberative AI and the discussion process
-Overall, participants provided both positive and negative feedback on Deliberative AI and the deliberative discussion process.
-43 out of 48 participants thought the deliberative discussion helped them make more informed decisions.
-Specifically, discussion with AI help humans “identify AI’s problems” (21/48), “consider different aspects” (10/48), “reflect and realize their mistakes” (15/48).
-For instance, Participant 1 (Male, 32) pointed out issues with AI’s knowledge:
-”The AI relies way too much on GPA scores but underestimates the role of recommendation letters.
-It did not give me convincing justifications.
-I cannot rely too much on the AI’s opinion on these two dimensions.” (P1, Male, 32)
-Participant 19 (Male, 42) appreciated how the discussions inspired self-reflection:
-”The AI can cause you to doubt what you believe to be sufficient for an opinion, especially according to the information/data given.” (P19, Male, 42)
-These findings align with the content of the human-AI conversations we collected.
-In 32% of the conversations, participants expressed doubts or questions to AI, in 17% of the conversations, participants acknowledged AI’s arguments, and in 15% of the conversations, participants engaged in reflection and self-correction.
-Besides, 18 participants acknowledged that discussing with AI can “offer new knowledge, insights, and fresh perspectives”. For example, P5 (Male, 45) said
-”It provided some information that I actually did not know (such as the percentile, how similar stats impacted other decisions in the past, etc.), which I thought were extremely helpful.”
-Moreover, nine participants explicitly acknowledged that discussion “helps them identify biases”. For example,
-”There were times where it made me realize the AI probably had an inherent bias so all in all I think it helped to actually make me pause and reflect.” (P35, Female, 29)
-This observation is substantiated by an examination of participants’ conversation histories during the discussion process.
-Remember that the AI model assigns a positive feature importance to [Country belongs to the US] while a negative feature importance to feature [Country belongs to Asia].
-We are gratified to note that a substantial portion of participants (15 out of 48) identified biases in AI’s opinions.
-For instance, P3 (Male, 42) questioned, “Why does their country matter?
-They are being penalized for being from a certain country, which could be interpreted as being racist.” Similarly, P17 (Female, 60) inquired, “Could you explain why you see an engineering major and Asian origin as negatives?” These insightful conversations indicate that with Deliberative AI, participants could identify the AI model’s biases in the decision-making process.
-What’s more, 15 participants thought the discussion “promoted a balance between AI’s objective knowledge and human’s subjective opinions”. For example,
-”It provides an opportunity to consider multiple perspectives and statistical data, which can lead to a more informed and balanced decision-making process.” (P2, Male, 42)
-”I think that receiving the feedback with the data behind the AI is a helpful visual reminder of how much above/below the average the applicant was.
-Being able to rely on the AI for this and focusing on my own opinions that are more subjective helped me to balance.” (P48, Male, 28)
-However, we also identified limitations in the current discussion process based on participants’ feedback.
-Regarding the discussion’s limitations, five participants felt their “agency gets reduced” as the AI prompted them to think other than passively waiting for their questions.
-Additionally, four participants found the deliberative discussion “mentally demanding” and believed the AI’s responses were overly verbose.
-These findings align with the results of the effects on participants’ user experience (see Sec.
-6.2.2).
-Regarding the Deliberative AI’s limitations, two common concerns were that it “relies too heavily on data over logic” (mentioned by 15 participants) and “fails to consider human reasoning” (mentioned by 10 participants).
-This feedback aligns with our analysis of participants’ conversations, where humans integrated personal experiences and logic into their decision-making.
-For instance, they considered factors such as “Business applications are highly competitive, so a higher GPA is required.” (P17, Female, 60), “A good SOP represents that the applicant has an in-depth understanding of the project and a strong willingness.” (P32, Male, 28), “Middle-ranking schools may give students higher GPA, so a GPA of 3.6 may not have a particularly positive impact.” (P29, Male, 50).
-In contrast, the AI’s responses lacked in-depth analysis of these human justifications.
-Furthermore, some participants expressed concerns about the AI’s inability to remember historical discussions (five), its failure to consider the bigger picture (three), and its lack of transparency (three).
-6.4.2. Perceptions of Deliberative AI’s updating
-Participants generally appreciated the feature of dynamical updating.
-25 out of 48 participants found that the AI updates “improved the user experience.” They described the experience as more interactive, akin to “interacting with a real person” (6 participants), where their opinions were “listened to and respected” (19 participants).
-Furthermore, these dynamic updates made the discussion process “more engaging” (8 participants) and the overall decision-making process “more collaborative” (5 participants).
-However, perceptions of how AI updated its opinions varied.
-While 11 participants felt the magnitude of AI updates was too sensible (e.g., “the AI updated too much”), 13 participants expressed uncertainty about “why and how the AI updated.” For instance, one participant mentioned,
-”I’m not sure if the AI really changed its opinions because of what I said or because it was programmed to do so in response to my judgments”. (P2, Male, 42)
-Interestingly, some participants believed that AI changed its views too easily, while others felt it didn’t compromise enough.
-Additionally, AI’s opinion updates influenced participants’ perceptions of AI.
-Eight participants believed that “AI could take qualitative information into account”, and five participants thought “AI is learning human knowledge from the discussion process”, which could “improve the quality of decision-making” (3 participants).
-Three participants reported that the lack of transparency in the AI update process made them “doubt the quality of the opinions given by the AI at the beginning”.
-Therefore, although AI updates imitate the behavior of deliberation between humans, in order to make AI updates more meaningful, it is necessary to consider the perceptions of different users and conduct a more targeted design.
-6.4.3. Opportunities for future improvements of Deliberative AI
-Participants offered insightful suggestions for enhancing Deliberative AI in the future.
-  • Deeper Understanding of Humans: 15 participants stressed the need for AI to “better comprehend humans’ thoughts and arguments”, desiring more nuanced and context-aware communication.
-  • Enhanced Memory: 8 participants emphasized the importance of AI “remembering previous interactions for more personalized and coherent conversations”.
-  • Holistic Problem Solving: 12 participants recommended that AI “think beyond statistics” and “provide arguments with broader considerations”, offering more comprehensive suggestions.
-  • Deeper Analysis: 10 participants advocated for AI to delve deeper into discussions, “conducting more thorough probing and “analyzing the implications beyond the superficial meaning of an attribute within a given context”.
-  • More Transparency: 8 participants expressed the desire for more transparency from AI, seeking explanations for “why AI’s opinions changed during conversations”.
-This transparency was seen as essential for building trust and understanding.
-Regarding the design of the human-AI discussion process, 8 participants suggested making the conversation “more conversational and human-like”, with 11 participants favoring “concise AI responses”.
-Additionally, 2 participants recommended “progressively disclosing AI’s arguments”, while 3 participants desired more user agency in leading the discussion process.
-In addition, 3 participants also suggested improving the dialogue generation speed and visual information display.
-Things to Note about Interpreting the Results.
-We explored the possible impact of Human-AI Deliberation on decision-making from different aspects.
-However, we caution readers to consider these results as exploratory findings.
-Since we did not recruit real members of the graduate admissions committee to participate in our study, we do not know whether Deliberative AI can help with real graduate admissions decisions.
-In addition, the generalizability of these results has yet to be verified.
-In our study, Deliberative AI only had 50% accuracy on the selected “challenging” task cases, so it is uncertain whether similar findings can appear on “easy” cases where AI has high performance.
-In addition, due to the length of discussion, we only asked participants to solve four task cases with Deliberative AI, which also limits the generalizability of the results.
-Nonetheless, we believe these exploratory findings are valuable in providing an initial understanding of the potential of Deliberative AI and the impact it may have on decision-making.
-Common issues identified in this study can shed light on future designs of Deliberative AI.
-7. Discussion
-In this paper, we introduce Human-AI Deliberation, which aims to address two common challenges in AI-assisted decision-making.
-First, we tackle the issue of humans’ insufficient analytical thinking when a recommendation is directly presented.
-Second, we address the problem of limited interaction support for resolving human-AI disagreements.
-Existing interfaces only allow individuals to accept or reject AI suggestions as a whole, hindering the utilization of human-AI nuanced knowledge and collaborative problem-solving.
-Our exploratory assessment in the context of an illustrative task reveals that Human-AI Deliberation has the potential to enhance decision accuracy and promote appropriate reliance on AI recommendations compared to traditional explainable AI assistants.
-Analysis of questionnaire ratings, human behaviors, and open-ended feedback provides valuable insights.
-We now delve into our key findings and discuss design implications and critical limitations.
-7.1. Deliberation as a New Paradigm Complementing Existing (X)AI Assistance
-The proposed Human-AI Deliberation aims at engaging humans in the deliberation process through conflict-oriented discussions.
-It can serve as a powerful supplement to existing AI assistance.
-On the one hand, Deliberative AI still plays the role of an assistant and its underlying component is based on existing XAI.
-Although it will refute, it always respects human opinions.
-People have the full agency as the final decision-maker.
-Additionally, Deliberative AI can be seen as an enhanced version of an AI assistant that engages people in analytical thinking.
-Some interaction design of Deliberative AI aligns with prior work, such as encouraging participants to form independent opinions before seeing AI suggestions (e.g., Cognitive Forcing Function (Buçinca et al., 2021)) and utilizing AI-generated questions to stimulate critical thinking (e.g., employing AI-framed questioning (Danry et al., 2023)).
-On the other hand, Human-AI Deliberation makes important contributions to stepping outside the boundaries of traditional AI/XAI for AI-assisted decision-making.
-People are no longer just recipients of AI suggestions and explanations, but need to actively participate in discussions and explain their own ideas, which contributes to mutual communication and transparency between humans and AI.
-This communication can lead to more informed and mature decisions as deliberations progress.
-We suggest that the design of future AI assistants should not only focus on the design of the AI side (such as improving the performance of AI or improving the presentation of AI auxiliary information), but also draw on human-to-human decision-making methods and theories to explore design opportunities, especially from the perspective of communication between humans and AI.
-7.2. Reducing Human Over-Reliance by Exposing AI Mistakes
-Human-AI Deliberation significantly diminishes participants’ tendency to over-rely on incorrect AI suggestions.
-A pivotal element is the adoption of cognitive forcing theory, as proposed by Buccinca et al.
-(Buçinca et al., 2021), which motivates individuals to formulate their opinions before being exposed to AI recommendations.
-This method effectively counters the anchoring bias—the initial influence of AI’s erroneous predictions on judgment—thereby encouraging more analytical, System 2 thinking  (Kahneman, 2011).
-Notably, our Explainable AI (XAI) baseline also requires participants to articulate their perspectives independently before receiving AI counsel and explanations.
-Despite this stringent baseline, our Deliberative AI approach shows enhanced effectiveness in reducing over-reliance.
-Thus, the observed decrease in over-reliance within our Deliberative AI extends beyond mere cognitive forcing, attributed also to the promotion of thoughtful deliberation between individuals and AI amidst contrasting viewpoints.
-Within the Deliberative AI, as participants navigate through conflicting opinions with AI, they become increasingly cognizant of AI limitations, further mitigating over-reliance.
-Furthermore, 31% of conversations involve participants expressing skepticism towards the AI’s opinions and logic.
-However, the decline in over-reliance is not due to people blindly doubting AI.
-Notably, our strategy significantly reduces over-reliance on AI without increasing under-reliance.
-This indicates that through deliberative discourse, participants critically assess AI’s opinions on a case-by-case basis, integrating accurate AI insights while amending their own assessments as necessary and identifying instances where AI judgments may be erroneous.
-Consequently, we recommend that designers of AI-assisted decision-making systems prioritize transparency and explicitly highlight AI’s potential errors, rather than merely hinting that “AI may make errors”.
-7.3. Human-AI Conflict Resolving Is The Key to Improving Collaboration
-We contend that resolving conflicts is inherently more beneficial for enhancing decision-making processes than merely seeking consensus.
-Our methodology, Human-AI Deliberation, emphasizes the resolution of disputes between humans and AI, a dimension presently overlooked in AI-assisted decision-making frameworks.
-We argue that conflict serves as a crucial element in human-AI collaboration, unveiling inherent flaws and biases.
-This conflict-centric approach yields multiple advantages: it enhances the accuracy of decision-making by mitigating over-reliance on AI, fosters introspection, and facilitates the reconciliation of differing opinions with AI recommendations.
-Furthermore, by prioritizing conflict, humans can better recognize biases in AI interpretations, thereby advancing fairness in decision-making processes (Kramer et al., 1990; Hochman et al., 2015).
-Nevertheless, an emphasis on conflict resolution—despite its apparent benefits—may diminish user satisfaction.
-Engaging in deliberation with AI, particularly around divergent viewpoints, necessitates that individuals critically evaluate their own perspectives and articulate their arguments.
-This process amplifies the perception of ”dissenting voices,” potentially compromising the user experience.
-Hence, we advocate for a balanced approach in future design paradigms: one that centers on conflict resolution while simultaneously considering strategies to enhance user engagement with conflicts without detracting from the overall experience.
-7.4. Obstacles to Discussion: Humans and AI Think Differently
-Humans Use Heuristics and Logic, while AI Relies on Data: Combining LLM and DS-Model empowers Deliberative AI in dynamic discussions with humans.
-However, AI’s data-centric approach clashes with human decision-making, which integrates personal understanding and experiences (Ghattas et al., 2014), logic (Evans, 2002), heuristics (Doswell, 2004), and creativity (Simon, 1990, 1997).
-To address this, we propose DS-Model guiding LLMs for data-related discussions and LLMs autonomously handling data-irrelevant matters based on DS-Model’s predictions.
-While effective, this approach faces limitations in engaging users with strong subjective perspectives and reasoning (Liu et al., 2023).
-Existing Explanation Methods Are Insufficient to Enable Discussion: Current XAI methods, like local feature importance (Liao and Varshney, 2021), inadequately clarify specific feature contributions.
-Our solution, based on training data, provides data-related insights, such as data distribution and value comparisons, to empower AI in feature-level discussions.
-However, challenges arise.
-First, patterns inferred from data may not precisely align with AI’s explanations.
-Second, users often rely on various forms of reasoning beyond data, including subjective experiences (Doswell, 2004; Ghattas et al., 2014).
-This finding is aligned with Miller et al.’s work (Miller, 2019), ”probabilities are not as important as causal links.” Hence, offering data-driven insights as meta-explanations may not fully align with users’ thinking styles or replicate human-like conversations.
-Therefore, future human-AI deliberation designs should try to cope with humans’ intuition and cognitive processes and design human-compatible AI explanations to empower more effective human-AI discussions.
-7.5. Ethical Concerns in Deliberative AI Design
-Responsibly Utilizing LLMs for Assisting Human Decision-Making.
-LLMs are advancing rapidly but face a critical issue known as “hallucination”, where they generate seemingly credible but false information (Ji et al., 2023).
-Engaging in discussions with users solely based on LLM-generated opinions, without proper fine-tuning of task data, is highly irresponsible.
-Even with fine-tuning, unpredictable responses can occur.
-To address this, we advocate for a responsible approach that combines LLMs and DS-Model.
-DS-Model’s output tightly controls LLMs’ responses, with LLMs primarily acting as a communication bridge between users and DS-Model.
-While this approach restricts LLMs’ flexibility to some extent, it enhances security and controllability.
-However, it’s essential to note that even with DS-Model and training data guiding LLMs’ responses, the potential for error information remains.
-Therefore, researchers planning to employ LLMs in decision-making should exercise caution and transparency.
-It’s crucial to inform users about the possibility of inaccurate information and potential pitfalls associated with LLM-generated responses.
-Concerns Regarding AI Updating.
-Research on human-AI discussions highlights a desire for AI to actively engage with human opinions (Zheng et al., 2023).
-To address this, we’ve implemented an AI opinion update mechanism, allowing AI to adapt its stance based on user arguments’ rationality and prediction uncertainty.
-Although users generally appreciate the AI’s responsiveness and the feeling of being heard and respected, ethical concerns have emerged, particularly regarding accountability.
-One pivotal question arises: Who should be held accountable if the AI initially had the correct prediction but updated to an incorrect one after engaging in discussions with the human?
-Furthermore, the AI’s ability to adapt its viewpoints may give humans the impression of learning and progress, despite the underlying AI model remaining unchanged.
-This can impact user expectations and potentially lead to an overestimation of AI capability.
-While the AI’s capacity to update opinions is promising, it’s crucial to remain vigilant about potential adverse outcomes during the design phase.
-Most importantly, future designs should prioritize transparency in the AI updating mechanism to help users fully comprehend the principles governing AI updates and effectively manage their expectations.
-7.6. Applicability and Generalizability
-Although the human-AI deliberation framework we introduced is potentially applicable and can theoretically be adapted to various decision-making scenarios, deliberation may not be suitable for all tasks.
-First, the deliberative process requires additional time and cognitive resources, making it less suitable for repetitive, low-stakes decisions such as content moderation (Lai et al., 2022).
-Instead, it may be more beneficial in high-stakes decision-making environments, such as healthcare (Lee et al., 2021), financial investments (Green and Chen, 2019), and criminal justice (Chiang et al., 2023), especially when dealing with complex or edge cases.
-Secondly, Human-AI Deliberation is better suited to handle challenging task cases that are difficult for both humans and AI.
-Furthermore, our proposed method of combining large language models (LLMs) with domain-specific (DS) models can be adapted to other decision-making scenarios.
-It does not limit the form of the DS model, which can range from a simple machine learning classifier to a complex neural network model or even another specific large model.
-7.7. Limitations and Future Work
-The study design of this paper has several limitations.
-First, the illustrative task (college admission decision) we chose disconnects with real college admission.
-On the one hand, limited by the available public data set, the data set we used does not contain the raw texts of some important dimensions, such as purpose statements, diversity statements, recommendation letters, etc.
-Thus, in our study, participants could only discuss the ratings on them rather than what information is included in them.
-Furthermore, the participants we recruited were not members of the actual admissions committee.
-Although we require participants to have at least a master’s degree and graduate application experience, there is still a considerable gap in expertise.
-Therefore, the graduate admission task is just an illustrative task to explore the instantiation and potential of the proposed Human-AI Deliberation framework.
-We need future studies to truly explore whether and why Deliberative AI yields improvement in real-world college admission decision-making.
-Second, the number of decision task samples used in our study is relatively limited.
-We tried to include 8 or 10 decision task samples in our pilot study, which resulted in participants taking more than one hour to complete all decision tasks.
-In addition, the users’ engagement dropped substantially after completing three or four tasks, and they simply dragged the slider on the interface to give their opinions without discussing the conflicting points with AI.
-We infer that long tasks lead to user fatigue, which will affect the validity of the study.
-Therefore, due to concerns about the length of the study, we only selected four task samples, which to some extent limits the generalizability and significance reliability of the experimental results.
-Future work needs to test the effects of Human-AI Deliberation through a long-term study so that more deliberative decision data can be collected.
-Third, the AI in our study is of low performance (50% accuracy), not a conventional AI that performs better than humans.
-We made this design choice because Human-AI Deliberation is more suitable for dealing with task cases that are challenging for both humans and AI.
-However, we acknowledge that the decision-making dynamics assisted by AI with different performances can be different.
-Thus, all our results are based on the premise that AI’s performance is relatively low and on par with humans’ independent decision-making performance.
-Therefore, we cannot say whether some of the positive outcomes of Deliberative AI (such as a reduction in people’s overdependence) will persist when the AI performs very well.
-Future work can further consider the variation of human and AI decision performance.
-Finally, our study shows the potential of Deliberative AI compared to XAI, but it is still unclear which components of Deliberative AI contribute to the benefits.
-We acknowledge that Deliberative AI has more features than the XAI condition.
-For example, Deliberative AI can present distribution information of the training data (e.g., percentiles) in the conversation (as evidence of AI in the Human-AI Deliberation framework), which XAI does not have.
-This is because we chose to apply the most commonly used XAI display method as the baseline.
-These XAI systems typically only provide suggestions and explanations without providing data distribution information.
-In addition, Deliberative AI consists of multiple critical components.
-A future ablation study could help investigate the role and importance of each component for a systematic comparison.
-8. Conclusion
-In this paper, we introduce the framework of Human-AI Deliberation, as a new paradigm of AI assistance for decision-making.
-Human-AI Deliberation encourages the externalization of thoughts, facilitates interactive deliberation between humans and AI, and allows for dynamic updates of decisions.
-To enable the deliberation, we present a novel AI assistant called Deliberative AI, which can identify differences in viewpoints, engage in comprehensive deliberation, and adapt its suggestions during discussions.
-We apply this framework to an illustrative task (graduate admissions decisions) and conduct an exploratory study to assess its potential impact on decision-making processes, outcomes, user perceptions, and experiences.
-Results indicate the potential of Deliberative AI to improve decision accuracy and promote more appropriate human reliance on AI.
-Additionally, we analyze participants’ behaviors and open-ended feedback to gain deeper insights into how users use and perceive Deliberative AI, uncovering areas for improvement.
-Lastly, we offer design implications based on our findings, highlighting the potential of Human-AI Deliberation as a valuable supplement to current AI-assisted decision-making.
-Appendices
-Appendix A Base Interface
-Figure 15.
-The baseline (traditional explainable AI) interface in our user study.
-The interface contains three parts.
-The top (A) is the applicant’s profile.
-The bottom left part (B) shows AI’s feature contribution explanation.
-The bottom right part (C) is for humans to (1) indicate their initial predictions, (2) see AI’s suggestions, and (3) indicate their final decisions.
-Note that AI’s suggestions and explanations are only shown after humans make their initial predictions.
-(All the dashed lines are only for illustration)
-[…]